### v1

# On Selection of Cross-Section Averages in Non-stationary Environments

Abstract
Information criteria (IC) are important tools in the literature of factor models that allow one to estimate a typically unknown number of latent factors. Although first proposed for the Principal Components setting in the seminal work by Bai and Ng, (2002), it has recently been shown that IC perform extremely well in Common Correlated Effects (CCE) and related setups with stationary factors. In particular, they can consistently select a sufficient set of cross-section averages (CAs) to approximate the factor space. Given that CAs can proxy nonstationary factors, it is tempting to believe that the consistency of IC continues to hold under such generality. This study is a cautionary tale for practitioners. We demonstrate formally and in simulations that IC has a severe underselection issue even under very mild forms of factor non-stationarity, which goes against the sentiment in the CAs literature.
1 Introduction
Consider the following set of $K$ variables that admits a factor structure for $i=1,\ldots,N$:
$\mathbf{Z}_{i}=\mathbf{F}\mathbf{C}_{i}+\mathbf{U}_{i},$ (1.1)
where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is the error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ is latent common factors that can be either stationary or not, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix. (1.1) nests several popular settings.
Example 1. Common Correlated Effects (CCE). Let $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables and $\mathbf{y}_{i}$ is the dependent variable. Here, $K=k+1$, and $i$ indexes individuals, such as countries or households. Then (1.1) represents a classical CCE setup of Pesaran, (2006):
$\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},,\quad\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i},$ (1.2)
where, $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter vector, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings). Both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, and this induces cross-section dependence among individuals and endogeneity as the error in $\mathbf{y}_{i}$ is correlated with $\mathbf{X}_{i}$. (1.1) nests (1.2) with $\mathbf{C}_{i}=[\boldsymbol{\Gamma}_{i}\boldsymbol{\beta}+\boldsymbol{\gamma}_{i},\boldsymbol{\Gamma}_{i}]$ and $\mathbf{U}_{i}=[\mathbf{V}_{i}\boldsymbol{\beta}+\boldsymbol{\varepsilon}_{i},\mathbf{V}_{i}]$.
Example 2. Distinct Correlated Effects (DCE). We can introduce distinct but correlated factors $\mathbf{F}=[\mathbf{F}_{\mathbf{x}},\mathbf{F}_{\mathbf{y}}]\in\mathbb{R}^{T\times(m_{x}+m_{y})}$ in Example 1, which makes the setup mode applicable:
$\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}_{\mathbf{y}}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},,\quad\mathbf{X}_{i}=\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$ (1.3)
as considered in Cui et al., (2022) or Stauskas and De Vos, (2025). Here, (1.1) nests (1.3) similarly to Example 1, but $\boldsymbol{\gamma}_{i}$ is modified accordingly, depending on whether $\mathbf{F}_{\mathbf{y}}\subset\mathbf{F}_{\mathbf{x}}$, vice versa or $\mathbf{F}_{\mathbf{x}}\cap\mathbf{F}_{\mathbf{y}}=\emptyset$.
Example 3. Block Structure (FRED-MD-QD , EA-MD-QD). The set of regressors in Examples 1 and 2 can be high-dimensional, but follow a natural group structure with $k$ blocks and $m$ underlying factors. Let $i_{j}$ index series belonging to a block $j$ for $j=1,\ldots,k$. Then $\mathbf{Z}_{i}$ is $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}=[\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{k}}]^{\prime}$. Here, $N$ is the number of series per block. Blocks, such as monetary, industrial or fiscal series can be found in FRED-MD (see McCracken and Ng,, 2016), FRED-QD (see McCracken and Ng,, 2020) or EA-MD-QD (see Barigozzi et al.,, 2024) datasets. These datasets have been very popular in factor-augmented forecasts (see e.g. Karabiyik et al.,, 2019, Stauskas and Westerlund,, 2022, or Morico and Stauskas,, 2025).
In all the cases above, we can estimate the factors by taking $K$ cross-section averages (CAs): $\widehat{\mathbf{F}}=\overline{\mathbf{Z}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g. Pesaran and Tosetti,, 2011). In Example 1, $\widehat{\mathbf{F}}$ is used as an additional regressor to merely control for the factors, and subsequently $\boldsymbol{\beta}$ is estimated with the least squares (LS), which constitutes the typical CCE estimator (see e.g. Karabiyik et al.,, 2017). In Example 2, only $\widehat{\mathbf{F}}_{\mathbf{x}}=\overline{\mathbf{X}}$ is used as a regressor, while $\mathbf{F}_{\mathbf{y}}$ is controlled by bootstrap (see Stauskas and De Vos,, 2025). In Example 3, coefficient of $\widehat{\mathbf{F}}$ itself matters for forecast evaluation (see e.g. Massacci and Kapetanios,, 2024).
A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties. Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings (see e.g. Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023). This property holds as long as the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs ($K\in\{k,k+1\}$ in Examples 1-3). However, there are reasons to ensure that $m=K$. If $m<K$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017). This situation is common in macro datasets (see De Vos and Stauskas,, 2024). In DCE, $\mathbf{F}_{\mathbf{x}}$ estimated with an excess error makes bootstrap inconsistent (see Stauskas and De Vos,, 2025). In Example 3, popular tests of forecast accuracy comparison have highly non-standard asymptotic distributions (see Stauskas and Westerlund,, 2022). More generally, inclusion of too many CAs might even lead to a model that over-fits the data. In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis et al.,, 2021; Juodis,, 2022, for a discussion). In response to this, Margaritella and Westerlund, (2023) were the first to demonstrate that Information Criteria (IC) of Bai and Ng, (2002) from Principal Components (PC) literature can be consistently applied in the pure CCE setting in Example 1. Subsequently, De Vos and Stauskas, (2024) showed similar results applicable to more general Examples 2 and 3.111Juodis, (2022) considers eigenvalue-based selection (see also Ahn and Horenstein,, 2013). In that study, factors are stationary, as well. Another set of results concern fixed-$T$, where properties of $\mathbf{F}$ do not play an asymptotic role. Also, the method focuses solely on the number of factors, but cannot select the CAs. We leave analysis of this approach for the future research.
In this study, we re-visit IC proposed in Margaritella and Westerlund, (2023) and De Vos and Stauskas, (2024). Given the power of CAs to proxy very general factor structures, it is only natural to evaluate IC under nonstationarity conditions. Contrary to the sentiment in the CCE/CAs literature, IC crumble under very mild forms of nostationarity. We demonstrate formally and in simulations that they exhibit a severe under-selection problem, which is independent of penalties used in the objective functions of IC. Moreover, nonstationarity of $\mathbf{F}$ cannot be accommodated in the objective functions a priori, as their very design is based on the idea that properties of factors and idiosyncratics should be similar. This study can be seen as a caution to practitioners to transform data and not to extrapolate attractive properties of CAs should they need to choose a number thereof.
2 Econometric Setup
To further operationalize our examples, we introduce $M$ which is a set of column indices of $\overline{\mathbf{Z}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{K\times g}$ picks the corresponding $g=\mathrm{cols}(\mathbf{q}_{M})$ averages in practice. For instance, $\overline{\mathbf{Z}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k+1$ CAs in the typical CCE setup. Moreover, let $\mathbf{q}_{M}=[\boldsymbol{0}_{g\times 1},\mathbf{q}_{M_{\mathbf{x}}}^{\prime}]^{\prime}$, then $\overline{\mathbf{Z}}\mathbf{q}_{M}=\overline{\mathbf{X}}\mathbf{q}_{M_{\mathbf{x}}}=\widehat{\mathbf{F}}_{M_{\mathbf{x}}}$ selects only from the set of $k$ regressors (or blocks), as required in Examples 2 and 3.222Throughout the remainder of the study, we use $m$ to denote the number of factors without the loss of generality. Clearly, this number can vary among our examples. Let accordingly $M_{0}$ ($M_{\mathbf{x},0}$) denote the set of averages from $\overline{\mathbf{Z}}$ ($\overline{\mathbf{X}}$) such that $\mathrm{rk}(\overline{\mathbf{C}}\mathbf{q}_{M_{0}})=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x},0}})=m$, and $|M_{0}|=|M_{\mathbf{x},0}|=m$, where $|M|$ denotes cardinality of an arbitrary set $M$. IC under consideration are given by
$\mathrm{IC}^{MW}(M)=\ln\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}\right)+g\cdot p_{N,T},$ (2.1)
$\mathrm{IC}^{DVS}(M_{\mathbf{x}})=\ln\det\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}\right)+g\cdot k\cdot p_{N,T},$ (2.2)
$p_{N,T}=\begin{cases}\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\\ \frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right).\end{cases}$ (2.3)
Here and further, MW represents Margaritella and Westerlund, (2023), while DVS stands for De Vos and Stauskas, (2024). $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix onto the orthogonal complement of $\mathbf{A}$, $\mathbf{A}^{+}$ is a Moore-Penrose (MP) inverse, while $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ with $\widehat{\boldsymbol{\beta}}$ is obtained with the CCE estimator under all available $k+1$ CAs. Finally, $p_{N,T}$ is a penalty term which goes to 0 with $(N,T)\to\infty$. Examples of feasible and most popular penalties are given in (2.3) for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002).
Let $\overline{M}$ ($\overline{M}_{\mathbf{x}}$) be the set of indices of all available CAs that $|\overline{M}|=k+1$ ($|\overline{M}_{\mathbf{x}}|=k$). Then
$\widehat{M}\hskip 2.84526pt(\widehat{M}_{\mathbf{x}})=\operatorname*{arg\,min}_{M\subseteq\overline{M}\hskip 2.84526pt(M_{\mathbf{x}}\subseteq\overline{M}_{\mathbf{x}})}\mathrm{IC}(M)\hskip 2.84526pt(\mathrm{IC}(M_{\mathbf{x}})),$ (2.4)
where $|\widehat{M}|\hskip 2.84526pt(|\widehat{M}_{\mathbf{x}}|)=g$ provides the estimator of the number of factors for both criteria. In MW and DVS, it is demonstrated that $\mathbb{P}(\widehat{M}=M_{0})\to 1$ and $\mathbb{P}(\widehat{M}_{\mathbf{x}}=M_{\mathbf{x},0})\to 1$, respectively, as $(N,T)\to\infty$. These results theoretically justify the use of IC previously very popular in the PC literature to select the number of factors. These findings are based on stationary $\mathbf{F}$, which has been considerably relaxed in the CCE/CAs literature. Therefore, our natural goal is to examine (2.1) and (2.2) under departures from this restrictive assumption. Throughout our analysis, we employ the following set of assumptions.
Assumption 1.
  • (a)
    $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009). In particular,
    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}+\frac{\mathbf{Q}}{T^{\tau}},\quad\mathbf{Q}<0\hskip 5.69054pt(diagonal),$
    where $\mathbf{u}_{f,t}$ is a zero-mean linear process.
  • (b)
    If $\tau=0$, then $|\lambda_{\mathrm{min}}(\mathbf{Q})|<2$, where $\lambda_{\mathrm{min}}$ is the minimum eigenvalue, in order to ensure that $\mathbf{R}_{f,T}=\mathbf{I}_{m}+\mathbf{Q}$ is inside the unit circle, and $\mathbf{u}_{f,t}$ is such that $\{\mathbf{f}_{t}\}$ has absolute summable autocovariances.
Assumption 2. Let $\mathbf{e}_{i,t}=(\varepsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}\in\mathbb{R}^{k+1}$. Then
  • (a)
    • (i)
      If $\tau\in(0,1)$, then $\{\mathbf{e}_{i,t}\}$ is uncorrelated over time with $\mathbb{E}(\mathbf{e}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\mathbf{e}_{i,t}\mathbf{e}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{ee,i,t}$ positive definite and $\mathbb{E}(\|\mathbf{e}_{i,t}\|^{4})<\infty$.
    • (ii)
      If $\tau=0$, then we let $\mathbf{e}_{i,t}=\mathbf{K}_{i}(L)\boldsymbol{\epsilon}_{i,t}=\sum_{j=0}^{\infty}\mathbf{K}_{i,j}\boldsymbol{\epsilon}_{i,t-j}$, where $\boldsymbol{\epsilon}_{i,t}$ is independent across $t$ with $\mathbb{E}(\boldsymbol{\epsilon}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\boldsymbol{\epsilon}_{i,t}\boldsymbol{\epsilon}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{\epsilon\epsilon,i,t}$ positive definite, $\mathbb{E}(\|\boldsymbol{\epsilon}_{i,t}\|^{4})<\infty$, and $\sum_{j=0}^{\infty}j^{1/2}\|\mathbf{K}_{i,j}\|<\infty$. Also, $\boldsymbol{\xi}_{t}=\mathrm{vec}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}-\mathbb{E}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right)\right)$ is strong mixing with coefficients of size $-bd(b-d)$ with $b>4$ and $b>d>2$, $\mathbb{E}(\|\boldsymbol{\xi}_{t}\|^{b})<\infty$, and $\lim_{T\to\infty}T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{s}^{\prime})$ is positive definite.
    • (iii)
      In both cases, $\varepsilon_{i,t}$ and $\mathbf{v}_{j,s}$ are independent for all $i,j,t,s$.
  • (b)
    We have $\left\|\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\mathbf{e}_{i,t}\mathbf{e}_{j,t}^{\prime}-\boldsymbol{\Sigma}_{ee}\right\|=o_{p}(1)$ $N\to\infty$ for a positive definite matrix $\boldsymbol{\Sigma}_{ee}$.
Assumption 3. $\mathbf{f}_{t}$ and $\mathbf{e}_{i,s}$ are independent for all $t$, $s$ and $i$.
Assumption 4. $\mathbf{C}_{i}$ is a matrix, such that $\mathbf{C}_{i}=\mathbf{C}+\boldsymbol{\eta}_{i},\mathrm{vec}(\boldsymbol{\eta}_{i})\sim IID(\boldsymbol{0}_{Km\times 1},\boldsymbol{\Omega})$ for positive semi-definite $\boldsymbol{\Omega}$. Also, $\overline{\mathbf{C}}=[\overline{\mathbf{C}}_{m},\overline{\mathbf{C}}_{-m}]$, where $\overline{\mathbf{C}}_{-m}\in\mathbb{R}^{m\times(K-m)}$ and $\overline{\mathbf{C}}_{m}\in\mathbb{R}^{m\times m}$ is full rank for all $N$, including $N\to\infty$, and $\|\mathbf{C}_{i}\|<\infty$. If $m=K$, then $\overline{\mathbf{C}}=\overline{\mathbf{C}}_{m}$.
Assumption 1 (a) parameterizes factors very flexibly and allows to perform comparative statics, since $\tau\to 1$ implies that the process becomes more persistent and approaches the usual local-to-unity process. We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allows us vary $\tau$ and document its effect on IC performance. This assumption also implies that $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$ which is a constant positive definite matrix. Its part (b) ensures that we return to the usual stationarity conditions under $\tau=0$. Assumption 2 (a) is split into two parts. If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}$ converges to a zero matrix too slow for our purposes, which would dominate and obscure the very effect of factor persistence on IC, and we would like to avoid this (see also Lemma 3.1 in Magdalinos and Phillips,, 2009). For completeness, we present simulation evidence under serial correlation, as well, which does not demonstrate a major negative effect on IC. The rest of the assumptions are standard, as they ensure mutual independence of factors and the error terms (Assumption 3), weak cross-section dependence of idiosyncratics (Assumption 2 (b)) and informativeness of the factor loadings (rank condition in Assumption 4). The loadings are random as in e.g. Pesaran, (2006), but they can be treated as deterministic parameters, as well.
To formulate our key theoretical result, let us denote the objective function component in (2.1) and (2.2) evaluated at some set of CAs $M$ by $\mathcal{V}_{j}$ for $j\in\{MW,DVS\}$. Similarly, let $\mathcal{V}_{j}^{0}$ be the same component evaluated at $M_{0}$. Proposition 1 below characterizes the asymptotic behavior of $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}$ under $M_{0}\subset M$ (over-specification) and $M\subset M_{0}$ (under-specification).333 Strictly speaking, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$. We only analyze the case of $M\subset M_{0}$ for brevity. As Margaritella and Westerlund, (2023) pointed out analysis would lead to the same conclusion in all the cases.
Proposition 1. Under Assumptions 1-4 as $(N,T)\to\infty$, we have that
$\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=\begin{cases}O_{p}(T^{\tau})\hskip 5.69054pt\textit{under}\hskip 5.69054ptM\subset M_{0},\\ O_{p}(C_{N,T}^{-2})\hskip 5.69054pt\textit{under}\hskip 5.69054ptM_{0}\subset M.\end{cases}$ (2.5)
Let us begin from the over-specification. In this case, the CAs are able to approximate all $m$ factors under Assumption 4. Also, for illustration purposes let $j=DVS$. Then $\mathcal{V}_{DVS}$ (and $\mathcal{V}_{DVS}^{0})$ is exactly (log-determinant of) the denominator of the CCE estimator, which is consistent for a positive definite matrix under general factors (see Westerlund,, 2018), and a similar argument applies to the objective function of MW. Thus, it is natural that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}$ is negligible. Moreover, the rate is already detected in Bai and Ng, (2002) under stationary $\mathbf{F}$ for PC, and it remains identical here. This rate is instructive and determines the properties of penalty $p_{N,T}$, which is the key motivation behind the functional forms in (2.3). Indeed, if $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$ as $(N,T)\to\infty$, Proposition 1 is the key ingredient to show that $\mathbb{P}(\mathrm{IC}^{DVS}(M_{\mathbf{x}})-\mathrm{IC}^{DVS}(M_{\mathbf{x},0})>0)\to 1$, and the same statement holds for MW. This means that the probability that IC are minimized at some other set containing $M_{0}$ goes to 0. In general, under $M_{0}\subset M$, the usual CCE logic prevails deceptively.
Under $M\subset M_{0}$, Assumption 4 is not satisfied, and the rank condition does not hold. This means that $m-g$ factors are not asymptotically projected out. Given that $\mathcal{V}_{j}$ is a quadratic form for both $j\in\{MW,DVS\}$, their second empirical moment is of order $O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a), but the objective function offers normalization by $T$ only. Exactly this imbalance induces divergence, and this implies that the probability that IC are minimized at some other set contained in $M_{0}$ does not necessarily go to 0. Under stationary $\mathbf{F}$, however, it is clear that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=O_{p}(1)$ and, in fact, $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}\to_{p}c>0$ as $(N,T)\to\infty$ (see the analysis of MW and DVS). Our Proposition 1 naturally provides the same result under $\tau=0$. Note that the failure of both IC is purely driven by the objective function and is not related to the penalty $p_{N,T}$. In fact, the only theoretical prescription for $p_{N,T}$ depends only on $C_{N,T}^{2}$ and not $\tau$, as given in (2.5).
The discussion above seems to suggest that incorporating the degree of non-stationarity of $\mathbf{F}$ into the objective function should solve the issue and allow keeping $p_{N,T}$ the same. However, this is not the case. Let us assume that $\tau$ is known and normalize the objective function by $T^{1+\tau}$, instead, as suggested by Assumption 1 (a). In the Supplement, we show that
$\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}=\ln\mathrm{det}\left[\mathbf{I}_{k}+\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x}}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}})\mathbf{X}_{i}}_{O_{p}(1)}\underbrace{\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}}\mathbf{X}_{i}\right)^{-1}}_{O_{p}(T^{\tau})}\right],$ (2.6)
which means that, again, the determinant does not converge to a strictly positive constant. The reason behind this is that the rank condition is satisfied in the second component of the product, and by following steps of Karabiyik et al., (2017) and Westerlund, (2018), we obtain
$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}}\mathbf{X}_{i}=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}+o_{p}(1)=O_{p}(T^{-\tau}),$ (2.7)
since the idiosyncratics are stationary. In effect, the new normalization induces imbalance elsewhere. Therefore, IC fundamentally depend not only on stationarity of factors, but also on stationarity of idiosyncratics. We confirm this in the additional Monte Carlo simulations presented in the Supplement, where the adjustment of the objective function becomes more harmful with $\tau$.
Remark 1. Throughout, we assume that $M_{0}$ ($M_{0,\mathbf{x}}$) is unique. However, because CAs do not have a natural ordering, the problem is only set-identified. It can be shown that both IC select the set that minimizes the asymptotic mean squared error of all the sets that satisfy Assumption 4 (see e.g. Corollary 3.1 in Margaritella and Westerlund,, 2023). In this study, the first order issue is ability of IC to select any $\widehat{M}$ with the cardinality of $m$.
3 Simulation Study
The data generating process of the simulation follows Margaritella and Westerlund, (2023):
$\mathbf{y}_{i}$ $=\mathbf{X}_{i}\boldsymbol{\beta}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},\quad\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i},$ (3.1)
where $\mathbf{X}_{i}=\left[\mathbf{x}_{i,1},...,\mathbf{x}_{i,T}\right]^{\prime}$ is a $T\times k$ matrix of observable regressors, $\boldsymbol{\beta}=\iota^{\prime}0.5$ a $k\times 1$ vector of unobserved parameters set to 0.5. $\mathbf{F}$ is a $T\times m$ matrix of $m$ unobserved common factors. For simplicity, we assume that $\mathbf{X}_{i}$ and $\mathbf{u}_{i}$ load on $m$ factors. The factor loadings are generated as $\boldsymbol{\Gamma}_{i}=[\mathbf{I}_{m},\mathbf{0}_{m\times(k-m)}]\psi_{i}$, that is, a matrix $m\times k$, with $\psi_{i}\sim N(1,1)$ and the elements of the vector $1\times m$ $\boldsymbol{\gamma}_{i}$ are drawn from $N(1,1)$.
In the case of time- and space-independent errors, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$. In the case of errors correlated weakly across time and cross-sections we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
$\mathbf{e}_{i,t}=(\epsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}=\mathbf{P}_{\rho}\mathbf{e}_{i,t-1}+\mathbf{z}_{\mathbf{w},i,t},\quad\mathbf{P}_{\rho}=\begin{bmatrix}\rho&\mathbf{0}_{1\times k}\\ \mathbf{0}_{k\times 1}&\rho_{v}\mathbf{I}_{k}\end{bmatrix},\quad\mathbf{z}_{\mathbf{w},i,t}=\begin{bmatrix}\sqrt{\frac{m\left(1-\rho^{2}\right)}{1+2J\kappa^{2}}}\boldsymbol{\epsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa\mathbf{w}_{i}^{\prime}\right)\\ \sqrt{\frac{\left(1-\rho_{v}^{2}\right)}{1+2J_{v}\kappa_{v}^{2}}}\boldsymbol{\Upsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa_{v}\boldsymbol{\tilde{\mathbf{w}}}_{i}^{\prime}\right)\end{bmatrix}$ (3.2)
where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero. $\rho$ is set to 0.5, $\kappa=0.2$ and $J=5$. The idiosyncratics in $\mathbf{X}_{i}$, $\mathbf{V}_{i}$, are also drawn as potentially weakly correlated over time and cross-sections. Here, $\mathbf{\Upsilon}_{t}$ is a $N\times k$ matrix, which stacks $\boldsymbol{\upsilon}_{i,t}\sim N(\mathbf{0}_{k},\mathbf{I}_{k})$ over $i$, and it is independent of $\boldsymbol{\epsilon}_{t}$. The vector $\boldsymbol{\iota}_{N,i}\in\mathbb{R}^{N\times 1}$ contains 1 in the $i$-th coordinate, and zeros elsewhere. $\rho_{v}=0$ for the time IID case and $\rho_{v}=0.5$ for weakly time-correlated idiosyncratics. In both cases, we allow for weak correlation across cross sections, $\kappa_{v}=0.2$, $J_{v}=5$ and $\tilde{\mathbf{w}}_{i}$ is defined as $\mathbf{w}_{i}$ above. Therefore, we also test correlated idiosyncratics over time, whereas our theory prescribes uncorrelated innovations for $\tau\in(0,1)$. The factors are drawn as:
$\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{Q}T^{-\tau},\quad\mathbf{Q}=diag(q_{1},...,q_{m}),q_{j}\sim U(0,2)$ (3.3)
$\mathbf{u}_{t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.4)
In the stationary factors case, the factors are drawn as a stable VAR with $\tau=0$. Two cases for non-stationary factors are considered: the near stationary case of $\tau=0.1$ and the case close to local-to-unity process $\tau=0.9$. All simulations have 1000 repetitions. In the Supplement, we provide additional simulation evidence of the eigenvalue growth ratio (ER) based on Juodis, (2022) for comparison. The focus of this method is the number of factors ($m$), and it does not select the sufficient set of CAs. Nevertheless, its performance crumbles with $\tau\to 1$, as well. Moreover, we present extended tables with extra $N,T=150,200$ for all our experiments.
Table 3 below presents the average of the number of CAs selected for the case of stationary factors ($\tau=0$), the near stationary case and the close to local-to-unity process. The idiosynchratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units. We start with the in the literature established case of stationary factors. In the small $T$ cases, $T=10$ and $T=20$, we first observe an overselection of the number of CAs, followed by an underselection. This pattern resembles the results in Margaritella and Westerlund, (2023) and is consistent across all values for $N$ and $\tau$ [^4] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$ If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors. Once the number of cross-sectional units and time periods passes 50, all four criteria select close to or exactly the correct number of factors.
Moving to the near-stationary case, the pattern remains the same. However, once moving to the near unit root process, we observe for all four criteria, overselection of the number of CAs for the small time series case of $T=10$, followed by underselection if $T=20$. The underselection persists even if $N$ and $T$ increase to 100. This showcases the collapse of all information criteria in the case of non-stationary factors.
Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
$\tau=0$ $\tau=0.1$ $\tau=0.9$
$N$ $T$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$
10 10 9.00 8.58 5.12 1.00 8.99 8.65 4.92 1.00 8.99 8.51 2.11 1.02
20 10 8.99 8.78 7.98 3.18 8.98 8.78 7.94 3.19 8.95 8.31 7.12 1.52
50 10 8.99 8.99 7.94 6.83 9.00 8.98 7.94 6.68 8.96 8.87 5.97 2.47
100 10 9.00 9.00 7.93 7.52 9.00 9.00 7.97 7.69 8.98 8.94 5.42 3.42
10 20 1.15 1.02 1.00 1.00 1.19 1.03 1.01 1.00 1.21 1.12 1.05 1.01
20 20 1.23 1.01 1.93 1.02 1.27 1.02 1.97 1.04 1.33 1.16 1.37 1.13
50 20 2.68 2.04 2.02 1.24 2.72 2.10 2.08 1.33 1.45 1.35 1.32 1.20
100 20 3.79 3.57 3.29 2.67 3.58 3.40 2.98 2.45 1.51 1.46 1.35 1.30
10 50 1.21 1.13 1.05 1.03 1.11 1.05 1.00 1.00 1.25 1.21 1.14 1.09
20 50 1.09 1.03 1.21 1.01 1.08 1.02 1.29 1.01 1.29 1.23 1.32 1.20
50 50 3.60 2.80 3.91 2.45 3.90 3.24 3.97 2.94 1.95 1.68 1.92 1.55
100 50 3.33 3.21 3.30 2.95 3.19 3.06 3.08 2.71 1.59 1.51 1.59 1.48
10 100 1.00 1.00 1.26 1.18 1.00 1.00 1.28 1.22 1.15 1.14 1.04 1.03
20 100 1.66 1.36 2.50 2.15 1.76 1.43 2.61 2.41 1.43 1.36 1.66 1.58
50 100 4.00 3.99 4.00 4.00 4.00 3.98 4.00 3.99 2.11 1.97 1.97 1.80
100 100 4.01 4.00 4.00 4.00 4.01 4.00 4.00 4.00 2.25 2.04 2.17 1.86
Table 1: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.2) and (2.1). For $IC^{DVS}_{1}$ and $IC^{MW}_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $IC^{DVS}_{2}$ and $IC^{MW}_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ uncorrelated over time, but weakly correlated across units, see (3.2).
To investigate the behavior of the IC over different levels of non-stationarity, we increase $\tau$ in $0.05$ steps from 00 to $0.95$. Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$. We also allow the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ to be uncorrelated over time ($\rho=\rho_{v}=0$, in blue) and weakly correlated ($\rho=\rho_{v}=0.5$, in red). The share of misselected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint. Neither of IC is strongly affected by weak dependence over time in the idiosyncratics. However, as shown in the Supplement, if the idiosynchratics in $\mathbf{X}_{i}$ are non-stationary as well, the IC equally fail to identify the correct number of CAs, with and without the adjustment of $T^{1+\tau}$. This again stresses the requirement for balanced stationarity in $\mathbf{F}$ and $\mathbf{V}_{i}$.
(a) $IC^{DVS}_{1}$
(b) $IC^{MW}_{1}$
Figure 1: Share of incorrect selected cross-sectional averages with increasing $\tau$. $N=T=100$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023). Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time ($\rho=\rho_{v}=0$) and weakly correlated if $\rho=\rho_{v}=0.5$. Both are weakly correlated across units.
4 Conclusion
As predicted by the theory and our simulation results, the inability of IC to accommodate non-stationary factors is evident. The number of CAs is severely underselected, independent of the number of time periods and units. The underselection has severe consequences in terms of our examples. For the CCE estimator, since the number of factors is larger than the number of CAs, the rank condition fails and therefore consistency of the CCE estimator as well. In DCE case, the bootstrap procedure cannot help accommodate different factors in $\mathbf{X}$ and $\mathbf{y}$ anymore. In forecasting, incorrectly estimated factors harm the forecast comparison procedures.
5 Appendix
5.1 Auxiliary Results
Lemma 1
Under Assumptions 1 - 4, we have that
$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\overline{\mathbf{E}}\right\|=O_{p}(N^{-1/2}T^{-\tau/2})$
as $(N,T)\to\infty$
Proof. To begin with, we can write $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\overline{\mathbf{E}}=\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}$. Then for some positive $\epsilon$,
$\mathbb{P}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right\|>\epsilon\right)$ $\leq\epsilon^{-1}\mathbb{E}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right\|\right)$
$=\frac{1}{\epsilon\sqrt{N}}\mathbb{E}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}(\sqrt{N}\overline{\mathbf{e}}_{t}^{\prime})\right\|\right)$
$\leq\frac{1}{\epsilon\sqrt{N}}\mathbb{E}\left(\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\left\|\mathbf{f}_{t}\right\|\left\|(\sqrt{N}\overline{\mathbf{e}}_{t})\right\|\right)$
$=\frac{1}{\epsilon\sqrt{N}}\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}(\left\|\mathbf{f}_{t}\right\|)\mathbb{E}\left(\left\|(\sqrt{N}\overline{\mathbf{e}}_{t})\right\|\right)$
$\leq\frac{1}{\epsilon\sqrt{N}T^{\tau/2}}\left(\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}(\left\|\mathbf{f}_{t}\right\|)^{2}\right)^{1/2}\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left(\left\|(\sqrt{N}\overline{\mathbf{e}}_{t})\right\|\right)^{2}\right)^{1/2}}_{O(1)}$
$=O(N^{-1/2}T^{-\tau/2}),$ (A.1)
because
$\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}(\left\|\mathbf{f}_{t}\right\|)^{2}$ $\leq\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left[\left(\sqrt{(\mathrm{tr}(\mathbf{f}_{t}\mathbf{f}_{t}^{\prime}))}\right)^{2}\right]$
$=\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}(\mathrm{tr}[\mathbf{f}_{t}\mathbf{f}_{t}^{\prime}])=O(1)$ (A.2)
by the concavity of function appearing in Jensen’s inequality. Note that this result gives a different rate than for similar terms in, for example, Pitarakis, (2023), which is brought down exactly by time dependence in $\{\overline{\mathbf{e}}_{t}\}$. Also note that independence between the factors and idiosyncratics is not strictly necessary, and we use it for simplicity. Indeed,
$\mathbb{P}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right\|>\epsilon\right)$ $\leq\frac{1}{\epsilon\sqrt{N}}\mathbb{E}\left(\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\left\|\mathbf{f}_{t}\right\|\left\|(\sqrt{N}\overline{\mathbf{e}}_{t})\right\|\right)$
$\leq\frac{1}{\epsilon\sqrt{N}}\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\left\|\mathbf{f}_{t}\right\|^{2}\right)^{1/2}\mathbb{E}\left(\left\|\sqrt{N}\overline{\mathbf{e}}_{t}\right\|^{2}\right)^{1/2}$
$\leq\frac{\sup_{t}\mathbb{E}\left(\left\|\sqrt{N}\overline{\mathbf{e}}_{t}\right\|^{2}\right)^{1/2}}{\epsilon\sqrt{N}T^{\tau/2}}\sup_{t}\mathbb{E}\left(\left\|T^{-\tau/2}\mathbf{f}_{t}\right\|^{2}\right)^{1/2}$
$=O(N^{-1/2}T^{-\tau/2}),$ (A.3)
because $\sup_{t}\mathbb{E}\left(\left\|T^{-\tau/2}\mathbf{f}_{t}\right\|^{2}\right)^{1/2}=O(1)$ by Lemma 3.1 in Magdalinos and Phillips, (2009).
Corollary 1
Under Assumptions 1 - 4, but $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ are uncorrelated over time, we have that
$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\overline{\mathbf{E}}\right\|=O_{p}(N^{-1/2}T^{-(1+\tau)/2})$
as $(N,T)\to\infty$
Proof. By assumption, we then have that $\overline{\mathbf{e}}_{t}$ is uncorrelated over time. Then by Markov’s inequality we obtain
$\mathbb{P}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right\|>\epsilon\right)$ $\leq\epsilon^{-2}\mathbb{E}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right\|^{2}\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\mathbb{E}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\mathbf{f}_{s}^{\prime}\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{t}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{NT^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(N\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{t}\right)\right]\right)$
$\leq\epsilon^{-2}\frac{\sup_{t}\mathbb{E}\left(N\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{t}\right)}{NT^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\right]\right)$
$\leq O(N^{-1}T^{-1-\tau})\times\underbrace{\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\prime}\right)\right]}_{O(1)}$
$=O(N^{-1}T^{-1-\tau}),$ (A.4)
which implies that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\overline{\mathbf{E}}\right\|=O_{p}(N^{-1/2}T^{-(1+\tau)/2})$ if $\overline{\mathbf{e}}_{t}$ is uncorrelated over time. Note that
$\mathbb{E}\left(N\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{t}\right)=\mathbb{E}\left(\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{j,t}\right)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathrm{tr}\left[\mathbb{E}\left(\mathbf{e}_{i,t}\mathbf{e}_{j,t}^{\prime}\right)\right]=O(1),$ (A.5)
uniformly in $t$ by our assumptions.
Lemma 2
Under Assumptions 1 - 4, we have that
$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-\tau/2})$
as $T\to\infty$
Proof. Similarly to Lemma 1, we have that by Markov’s and Cauchy-Schwarz inequalities
$\mathbb{P}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\right\|>\epsilon\right)$ $\leq\epsilon^{-1}\mathbb{E}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\right\|\right)$
$=\frac{1}{\epsilon}\mathbb{E}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\right\|\right)$
$\leq\frac{1}{\epsilon}\mathbb{E}\left(\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\left\|\mathbf{f}_{t}\right\|\left\|\mathbf{e}_{i,t}\right\|\right)$
$=\frac{1}{\epsilon}\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}(\left\|\mathbf{f}_{t}\right\|)\mathbb{E}\left(\left\|\mathbf{e}_{i,t}\right\|\right)$
$\leq\frac{1}{\epsilon T^{\tau/2}}\left(\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}(\left\|\mathbf{f}_{t}\right\|)^{2}\right)^{1/2}\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left(\left\|\mathbf{e}_{i,t}\right\|\right)^{2}\right)^{1/2}}_{O(1)}$
$=O(T^{-\tau/2}),$ (A.6)
as expected.
Corollary 2
Under Assumptions 1 - 4, but $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ are uncorrelated over time, we have that
$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{U}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$
as $T\to\infty$
Proof. We use an approach similar to the one in Corollary 1. Then by Markov’s inequality we obtain
$\mathbb{P}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\right\|>\epsilon\right)$ $\leq\epsilon^{-2}\mathbb{E}\left(\left\|\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\right\|^{2}\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\mathbb{E}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\mathbf{f}_{s}^{\prime}\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,t}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,t}\right)\right]\right)$
$\leq\epsilon^{-2}\frac{\sup_{t}\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,t}\right)}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\right]\right)$
$\leq O(T^{-1-\tau})\times\underbrace{\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\prime}\right)\right]}_{O(1)}$
$=O(T^{-1-\tau}),$ (A.7)
which implies that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$.
5.2 Proof of Proposition 1
In the proof we focus on the IC of De Vos and Stauskas, (2024) ($j=DVS$). Therefore, we will have $K=k$ throughout. The argument for the IC of Margaritella and Westerlund, (2023) ($j=MW$) is very similar, and we only need to take into account that we are working with the usual CCE setup, where $\widehat{\boldsymbol{\beta}}=\boldsymbol{\beta}+o_{p}(1)$ under the full set of $k+1$ CAs. We will comment on this at the end of the proof. For notational simplicity, we will use $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ without the subscript $\mathbf{x}$. Also, because IC of DVS covers Examples 2 and 3 (see the main text), we denote factors as $\mathbf{F}_{\mathbf{x}}$.
For brevity, let $\overline{\mathbf{Q}}_{M}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}$, where $\widehat{\mathbf{F}}_{M_{\mathbf{x}}}=\overline{\mathbf{X}}\mathbf{q}_{M}=(\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}+\overline{\mathbf{V}})\mathbf{q}_{M}$ for the selector matrix $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$. Then, for $M_{0}$ being the true subset of CAs such that $g=m$, let us characterize difference of two Information Criteria:
$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\ln\mathrm{det}\left(\overline{\mathbf{Q}}_{M}\right)-\ln\mathrm{det}\left(\overline{\mathbf{Q}}_{M_{0}}\right)+k(g-m)p_{N,T}$
$=\ln\left[\frac{\mathrm{det}(\overline{\mathbf{Q}}_{M})}{\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}})}\right]+k(g-m)p_{N,T}$
$=\ln\left[\det(\overline{\mathbf{Q}}_{M})\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}}^{-1})\right]+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\overline{\mathbf{Q}}_{M}\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+\overline{\mathbf{Q}}_{M}\overline{\mathbf{Q}}_{M_{0}}^{-1}-\mathbf{I}_{k}\right)+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+\overline{\mathbf{Q}}_{M}\overline{\mathbf{Q}}_{M_{0}}^{-1}-\overline{\mathbf{Q}}_{M_{0}}\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
$=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+k(g-m)p_{N,T},$ (A.8)
where we used $\ln(a)-\ln(b)=\ln\left(\frac{a}{b}\right)$ and $\mathrm{det}(\mathbf{A}\mathbf{B})=\mathrm{det}(\mathbf{A})\mathrm{det}(\mathbf{B})$. This implies that we will examine the asymptotic behavior of $\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$. Its behavior is determined by whether $M_{0}\subset M$ (over-specification) or $M\subset M_{0}$ (under-specification). In the latter case, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$. We only analyze the case of $M\subset M_{0}$ as in Margaritella and Westerlund, (2023) or De Vos and Stauskas, (2024), because other cases will lead to the same asymptotic conclusions.
5.3 Part $M\subset M_{0}$
By following De Vos and Stauskas, (2024), we have
$\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$ $=\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]+\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
$=\mathbf{I}-\mathbf{II}+\mathbf{III}.$ (A.9)
where, for instance, $\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}$ means that it is evaluated at $\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}$. We begin with
$\mathbf{I}$ $=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}$
$=\mathbf{I}_{a}+\mathbf{I}_{b}+\mathbf{I}_{c}+\mathbf{I}_{d}.$ (A.10)
where the expansion $\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}$ is the key. Note how
$T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}$ $=\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}+\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}+\mathbf{q}_{M}^{\prime}T^{-(1+\tau)}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}$
$+\mathbf{q}_{M}^{\prime}TT^{-(1+\tau)}T^{-1}\overline{\mathbf{V}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}$
$=\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}+O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(1+\tau)/2})$
$=\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}+O_{p}(\xi^{-1}_{N,T,\tau})$ (A.11)
by Corollary 2, where we defined $\xi_{N,T,\tau}^{-1}=N^{-1}T^{-\tau}+N^{-1/2}T^{-(1+\tau)/2}$. The remainder has a slightly different rate than in De Vos and Stauskas, (2024) (see e.g. 2.184 in their Online Supplement), because the idiosyncratic component is still stationary, and so
$\left\|T^{-(1+\tau)}\overline{\mathbf{V}}^{\prime}\overline{\mathbf{V}}\right\|=N^{-1}T^{-\tau}\left\|NT^{-1}\overline{\mathbf{V}}^{\prime}\overline{\mathbf{V}}\right\|=O_{p}(N^{-1}T^{-\tau}).$ (A.12)
Moreover, note that
$N\times\xi_{N,T,\tau}^{-1}$ $=N\times(O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(1+\tau)/2}))$
$=O_{p}(T^{-\tau})+O_{p}(\sqrt{N}T^{-1/2}T^{-\tau/2})=O_{p}(1)$ (A.13)
without further restrictions (but negligible under $TN^{-1}=O(1)$. Also,
$T^{\tau}\times\xi_{N,T,\tau}^{-1}$ $=T^{\tau}\times(O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(1+\tau)/2}))$
$=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})=o_{p}(1)$ (A.14)
without further restrictions as $\tau\in(0,1)$. Both of these results will be important when determining the orders of many terms.
Next, in case of $M\subset M_{0}$, we have that $\mathrm{rk}\left(\frac{1}{T^{1+\tau}}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)=m$ even asymptotically, and hence
$\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}=\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}+O_{p}(\xi_{N,T,\tau}^{-1}).$ (A.15)
This leads to the expansion
$\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}$ $=\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}+\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}$
$+\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}$
$+\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime},$ (A.16)
which now takes into account the mildly integrated factors. Then, we can continue with
$\left\|\mathbf{I}_{a}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
$\leq\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}}_{O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(2\tau+1)/2})}$
$+2\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(N^{-1}T^{-(1+\tau)/2})+O_{p}(N^{-1/2}T^{-(2+\tau)/2})}$
$+\underbrace{T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|^{2}}_{O_{p}(T^{-1}\xi_{N,T,\tau}^{-1})}$
$=O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(2\tau+1)/2}),$ (A.17)
which is the dominating order. Next, we have
$\left\|\mathbf{I}_{b}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$\leq\underbrace{N^{-1}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}((NT)^{-1})}$
$+\underbrace{2T^{(\tau-1)/2}N^{-1/2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{(\tau-1)/2}N^{-1/2})}$
$+\underbrace{T^{\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{x}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})}$
$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.18)
which is a slowly decaying term dominated by $T^{(\tau-1)/2}=o(1)$, since $\tau\in(0,1)$. We let $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$, and we move on to
$\left\|\mathbf{I}_{c}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$\leq\underbrace{T^{-(1+\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{o_{p}(C_{N,T}^{-2})}$
$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(C_{N,T}^{-2})}$
$+\underbrace{N^{-1/2}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1})}$
$+\underbrace{T^{(\tau-1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{T^{(\tau-1)/2}\xi_{N,T,\tau}^{-1})}$
×$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
$=O_{p}(C_{N,T}^{-2}),$ (A.19)
which is negligible. Ultimately,
$\left\|\mathbf{I}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
$\leq\underbrace{T^{-(1+\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{o_{p}(C_{N,T}^{-2})}$
$+\underbrace{N^{-1/2}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1})}$
$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(C_{N,T}^{-2})}$
$+\underbrace{T^{(\tau-1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(T^{\tau/2}C_{N,T^{\tau}}^{-2})}$
×$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
$=O_{p}(C_{N,T}^{-2}),$ (A.20)
which has an identical asymptotic behavior to the one of $\mathbf{I}_{c}$. Hence, overall,
$\left\|\mathbf{I}\right\|=\left\|\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right\|=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.21)
which is the slowest decaying term. Importantly, we also have that
$\left\|\mathbf{I}\mathbf{I}\right\|=\left\|\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right\|=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.22)
as in this case $m=g$, and therefore the asymptotic behavior of this term will be identical.
We are left to analyze $\mathbf{III}$. Because we are in the case of under-specification, we can decompose $\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}$ into $\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}=\mathbf{A}=[\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M},\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}]=[\mathbf{B},\mathbf{C}]$, where $\mathbf{q}_{M^{C}}$ is the complement selector matrix that is $k\times(m-g)$. By utilizing this representation, we can decompose the projection matrix $\mathbf{P}_{\mathbf{A}}=\mathbf{P}_{\mathbf{B}}+\mathbf{P}_{\mathbf{M}_{\mathbf{B}}\mathbf{C}}$. By following this, we obtain
$\mathbf{III}$ $=\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}+\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$=\mathbf{III}_{a}+\mathbf{III}_{b}+\mathbf{III}_{c}+\mathbf{III}_{d},$ (A.23)
where the term $\mathbf{III}_{d}$ is particularly problematic. By using random loadings assumption, it explicitly can be written as
$\mathbf{III}_{d}$ $=\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$=T^{-1}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}$
$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}$
$=\mathbf{III}^{1}_{d}+\mathbf{III}^{2}_{d}+\mathbf{III}_{d}^{3}+\mathbf{III}^{4}_{d}.$ (A.24)
Note that the key component in all $\mathbf{III}^{1}_{d}-\mathbf{III}^{4}_{d}$ is $\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$, where
$T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$ $=T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}-T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}$
$=T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}-T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}$
$\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}}-\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}},$ (A.25)
as $T\to\infty$, which is positive definite. Now, moving on to $\mathbf{III}^{1}_{d}$, we obtain
$\mathbf{III}^{1}_{d}$ $=T^{-1}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$
$=T^{\tau}\boldsymbol{\Gamma}^{\prime}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}$
$=O_{p}(T^{\tau}),$ (A.26)
which means that the whole term is divergent, and the same applies to $\mathbf{III}^{1}_{d}$. In particular, $\mathbf{III}^{4}_{d}$ has the same order, since
$\left\|\mathbf{III}^{4}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|$
$\leq T^{\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|^{2}$
×$\times\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}T^{-(1+\tau)}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
$=O_{p}(T^{\tau}),$ (A.27)
while
$\left\|\mathbf{III}^{2}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|$
$\leq T^{\tau}\left\|\boldsymbol{\Gamma}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
×$\times\underbrace{\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|}_{O_{p}(N^{-1/2})}$
$=O_{p}(T^{\tau}N^{-1/2}),$ (A.28)
which is negligible only if $\tau\in(0,0.5)$ under $TN^{-1}=O(1)$ (to relate this term to (5.3) and the subsequent similar terms), and the same condition applies to $\mathbf{III}^{3}_{d}$:
$\left\|\mathbf{III}^{3}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}\right\|$
$\leq T^{\tau}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}\right\|$
×$\times\underbrace{\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|}_{O_{p}(N^{-1/2})}$
$=O_{p}(T^{\tau}N^{-1/2}),$ (A.29)
as welll. Next, we move on to $\mathbf{III}_{a}$, and for this we firstly analyze
$\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$ $\leq\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|$
$+\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
$\leq\underbrace{\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{O_{p}(1)}$
$+\underbrace{T^{-(1+\tau)/2}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|}_{O_{p}(T^{-(1+\tau)/2})}$
$=O_{p}(1).$ (A.30)
Next,
$\left\|\mathbf{III}_{a}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
$\leq T^{-1}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|$
$=O_{p}(T^{-1}),$ (A.31)
which would not be negligible under correlated idiosyncratics (see Lemma 2). The remaining two terms, are slowly vanishing. In particular,
$\left\|\mathbf{III}_{b}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
$\leq T^{(\tau-1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
×$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
$=O_{p}(T^{(\tau-1)/2}),$ (A.32)
and
$\left\|\mathbf{III}_{c}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
$\leq T^{(\tau-1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
×$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
$=O_{p}(T^{(\tau-1)/2}),$ (A.33)
which is just a transpose of $\mathbf{III}_{b}$. Note again, that both of the terms would slowly diverge under the conditions of Lemma 2. Therefore, by combining our interim results, we obtain
$\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$ $=\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]+\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})+O_{p}(T^{\tau})+O_{p}(T^{(\tau-1)/2})$
$=O_{p}(T^{\tau}),$ (A.34)
which is the dominant order. This implies that
$T^{-\tau}(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}})$ $=T^{-\tau}\mathbf{III}_{d}+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-(\tau+1)/2})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-(\tau+1)/2}),$ (A.35)
which converges to a positive semi-definite matrix. However, note how by using the fact that $\overline{\mathbf{Q}}_{M_{0}}=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})$, we have that
$T^{-\tau}\overline{\mathbf{Q}}_{M_{0}}$ $=T^{-\tau}\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=T^{-\tau}\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=\frac{1}{T^{1+\tau}}\frac{1}{N}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{X}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=\frac{1}{T^{1+\tau}}\frac{1}{N}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{V}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}-T^{-1}\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=O_{p}(T^{-\tau})+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-1})$
$=o_{p}(1),$ (A.36)
because the idiosyncratics are stationary. Combining (5.3) and (5.3), we get
$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
$=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right](T^{-\tau}\overline{\mathbf{Q}}_{M_{0}})^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+o(1)$
↛$\not\to_{p}c>0,$ (A.37)
as in total $T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right](T^{-\tau}\overline{\mathbf{Q}}_{M_{0}})^{-1}$ does not converge to a positive semi-definite matrix. Therefore
$\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0\right)\not\to 0,$ (A.38)
and the under-specification risk is not eliminated.
5.4 Part $M_{0}\subset M$
We now move to the case when $M_{0}\subset M$ (over-specification). To analyze this case we will introduce additional notation. We decompose the selector matrix $\mathbf{q}_{M}=[\mathbf{q}_{M_{0}},\mathbf{q}_{M_{0}^{C}}]$, where $\mathbf{q}_{M_{0}}\in\mathbb{R}^{k\times m}$ and $\mathbf{q}_{M_{0}^{C}}\in\mathbb{R}^{k\times(g-m)}$, where the latter selector corresponds to $g-m$ excess (over-selected) averages. By using this, we can further decompose
$\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}=\begin{bmatrix}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}},\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}^{C}}\end{bmatrix}=\begin{bmatrix}\overline{\boldsymbol{\Gamma}}_{M_{0}},\overline{\boldsymbol{\Gamma}}_{M_{0}^{C}}\end{bmatrix}=\overline{\boldsymbol{\Gamma}}_{M}$ (A.39)
for short-hard notation. Then, we have
$\widehat{\mathbf{F}}_{M_{\mathbf{x}}}=\overline{\mathbf{X}}\mathbf{q}_{M}=\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}+\overline{\mathbf{V}}\mathbf{q}_{M}=\mathbf{F}_{\mathbf{x}}\begin{bmatrix}\overline{\boldsymbol{\Gamma}}_{M_{0}},\overline{\boldsymbol{\Gamma}}_{M_{0}^{C}}\end{bmatrix}+\begin{bmatrix}\overline{\mathbf{V}}_{M_{0}},\overline{\mathbf{V}}_{M_{0}^{C}}\end{bmatrix},$ (A.40)
where $\overline{\mathbf{V}}\mathbf{q}_{M}=\overline{\mathbf{V}}_{M}$ is partitioned accordingly. Similarly to the case of using the total available set of CAs ($k$), we introduce in spirit of Karabiyik et al., (2017) the following rotation matrix
$\overline{\mathbf{H}}_{M}=\begin{bmatrix}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}&-\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\overline{\boldsymbol{\Gamma}}_{M_{0}^{C}}\\ \mathbf{0}_{(g-m)\times m}&\mathbf{I}_{g-m}\end{bmatrix}=\begin{bmatrix}\overline{\mathbf{H}}_{M_{0}},\overline{\mathbf{H}}_{M_{0}^{C}}\end{bmatrix},$ (A.41)
with the obvious definitions of $\overline{\mathbf{H}}_{M_{0}}$ and $\overline{\mathbf{H}}_{M_{0}^{C}}$ and $\mathrm{rk}(\overline{\mathbf{H}}_{M})=g$. In what follows, the post-multiplication by this matrix leads to
$\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\overline{\mathbf{H}}_{M}=\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M}+\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M}=\begin{bmatrix}\mathbf{F}_{\mathbf{x}},\mathbf{0}_{T\times(g-m)}\end{bmatrix}+\begin{bmatrix}\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}},\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}^{C}}\end{bmatrix}.$ (A.42)
Finally, we introduce $\mathbf{D}_{M}=\mathrm{diag}(\mathbf{I}_{m},\sqrt{N}\mathbf{I}_{g-m})$, such that
$\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}=\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M}\mathbf{D}_{M}+\overline{\mathbf{V}}\mathbf{q}_{M}\mathbf{D}_{M}$ $=\begin{bmatrix}\mathbf{F}_{\mathbf{x}},\mathbf{0}_{T\times(g-m)}\end{bmatrix}+\begin{bmatrix}\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}},\sqrt{N}\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}^{C}}\end{bmatrix}$
$=\begin{bmatrix}\mathbf{F}_{\mathbf{x}},\mathbf{0}_{T\times(g-m)}\end{bmatrix}+\begin{bmatrix}\overline{\mathbf{V}}_{M_{1}}^{0},\overline{\mathbf{V}}_{M_{2}}^{0}\end{bmatrix}$
$=\mathbf{F}_{\mathbf{x}}^{0}+\begin{bmatrix}\overline{\mathbf{V}}_{M_{1}}^{0},\overline{\mathbf{V}}_{M_{2}}^{0}\end{bmatrix}$
$=\mathbf{F}_{\mathbf{x}}^{0}+\overline{\mathbf{V}}_{M}^{0},$ (A.43)
where $\left\|\overline{\mathbf{V}}_{M_{1}}^{0}\right\|=O_{p}(N^{-1/2})$, but $\left\|\overline{\mathbf{V}}_{M_{2}}^{0}\right\|=O_{p}(1)$, such that the last $g-m$ columns are non-degenerate.
Further, let us introduce $\mathbf{G}_{T}=\mathrm{diag}(T^{(1+\tau)/2}\mathbf{I}_{m},T^{1/2}\mathbf{I}_{g-m})$. Then, we work out the asymptotic limit of $\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}$. Particularly,
$\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}$ $=\mathbf{G}_{T}^{-1}\mathbf{F}_{\mathbf{x}}^{0\prime}\mathbf{F}_{\mathbf{x}}^{0}\mathbf{G}_{T}^{-1}+\mathbf{G}_{T}^{-1}\mathbf{F}_{\mathbf{x}}^{0\prime}\overline{\mathbf{V}}_{M}^{0}\mathbf{G}_{T}^{-1}+\mathbf{G}_{T}^{-1}\overline{\mathbf{V}}_{M}^{0\prime}\mathbf{F}_{\mathbf{x}}^{0}\mathbf{G}_{T}^{-1}+\mathbf{G}_{T}^{-1}\overline{\mathbf{V}}_{M}^{0\prime}\overline{\mathbf{V}}_{M}^{0}\mathbf{G}_{T}^{-1}$
$=\mathbf{S}+O_{p}(N^{-1/2}T^{-\tau/2})+O_{p}(T^{-1/2}),$ (A.44)
where $\mathbf{S}=\mathrm{diag}(T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}},T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})\in\mathbb{R}^{g\times g}$. The rate comes from the leading, but vanishing components of the last three terms on the right-hand side of (5.4). Specifically,
$\left\|T^{-1-\tau/2}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\leq T^{-\tau/2}N^{-1/2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\mathbf{H}}_{M_{0}}\right\|\left\|\overline{\mathbf{H}}_{M_{0}^{C}}\right\|\left\|NT^{-1}\overline{\mathbf{V}}^{\prime}\overline{\mathbf{V}}\right\|=O_{p}(N^{-1/2}T^{-\tau/2}),$ (A.45)
$\left|T^{-1-\tau/2}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}_{M_{2}}^{0}\right\|\leq\sqrt{N}T^{-1/2}\underbrace{\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\right\|}_{O_{p}(N^{-1/2})}\left\|\mathbf{q}_{M}\right\|\left\|\overline{\mathbf{H}}_{M_{0}^{C}}\right\|=O_{p}(T^{-1/2}),$ (A.46)
$\left\|T^{-(1+\tau)}\overline{\mathbf{V}}^{0\prime}_{M_{1}}\overline{\mathbf{V}}^{0}_{M_{1}}\right\|\leq N^{-1}T^{-\tau}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\mathbf{H}}_{M_{0}}\right\|^{2}\left\|NT^{-1}\overline{\mathbf{V}}^{\prime}\overline{\mathbf{V}}\right\|=O_{p}(N^{-1}T^{-\tau}),$ (A.47)
$\left|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\leq N^{-1/2}T^{-(1+\tau)/2}\underbrace{\left\|\sqrt{N}T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\right\|}_{O_{p}(1)}\left\|\mathbf{q}_{M}\right\|\left\|\overline{\mathbf{H}}_{M_{0}^{C}}\right\|=O_{p}(N^{-1/2}T^{-(1+\tau)/2})$ (A.48)
by Corollary (1). This implies that
$\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|=O_{p}(N^{-1/2}T^{-\tau/2})+O_{p}(T^{-1/2}),$ (A.49)
so the rate result applies to the MP inverses, as well.
In what follows, we will use a slightly different path than De Vos and Stauskas, (2024) in order to sharpen some rates. Let us use the fact that $\mathbf{q}_{M_{0}}$ gives the minimal set of averages that asymptotically span the space of $\mathbf{F}_{\mathbf{x}}$. This means that $\mathrm{rk}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}_{M_{0}})=m$ and this matrix is $m\times m$. This means that
$\widehat{\mathbf{F}}_{M_{0}}=\overline{\mathbf{X}}\mathbf{q}_{M_{0}}=\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}_{M_{0}}+\overline{\mathbf{V}}_{M_{0}}\Longleftrightarrow\mathbf{F}_{\mathbf{x}}=(\widehat{\mathbf{F}}_{M_{0}}-\overline{\mathbf{V}}_{M_{0}})\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}.$ (A.50)
Next, observe that
$\overline{\mathbf{Q}}_{M_{0}}$ $=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\mathbf{X}_{i}=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i})^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}(\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i})$
$=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i}),$ (A.51)
where we used the fact that $\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\mathbf{F}}_{M_{0}}=\mathbf{0}_{T\times m}$. The trick is to recognize that since $M_{0}\subset M$, we also have that $\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\widehat{\mathbf{F}}_{M_{0}}=\mathbf{0}_{T\times m}$, because, based on (A.40) and the block-wise formula for a projection matrix, we have
$\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\widehat{\mathbf{F}}_{M_{0}}$ $=\left(\mathbf{I}_{T}-\mathbf{P}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}-\mathbf{P}_{\mathbf{M}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}\overline{\mathbf{X}}\mathbf{q}_{M_{0}^{C}}}\right)\widehat{\mathbf{F}}_{M_{0}}$
$=\left(\mathbf{M}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}-\mathbf{P}_{\mathbf{M}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}\overline{\mathbf{X}}\mathbf{q}_{M_{0}^{C}}}\right)\overline{\mathbf{X}}\mathbf{q}_{M_{0}}$
$=\mathbf{M}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}\overline{\mathbf{X}}\mathbf{q}_{M_{0}}-\mathbf{M}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}\overline{\mathbf{X}}\mathbf{q}_{M_{0}^{C}}(\mathbf{q}_{M_{0}^{C}}^{\prime}\overline{\mathbf{X}}^{\prime}\mathbf{M}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}\overline{\mathbf{X}}\mathbf{q}_{M_{0}^{C}})^{+}\mathbf{q}_{M_{0}^{C}}^{\prime}\overline{\mathbf{X}}^{\prime}\mathbf{M}_{\overline{\mathbf{X}}\mathbf{q}_{M_{0}}}\overline{\mathbf{X}}\mathbf{q}_{M_{0}}$
$=\mathbf{0}_{T\times m}$ (A.52)
and so
$\overline{\mathbf{Q}}_{M}$ $=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i})^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}(\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i})$
$=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i}).$ (A.53)
This implies that by appropriately adding and subtracting, we can write
$\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$ $=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})$
$=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})$
$-\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1})$
$=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})$
$-\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})(\mathbf{V}_{i}-\overline{\mathbf{V}}_{M_{0}}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i})$
$=\mathbf{I}-\mathbf{II},$ (A.54)
where in the second-to-last equality we used the MP inverse properties:
$(\mathbf{F}_{\mathbf{x}}^{0\prime}\mathbf{F}_{\mathbf{x}}^{0})^{+}=\left[\begin{array}[]{cc}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}&\mathbf{0}_{m\times(g-m)}\\ \mathbf{0}_{(g-m)\times m}&\mathbf{0}_{(g-m)}\end{array}\right]^{+}=\left[\begin{array}[]{cc}(\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}&\mathbf{0}_{m\times(g-m)}\\ \mathbf{0}_{(g-m)\times m}&\mathbf{0}_{(g-m)}\end{array}\right],$ (A.59)
leading to
$\mathbf{P}_{\mathbf{F}^{0}_{\mathbf{x}}}$ $=\mathbf{F}_{\mathbf{x}}^{0}(\mathbf{F}_{\mathbf{x}}^{0\prime}\mathbf{F}_{\mathbf{x}}^{0})^{+}\mathbf{F}_{\mathbf{x}}^{0\prime}=\left[\begin{array}[]{cc}\mathbf{F}_{\mathbf{x}},\mathbf{0}_{T\times(g-m)}\end{array}\right]\left[\begin{array}[]{cc}(\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}&\mathbf{0}_{m\times(g-m)}\\ \mathbf{0}_{(g-m)\times m}&\mathbf{0}_{(g-m)}\end{array}\right]\left[\begin{array}[]{cc}\mathbf{F}_{\mathbf{x}}^{\prime}\\ \mathbf{0}_{(g-m)\times T}\end{array}\right]$ (A.65)
$=\mathbf{F}_{\mathbf{x}}(\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\mathbf{F}_{\mathbf{x}}^{\prime}=\mathbf{P}_{\mathbf{F}_{\mathbf{x}}},$ (A.66)
where $(\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}=(\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{-1}$, which is bounded in probability for finite $T$. Hence, we now need to obtain the order of $\mathbf{II}$, because it dominates $\mathbf{I}$ due to the case of $m<g$. To achieve this, we start with the decomposition
$\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}$ $=T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0}(T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})^{+}\overline{\mathbf{V}}_{M_{2}}^{0\prime}+T^{-(1+\tau)}\overline{\mathbf{V}}_{M_{1}}^{0}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\overline{\mathbf{V}}_{M_{1}}^{0\prime}$
$+T^{-(1+\tau)}\overline{\mathbf{V}}_{M_{1}}^{0}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\mathbf{F}_{\mathbf{x}}^{\prime}+T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\overline{\mathbf{V}}_{M_{1}}^{0\prime}$
$+\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}((\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+})\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}.$ (A.67)
Hence, if we insert (5.4) into $\mathbf{II}$, we obtain $\mathbf{II}=\mathbf{II}^{1}-\mathbf{II}^{2}-\mathbf{II}^{3}+\mathbf{II}^{4}$, and so
$\left\|\mathbf{II}^{1}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})\mathbf{V}_{i}\right\|$
$\leq\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}_{M_{2}}^{0}\right\|^{2}\left\|(T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})^{+}\right\|+T^{-\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|$
$+2\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
$+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1/2}\mathbf{V}_{i}^{\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|^{2}\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|}_{=o_{p}(C_{N,T})^{-2}}$
$=O_{p}(N^{-1})+O_{p}(T^{-1})$
$=O_{p}(C_{N,T}^{-2}),$ (A.68)
which is driven by the first term, because $\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}_{M_{2}}^{0}\right\|=O_{p}(N^{-1/2})+O_{p}(T^{-1/2})$, whereas
$\left\|T^{-1/2}\mathbf{V}_{i}^{\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|$ $\leq\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}^{0}_{M}\right\|\left\|\sqrt{T}\mathbf{G}_{T}^{-1}\right\|+T^{-1/2}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|$
$=O_{p}(N^{-1/2})+O_{p}(T^{-1/2})=O_{p}(C_{N,T}^{-1}),$ (A.69)
and so the last term in (5.4) is $o_{p}(C_{N,T}^{-2})$. The other terms are clearly dominated. Next, we go to
$\left\|\mathbf{II}^{2}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\overline{\mathbf{V}}_{M_{0}}^{0\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})\mathbf{V}_{i}\right\|$
$\leq\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|\left\|T^{-1}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0}\right\|\left\|(T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})^{+}\right\|\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}_{M_{2}}^{0}\right\|}_{=O_{p}(C_{N,T}^{-2})}$
$+T^{-\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|\left\|T^{-1}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|$
$+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|\left\|T^{-1}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{=O_{p}(N^{-1}T^{-(1+\tau)/2})}$
$+\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|\left\|T^{-(1+\tau)}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|\left\|T^{-1}\overline{\mathbf{V}}_{M_{1}}^{0\prime}\mathbf{V}_{i}\right\|$
$+\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|\left\|T^{-1/2}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right|\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|\left\|T^{-1/2}\mathbf{V}_{i}^{\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|$
$=O_{p}(C_{N,T}^{-2}),$ (A.70)
since
$\left\|T^{-1/2}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|$ $\leq(NT)^{-1/2}\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\mathbf{F}_{\mathbf{x}}\right\|+N^{-1/2}\left\|\sqrt{T}\mathbf{G}_{T}^{-1}\right\|\left\|T^{-1}\sqrt{N}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\overline{\mathbf{V}}_{M}^{0}\right\|$
$=O_{p}(N^{-1/2}),$ (A.71)
and so the overall order is dominated by the first and third terms in (5.4). Further, note that $\mathbf{II}^{3}$ is just a transpose of $\mathbf{II}^{2}$ and so
$\left\|\mathbf{II}^{3}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})\overline{\mathbf{V}}_{M_{0}}^{0}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i}\right\|=O_{p}(C_{N,T}^{-2}).$ (A.72)
Eventually we move on to
$\left\|\mathbf{II}^{4}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\overline{\mathbf{V}}_{M_{0}}^{0\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})\overline{\mathbf{V}}_{M_{0}}^{0}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i}\right\|$
$\underbrace{\leq\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0}\right\|^{2}\left\|(T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})^{+}\right\|}_{=O_{p}(N^{-1})}$
$+T^{-\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|$
$+2\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}_{M_{0}}^{0}\right\|$
$+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\right\|^{2}\left\|T^{-1/2}\overline{\mathbf{V}}_{M_{0}}^{0\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|^{2}\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|}_{O_{p}(N^{-3/2}T^{-\tau/2})+O_{p}(N^{-1}T^{-1/2})}$
$=O_{p}(N^{-1}),$ (A.73)
which is decided by the first term, which contributes to the order of $O_{p}(C_{N,T}^{-2})$ (note that provided that we assume $TN^{-1}=O(1)$ this term itself can also be seen as $O_{p}(C_{N,T}^{-2})$). Thus, overall we obtain
$\left\|\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right\|$ $=O_{p}(C_{N,T}^{-2}).$ (A.74)
Now, we put the results together in a fashion similar to De Vos and Stauskas, (2024). In particular, because of (A.50), we can easily demonstrate that $\left\|\overline{\mathbf{Q}}_{M_{0}}^{-1}\right\|=O_{p}(1)$ (see e.g. the analysis of CCEP denominator in e.g. Stauskas, (2023), where even more general factors are considered). Then
$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+O_{p}(C_{N,T}^{-2})\right)+k(g-m)p_{N,T}$
$=\ln\mathrm{det}(\mathbf{I}_{k})+O_{p}(C_{N,T}^{-2})+k(g-m)p_{N,T}$
$\to_{p}0,$ (A.75)
where the approximation comes from p. 119 in Paulsen, (1984):
$\ln\mathrm{det}(\mathbf{I}+O_{p}(n^{-1}))=O_{p}(n^{-1})$
for some integer $n$. However, given that $p_{N,T}O_{p}(C_{N,T}^{-2})\to\infty$, then we have that
$p_{N,T}^{-1}(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0}))=k(g-m)+o_{p}(1)$ (A.76)
and so
$\mathbb{P}\left(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0}))<0)\right)\to 0,$ (A.77)
and so the overspecification risk is eliminated.
5.5 Proof Changes for IC of Margaritella and Westerlund, (2023)
Note that IC of MW is based on the residuals from the first stage CCE regression under the full set of available $k+1$ CAs. In particular, let us assume homogeneous $\boldsymbol{\beta}$. Then
$\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}.$ (A.78)
We then obtain
$\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})$ $=\underbrace{\ln\left[1+\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}\right)\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}\right)^{-1}\right]}_{\mathcal{V}_{MW}-\mathcal{V}_{MW}^{0}}$
$+(g-m)\cdot p_{N,T},$ (A.79)
where, by inserting (A.78) we further analyze
$\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}$ $=\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
$-\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
$-\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
$+\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}).$ (A.80)
Clearly,
$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq(NT)^{-1}\underbrace{\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|}_{O_{p}(T^{\tau})}$
=$=\begin{cases}O_{p}(N^{-1}T^{\tau-1})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.81)
because the rate of $\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$ was already derived for the IC of DVS under both cases. Also, $\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see Westerlund,, 2018), and thus the $\sqrt{NT}$-consistency will hold under our assumptions, as well. Similarly,
$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|$ $\leq(NT)^{-1/2}\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$
×$\times\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|}_{O_{p}(T^{\tau})}$
=$=\begin{cases}O_{p}(N^{-1/2}T^{\tau-1/2}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.82)
because asymptotically $\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ behaves identically to $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, and thus the dominating component will have the same rate. Eventually,
[…]
which will be the dominating term.
If the parameter vector is heterogeneous, e.g. $\boldsymbol{\beta}_{i}=\boldsymbol{\beta}+\boldsymbol{\upsilon}_{i}$, where $\boldsymbol{\upsilon}_{i}$ is IID and mean-zero, then
$\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},$ (A.84)
where now $\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see e.g. Stauskas,, 2023). This means that now, for example,
$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq N^{-1}\underbrace{\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|}_{O_{p}(T^{\tau})}$
=$=\begin{cases}O_{p}(N^{-1}T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.85)
and other terms will behave similarly, while
$\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\upsilon}_{i}^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}\right\|=\begin{cases}O_{p}(T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.86)
[…]
as before, whereas the cross-terms are of lower order that could be made negligible by imposing restrictions on $N$ and $T$ divergence rate.
5.6 Additional Monte Carlo Results
5.6.1 Independent Idiosyncratics
Table 2: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.2) and (2.1). For $IC^{DVS}_{1}$ and $IC^{MW}_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $IC^{DVS}_{2}$ and $IC^{MW}_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
Table 3: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
5.6.2 Non-Stationary Idiosyncratics in $\mathbf{X}$
Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
$\tau=0$ $\tau=0.1$ $\tau=0.9$
$N$ $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
10 10 9.00 8.58 5.12 1.00 8.99 8.65 4.93 1.00 8.98 8.52 2.14 1.02
20 10 8.99 8.78 7.98 3.18 8.98 8.79 7.94 3.23 8.95 8.34 7.12 1.52
50 10 8.99 8.99 7.94 6.83 9.00 8.98 7.94 6.71 8.96 8.86 5.96 2.50
100 10 9.00 9.00 7.93 7.52 9.00 9.00 7.96 7.67 8.98 8.93 5.46 3.41
150 10 9.00 9.00 7.99 7.97 9.00 9.00 7.99 7.89 8.98 8.97 4.95 3.44
200 10 9.00 9.00 8.00 7.99 9.00 9.00 8.00 8.00 9.00 9.00 5.56 4.08
10 20 1.15 1.02 1.00 1.00 1.19 1.03 1.01 1.00 1.20 1.12 1.05 1.01
20 20 1.23 1.01 1.93 1.02 1.27 1.02 1.97 1.04 1.33 1.15 1.38 1.14
50 20 2.68 2.04 2.02 1.24 2.73 2.11 2.09 1.32 1.45 1.35 1.32 1.20
100 20 3.79 3.57 3.29 2.67 3.60 3.41 2.99 2.44 1.53 1.48 1.36 1.30
150 20 3.25 3.14 2.94 2.76 3.54 3.46 3.33 3.13 1.55 1.51 1.43 1.38
200 20 3.73 3.63 3.43 3.27 3.95 3.92 3.79 3.70 1.84 1.78 1.62 1.57
10 50 1.21 1.13 1.05 1.03 1.11 1.05 1.00 1.00 1.25 1.21 1.14 1.09
20 50 1.09 1.03 1.21 1.01 1.08 1.02 1.28 1.01 1.30 1.24 1.33 1.20
50 50 3.60 2.80 3.91 2.45 3.90 3.27 3.97 2.95 1.93 1.66 1.91 1.53
100 50 3.33 3.21 3.30 2.95 3.18 3.05 3.07 2.70 1.59 1.51 1.59 1.47
150 50 3.99 3.98 3.98 3.95 3.95 3.91 3.93 3.83 1.57 1.51 1.60 1.51
200 50 4.00 4.00 3.99 3.98 3.98 3.97 3.94 3.92 1.89 1.83 1.65 1.60
10 100 1.00 1.00 1.26 1.18 1.00 1.00 1.27 1.22 1.16 1.15 1.04 1.03
20 100 1.66 1.36 2.50 2.15 1.77 1.45 2.61 2.41 1.42 1.37 1.66 1.58
50 100 4.00 3.99 4.00 4.00 4.00 3.98 4.00 3.99 2.10 1.96 1.99 1.81
100 100 4.01 4.00 4.00 4.00 4.01 4.00 4.00 4.00 2.24 2.03 2.19 1.86
150 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.96 2.77 2.97 2.68
200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.54 2.43 2.54 2.35
10 150 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.12 1.10 1.10 1.08
20 150 1.01 1.00 3.09 2.13 1.01 1.01 3.08 2.14 1.16 1.14 1.62 1.52
50 150 3.27 3.17 3.01 2.78 3.99 3.95 3.95 3.89 1.50 1.42 2.20 2.01
100 150 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.25 2.08 2.74 2.38
150 150 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.92 2.70 2.86 2.65
200 150 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.10 2.94 3.13 2.96
10 200 1.16 1.08 1.02 1.01 1.28 1.16 1.02 1.01 1.35 1.33 1.24 1.23
20 200 2.98 2.96 3.97 3.86 2.95 2.90 3.90 3.74 1.64 1.59 1.55 1.50
50 200 4.07 4.04 4.00 4.00 4.07 4.03 4.00 4.00 1.67 1.59 1.84 1.74
100 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.45 2.34 2.50 2.25
150 200 4.00 4.00 3.99 3.98 4.00 4.00 3.95 3.91 2.22 2.11 2.49 2.36
200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.40 3.21 3.39 3.19
Table 4: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ are generated same as factors. $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
Table 5: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ are generated same as factors. $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
5.6.3 Objective Function Adjustment
Table 6: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ DVS criteria from De Vos and Stauskas, (2024), see (2.2). IC objective function adjusted to $\mathrm{IC}^{DVS}(M_{\mathbf{x}})=\ln\det\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}\right)+g\cdot k\cdot p_{N,T}$. For $DVS_{1}$: $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
5.6.4 Correlated Idiosyncratics
Table 7: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are weakly correlated time and units units.
Table 8: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are weakly correlated time and units units.
5.6.5 Eigenvalue Ratio
Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
$\tau=0$ $\tau=0.1$ $\tau=0.9$
$N$ $T$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$
10 10 6.80 7.12 7.87 5.98 6.79 7.05 7.67 5.90 4.33 6.17 5.93 5.31
20 10 7.21 7.41 8.18 5.95 7.14 7.42 8.11 5.81 5.43 6.94 7.05 5.75
50 10 6.83 7.41 8.05 4.72 6.73 7.40 7.99 4.63 4.57 6.85 6.20 4.85
100 10 6.71 7.41 7.62 3.88 6.69 7.40 7.73 3.86 3.98 6.34 5.67 3.76
150 10 6.33 7.12 7.60 4.15 6.28 7.09 7.54 4.26 3.48 5.61 5.60 4.12
200 10 6.02 6.68 7.73 4.11 5.97 6.62 7.65 4.07 3.09 4.75 5.27 3.78
10 20 3.14 6.95 2.19 4.46 3.11 6.84 2.12 4.37 2.92 4.76 1.90 3.47
20 20 6.01 7.62 3.41 3.00 5.94 7.62 3.58 2.99 3.48 6.87 2.38 3.07
50 20 4.96 7.64 2.52 3.69 4.83 7.63 2.75 3.67 2.54 6.46 1.86 2.75
100 20 4.14 7.33 3.35 4.00 4.05 7.37 3.24 3.89 1.57 5.73 1.48 3.19
150 20 3.81 6.02 3.43 3.78 3.95 6.09 3.38 3.76 1.54 3.97 1.54 3.55
200 20 3.99 4.92 3.69 3.92 4.02 4.90 3.82 3.97 1.43 3.06 1.68 3.40
10 50 4.46 7.84 1.20 2.88 4.32 7.84 1.24 2.85 2.61 6.26 1.33 2.82
20 50 6.90 7.94 1.66 2.92 6.88 7.93 1.70 2.88 4.16 7.35 1.73 2.89
50 50 3.97 7.82 2.64 3.60 4.00 7.83 2.81 3.52 1.95 7.36 1.53 3.15
100 50 3.69 5.26 2.66 4.51 3.55 5.55 2.58 4.35 1.63 4.66 1.39 2.96
150 50 4.00 4.59 3.19 4.13 3.95 4.85 3.01 4.16 1.43 3.46 1.34 2.71
200 50 4.00 4.02 3.28 2.92 4.00 4.03 3.23 2.97 1.29 2.50 1.37 2.80
10 100 3.00 7.87 3.39 3.95 2.98 7.87 3.28 3.81 2.65 6.09 1.51 3.42
20 100 6.03 7.99 1.19 3.71 6.27 7.99 1.23 3.53 3.91 7.38 1.53 2.84
50 100 3.92 7.98 1.55 3.38 4.00 7.98 1.78 3.41 2.42 7.42 1.42 3.05
100 100 4.00 5.80 2.44 3.00 4.00 5.37 2.36 2.95 1.67 5.22 1.49 2.66
150 100 4.00 4.00 2.64 5.50 4.00 4.00 2.64 5.42 1.67 3.40 1.72 2.46
200 100 4.00 4.00 4.08 4.76 4.00 4.00 4.08 4.24 1.54 2.70 1.59 2.23
10 150 4.00 6.42 3.18 3.24 4.00 6.92 3.08 3.38 2.21 5.25 1.47 2.60
20 150 6.42 8.00 1.05 4.77 6.49 8.00 1.12 4.57 5.17 7.41 1.66 3.23
50 150 3.60 7.91 2.47 2.98 3.96 7.90 2.41 2.97 3.29 6.54 1.37 2.84
100 150 4.00 4.52 2.87 4.89 4.00 4.29 2.73 4.92 1.66 5.51 1.67 2.36
150 150 4.00 4.00 2.78 5.53 4.00 4.00 2.57 5.45 1.78 3.75 1.82 2.57
200 150 4.00 4.00 4.51 5.24 4.00 4.00 4.47 5.23 1.79 2.88 1.69 2.34
10 200 4.00 7.98 3.52 3.46 4.00 7.97 3.54 3.37 2.22 6.99 1.83 2.91
20 200 4.06 8.00 2.69 3.77 4.12 8.00 2.79 3.65 4.38 7.42 1.90 2.81
50 200 4.00 8.00 2.05 4.32 4.00 8.00 2.05 3.97 3.77 7.62 1.59 2.77
100 200 4.00 4.24 2.05 5.25 4.00 4.23 1.94 5.30 1.93 5.17 1.63 2.49
150 200 3.99 3.95 2.25 4.33 3.97 3.89 2.36 3.79 1.77 4.24 1.49 1.79
200 200 4.00 4.00 4.26 5.50 4.00 4.00 4.29 5.46 1.78 2.64 1.76 2.48
Table 9: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$. $ER(\mathbf{X})$ selects CA from $\overline{\mathbf{X}}$, $ER(\mathbf{Z})$ from $(\overline{\mathbf{y}},\overline{\mathbf{X}})$. In case of $\widetilde{\mathbf{Z}}$ and $\widetilde{\mathbf{X}}$, the data are scaled by $\widehat{\boldsymbol{\Sigma}}_{\mathbf{Z}}^{-1/2}$, where $\widehat{\boldsymbol{\Sigma}}_{\mathbf{Z}}=\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{Z}_{i}-\overline{\mathbf{Z}})^{\prime}(\mathbf{Z}_{i}-\overline{\mathbf{Z}})$, see Juodis, (2022). Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.

### v2

@@
 Abstract
-Information criteria (IC) are important tools in the literature of factor models that allow one to estimate a typically unknown number of latent factors.
-Although first proposed for the Principal Components setting in the seminal work by Bai and Ng, (2002), it has recently been shown that IC perform extremely well in Common Correlated Effects (CCE) and related setups with stationary factors.
-In particular, they can consistently select a sufficient set of cross-section averages (CAs) to approximate the factor space.
-Given that CAs can proxy nonstationary factors, it is tempting to believe that the consistency of IC continues to hold under such generality.
-This study is a cautionary tale for practitioners.
-We demonstrate formally and in simulations that IC has a severe underselection issue even under very mild forms of factor non-stationarity, which goes against the sentiment in the CAs literature.
+Information criteria (IC) have been widely used in factor models to estimate an unknown number of latent factors.
+It has recently been shown that IC perform well in Common Correlated Effects (CCE) and related setups in selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors.
+As CAs can proxy non-stationary factors, it is tempting to claim such generality of IC, too.
+We show formally and in simulations that IC have a severe underselection issue even under very mild forms of factor non-stationarity, which goes against the sentiment in the literature.
+Keywords: Information criteria, factors, CCE, cross-section averages, non-stationary data
+MOS subject classification: 91B84
 1 Introduction
@@
 Here, $K=k+1$, and $i$ indexes individuals, such as countries or households.
-Then (1.1) represents a classical CCE setup of Pesaran, (2006):
-$\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},,\quad\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i},$ (1.2)
-where, $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter vector, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings).
+Then (1.1) represents a classical CCE setup of Pesaran, (2006) with $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ and $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$.
+Here, $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter vector, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings).
 Both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, and this induces cross-section dependence among individuals and endogeneity as the error in $\mathbf{y}_{i}$ is correlated with $\mathbf{X}_{i}$.
-(1.1) nests (1.2) with $\mathbf{C}_{i}=[\boldsymbol{\Gamma}_{i}\boldsymbol{\beta}+\boldsymbol{\gamma}_{i},\boldsymbol{\Gamma}_{i}]$ and $\mathbf{U}_{i}=[\mathbf{V}_{i}\boldsymbol{\beta}+\boldsymbol{\varepsilon}_{i},\mathbf{V}_{i}]$.
+(1.1) nests this model with $\mathbf{C}_{i}=[\boldsymbol{\Gamma}_{i}\boldsymbol{\beta}+\boldsymbol{\gamma}_{i},\boldsymbol{\Gamma}_{i}]$ and $\mathbf{U}_{i}=[\mathbf{V}_{i}\boldsymbol{\beta}+\boldsymbol{\varepsilon}_{i},\mathbf{V}_{i}]$.
 Example 2.
 Distinct Correlated Effects (DCE).
-We can introduce distinct but correlated factors $\mathbf{F}=[\mathbf{F}_{\mathbf{x}},\mathbf{F}_{\mathbf{y}}]\in\mathbb{R}^{T\times(m_{x}+m_{y})}$ in Example 1, which makes the setup mode applicable:
-$\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}_{\mathbf{y}}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},,\quad\mathbf{X}_{i}=\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$ (1.3)
-as considered in Cui et al., (2022) or Stauskas and De Vos, (2025).
-Here, (1.1) nests (1.3) similarly to Example 1, but $\boldsymbol{\gamma}_{i}$ is modified accordingly, depending on whether $\mathbf{F}_{\mathbf{y}}\subset\mathbf{F}_{\mathbf{x}}$, vice versa or $\mathbf{F}_{\mathbf{x}}\cap\mathbf{F}_{\mathbf{y}}=\emptyset$.
+We can introduce distinct but correlated factors $\mathbf{F}=[\mathbf{F}_{\mathbf{x}},\mathbf{F}_{\mathbf{y}}]\in\mathbb{R}^{T\times(m_{x}+m_{y})}$ in Example 1, which makes the setup mode applicable: $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}_{\mathbf{y}}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$, $\mathbf{X}_{i}=\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$ as considered in Cui et al., (2022) or Stauskas and De Vos, (2025).
+Here, (1.1) nests such model similarly to Example 1, but $\boldsymbol{\gamma}_{i}$ is modified accordingly, depending on whether $\mathbf{F}_{\mathbf{y}}\subset\mathbf{F}_{\mathbf{x}}$, vice versa or $\mathbf{F}_{\mathbf{x}}\cap\mathbf{F}_{\mathbf{y}}=\emptyset$.
 Example 3.
 Block Structure (FRED-MD-QD , EA-MD-QD).
-The set of regressors in Examples 1 and 2 can be high-dimensional, but follow a natural group structure with $k$ blocks and $m$ underlying factors.
-Let $i_{j}$ index series belonging to a block $j$ for $j=1,\ldots,k$.
-Then $\mathbf{Z}_{i}$ is $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}=[\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{k}}]^{\prime}$.
+The regressors in Examples 1 and 2 can be high-dimensional, but follow a natural structure of $k$ blocks and $m$ underlying factors.
+Let $i_{j}$ index series in block $j$ for $j=1,\ldots,k$.
+Then $\mathbf{Z}_{i}=\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}=[\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{k}}]^{\prime}$.
 Here, $N$ is the number of series per block.
-Blocks, such as monetary, industrial or fiscal series can be found in FRED-MD (see McCracken and Ng,, 2016), FRED-QD (see McCracken and Ng,, 2020) or EA-MD-QD (see Barigozzi et al.,, 2024) datasets.
+Monetary, industrial or fiscal blocks of series can be found in FRED-MD (see McCracken and Ng,, 2016), FRED-QD (see McCracken and Ng,, 2020) or EA-MD-QD (see Barigozzi et al.,, 2024) datasets.
 These datasets have been very popular in factor-augmented forecasts (see e.g.
-Karabiyik et al.,, 2019, Stauskas and Westerlund,, 2022, or Morico and Stauskas,, 2025).
+Karabiyik and Westerlund,, 2021, or Morico and Stauskas,, 2025).
 In all the cases above, we can estimate the factors by taking $K$ cross-section averages (CAs): $\widehat{\mathbf{F}}=\overline{\mathbf{Z}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g.
@@
 In Example 2, only $\widehat{\mathbf{F}}_{\mathbf{x}}=\overline{\mathbf{X}}$ is used as a regressor, while $\mathbf{F}_{\mathbf{y}}$ is controlled by bootstrap (see Stauskas and De Vos,, 2025).
-In Example 3, coefficient of $\widehat{\mathbf{F}}$ itself matters for forecast evaluation (see e.g.
+In Example 3, the coefficient of $\widehat{\mathbf{F}}$ itself matters for forecast evaluation (see e.g.
 Massacci and Kapetanios,, 2024).
 A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties.
-Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings (see e.g.
+Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings without any modifications (see e.g.
 Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023).
-This property holds as long as the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs ($K\in\{k,k+1\}$ in Examples 1-3).
+This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs ($K\in\{k,k+1\}$ in Examples 1-3).
 However, there are reasons to ensure that $m=K$.
@@
 In Example 3, popular tests of forecast accuracy comparison have highly non-standard asymptotic distributions (see Stauskas and Westerlund,, 2022).
-More generally, inclusion of too many CAs might even lead to a model that over-fits the data.
-In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis et al.,, 2021; Juodis,, 2022, for a discussion).
-In response to this, Margaritella and Westerlund, (2023) were the first to demonstrate that Information Criteria (IC) of Bai and Ng, (2002) from Principal Components (PC) literature can be consistently applied in the pure CCE setting in Example 1.
-Subsequently, De Vos and Stauskas, (2024) showed similar results applicable to more general Examples 2 and 3.111Juodis, (2022) considers eigenvalue-based selection (see also Ahn and Horenstein,, 2013).
-In that study, factors are stationary, as well.
-Another set of results concern fixed-$T$, where properties of $\mathbf{F}$ do not play an asymptotic role.
-Also, the method focuses solely on the number of factors, but cannot select the CAs.
+In general, too many CAs might even lead to a model that over-fits the data.
+In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion).
+In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that Information Criteria (IC) of Bai and Ng, (2002) from Principal Components (PC) literature can be applied in the CCE setting in Example 1 under stationary factors to identify an optimal set of CAs.
+De Vos and Stauskas, (2024) (DVS) showed similar results applicable to more general Examples 2 and 3.111Juodis, (2022) considers eigenvalue-based methods of Ahn and Horenstein, (2013)) under stationarity.
+The focus there is the number of factors but not the optimal set of CAs.
 We leave analysis of this approach for the future research.
-In this study, we re-visit IC proposed in Margaritella and Westerlund, (2023) and De Vos and Stauskas, (2024).
-Given the power of CAs to proxy very general factor structures, it is only natural to evaluate IC under nonstationarity conditions.
-Contrary to the sentiment in the CCE/CAs literature, IC crumble under very mild forms of nostationarity.
+MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC.
+The power of CAs to proxy general factor structures makes it natural to evaluate this statement and re-visit IC proposed in MW and DVS.
+We show that, contrary to the sentiment in the CCE/CAs literature, IC crumble under very mild forms of nostationarity.
 We demonstrate formally and in simulations that they exhibit a severe under-selection problem, which is independent of penalties used in the objective functions of IC.
@@
 2 Econometric Setup
-To further operationalize our examples, we introduce $M$ which is a set of column indices of $\overline{\mathbf{Z}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{K\times g}$ picks the corresponding $g=\mathrm{cols}(\mathbf{q}_{M})$ averages in practice.
+To further operationalize our examples, we introduce $M$ which is a set of column indices of $\overline{\mathbf{Z}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{K\times g}$ picks the corresponding $g$ averages in practice.
 For instance, $\overline{\mathbf{Z}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k+1$ CAs in the typical CCE setup.
@@
 Clearly, this number can vary among our examples.
-Let accordingly $M_{0}$ ($M_{\mathbf{x},0}$) denote the set of averages from $\overline{\mathbf{Z}}$ ($\overline{\mathbf{X}}$) such that $\mathrm{rk}(\overline{\mathbf{C}}\mathbf{q}_{M_{0}})=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x},0}})=m$, and $|M_{0}|=|M_{\mathbf{x},0}|=m$, where $|M|$ denotes cardinality of an arbitrary set $M$.
+Let accordingly $M_{0}$ ($M_{\mathbf{x},0}$) denote the set of averages from $\overline{\mathbf{Z}}$ ($\overline{\mathbf{X}}$) such that $\mathrm{rank}(\overline{\mathbf{C}}\mathbf{q}_{M_{0}})=\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x},0}})=m$, and $|M_{0}|=|M_{\mathbf{x},0}|=m$, where $|M|$ denotes cardinality of an arbitrary set $M$.
 IC under consideration are given by
@@
 $\mathrm{IC}^{DVS}(M_{\mathbf{x}})=\ln\det\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}\right)+g\cdot k\cdot p_{N,T},$ (2.2)
-$p_{N,T}=\begin{cases}\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\\ \frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right).\end{cases}$ (2.3)
-Here and further, MW represents Margaritella and Westerlund, (2023), while DVS stands for De Vos and Stauskas, (2024).
-$\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix onto the orthogonal complement of $\mathbf{A}$, $\mathbf{A}^{+}$ is a Moore-Penrose (MP) inverse, while $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ with $\widehat{\boldsymbol{\beta}}$ is obtained with the CCE estimator under all available $k+1$ CAs.
+$p_{N,T}\in\left\{\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right)\right\}$ (2.3)
+where $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix onto the orthogonal complement of $\mathbf{A}$, $\mathbf{A}^{+}$ is a Moore-Penrose (MP) inverse, while $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ with $\widehat{\boldsymbol{\beta}}$ is obtained with the CCE estimator under all available $k+1$ CAs.
 Finally, $p_{N,T}$ is a penalty term which goes to 0 with $(N,T)\to\infty$.
 Examples of feasible and most popular penalties are given in (2.3) for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002).
+Note that (2.1) - (2.2) are functions of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
 Let $\overline{M}$ ($\overline{M}_{\mathbf{x}}$) be the set of indices of all available CAs that $|\overline{M}|=k+1$ ($|\overline{M}_{\mathbf{x}}|=k$). Then
@@
 Also, $\boldsymbol{\xi}_{t}=\mathrm{vec}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}-\mathbb{E}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right)\right)$ is strong mixing with coefficients of size $-bd(b-d)$ with $b>4$ and $b>d>2$, $\mathbb{E}(\|\boldsymbol{\xi}_{t}\|^{b})<\infty$, and $\lim_{T\to\infty}T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{s}^{\prime})$ is positive definite.
+Here and throughout, $\left\|\mathbf{A}\right\|=\sqrt{\mathrm{trace}(\mathbf{A}^{\prime}\mathbf{A})}$ is Frobenius norm.
     • (iii)
@@
 If $m=K$, then $\overline{\mathbf{C}}=\overline{\mathbf{C}}_{m}$.
-Assumption 1 (a) parameterizes factors very flexibly and allows to perform comparative statics, since $\tau\to 1$ implies that the process becomes more persistent and approaches the usual local-to-unity process.
-We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allows us vary $\tau$ and document its effect on IC performance.
-This assumption also implies that $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$ which is a constant positive definite matrix.
-Its part (b) ensures that we return to the usual stationarity conditions under $\tau=0$.
+Assumption 1 (a) treats factors very flexibly and allows us perform comparative statics, as $\tau\to 1$ increases persistence of the process.
+We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allow us vary $\tau$ and give theoretical guarantees, such as $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$, which is a constant positive definite matrix.
+Part (b) ensures that we return to the usual stationarity conditions under $\tau=0$.
 Assumption 2 (a) is split into two parts.
-If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}$ converges to a zero matrix too slow for our purposes, which would dominate and obscure the very effect of factor persistence on IC, and we would like to avoid this (see also Lemma 3.1 in Magdalinos and Phillips,, 2009).
-For completeness, we present simulation evidence under serial correlation, as well, which does not demonstrate a major negative effect on IC.
-The rest of the assumptions are standard, as they ensure mutual independence of factors and the error terms (Assumption 3), weak cross-section dependence of idiosyncratics (Assumption 2 (b)) and informativeness of the factor loadings (rank condition in Assumption 4).
+If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}$ converges to zero too slow for our purposes, which would dominate and obscure the very effect of factor persistence on IC, and we would like to avoid this (see Lemma 3.1 in Magdalinos and Phillips,, 2009).
+We present simulations under serial correlation in the Supplement, which does not demonstrate a major negative effect on IC.
+The rest of the assumptions are standard, as they ensure mutual independence of factors and the error terms (Assumption 3), weak cross-section dependence of idiosyncratics (Assumption 2 (b)) and informativeness of the loadings (rank condition in Assumption 4).
 The loadings are random as in e.g.
-Pesaran, (2006), but they can be treated as deterministic parameters, as well.
+Pesaran, (2006), but they can also be treated as deterministic parameters.
 To formulate our key theoretical result, let us denote the objective function component in (2.1) and (2.2) evaluated at some set of CAs $M$ by $\mathcal{V}_{j}$ for $j\in\{MW,DVS\}$.
@@
 Also, for illustration purposes let $j=DVS$.
-Then $\mathcal{V}_{DVS}$ (and $\mathcal{V}_{DVS}^{0})$ is exactly (log-determinant of) the denominator of the CCE estimator, which is consistent for a positive definite matrix under general factors (see Westerlund,, 2018), and a similar argument applies to the objective function of MW.
+Then $\mathcal{V}_{DVS}$ (and $\mathcal{V}_{DVS}^{0})$ is exactly (log-determinant of) the denominator of the CCE estimator.
+It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x}}})^{+}+o_{p}(1)$, and by inserting this into $\mathbf{X}_{i}$, we can show that
+$\mathcal{V}_{DVS}=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}\right)+o_{p}(1)=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}\right)+O_{p}(T^{-1}),$ (2.6)
+whose argument is consistent for a positive definite matrix under general factors (see Westerlund,, 2018), and a similar logic applies to the objective function of MW.
 Thus, it is natural that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}$ is negligible.
@@
 In general, under $M_{0}\subset M$, the usual CCE logic prevails deceptively.
-Under $M\subset M_{0}$, Assumption 4 is not satisfied, and the rank condition does not hold.
-This means that $m-g$ factors are not asymptotically projected out.
-Given that $\mathcal{V}_{j}$ is a quadratic form for both $j\in\{MW,DVS\}$, their second empirical moment is of order $O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a), but the objective function offers normalization by $T$ only.
+Under $M\subset M_{0}$, the rank condition in Assumption 4 is not satisfied.
+This means that $m-g$ factors are not asymptotically projected out, and the representation in (2.6) is invalid.
+Given that $\mathcal{V}_{j}$ is a quadratic form for both $j\in\{MW,DVS\}$, their second empirical moment is of the order $O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a).
+However, the objective function offers normalization by $T$ only.
 Exactly this imbalance induces divergence, and this implies that the probability that IC are minimized at some other set contained in $M_{0}$ does not necessarily go to 0.
-Under stationary $\mathbf{F}$, however, it is clear that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=O_{p}(1)$ and, in fact, $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}\to_{p}c>0$ as $(N,T)\to\infty$ (see the analysis of MW and DVS).
+Under stationary $\mathbf{F}$, however, it is clear that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=O_{p}(1)$ and, in fact, $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}\to_{p}c>0$ as $(N,T)\to\infty$ as required (see the analysis of MW, DVS and Bai and Ng,, 2002).
 Our Proposition 1 naturally provides the same result under $\tau=0$.
-Note that the failure of both IC is purely driven by the objective function and is not related to the penalty $p_{N,T}$.
-In fact, the only theoretical prescription for $p_{N,T}$ depends only on $C_{N,T}^{2}$ and not $\tau$, as given in (2.5).
-The discussion above seems to suggest that incorporating the degree of non-stationarity of $\mathbf{F}$ into the objective function should solve the issue and allow keeping $p_{N,T}$ the same.
-However, this is not the case.
-Let us assume that $\tau$ is known and normalize the objective function by $T^{1+\tau}$, instead, as suggested by Assumption 1 (a).
-In the Supplement, we show that
-$\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}=\ln\mathrm{det}\left[\mathbf{I}_{k}+\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x}}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}})\mathbf{X}_{i}}_{O_{p}(1)}\underbrace{\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}}\mathbf{X}_{i}\right)^{-1}}_{O_{p}(T^{\tau})}\right],$ (2.6)
-which means that, again, the determinant does not converge to a strictly positive constant.
-The reason behind this is that the rank condition is satisfied in the second component of the product, and by following steps of Karabiyik et al., (2017) and Westerlund, (2018), we obtain
-$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}}\mathbf{X}_{i}=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}+o_{p}(1)=O_{p}(T^{-\tau}),$ (2.7)
-since the idiosyncratics are stationary.
-In effect, the new normalization induces imbalance elsewhere.
+Note that the failure of both IC is purely driven by the objective function, and it complicates the role of $p_{N,T}$.
+If $M_{0}\subset M$, $p_{N,T}$ should simply vanish slower than $C_{N,T}^{-2}$, but under $M\subset M_{0}$ the penalty has to counter divergence.
+This discussion suggests that accounting for non-stationarity of $\mathbf{F}$ in the objective function should solve the issue, but this is not the case.
+Assuming that $\tau$ is known, we normalize the objective function by $T^{1+\tau}$ as Assumption 1 (a) suggests.
+In the Supplement, we show that under $M\subset M_{0}$:
+$\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}=\ln\mathrm{det}\left[\mathbf{I}_{k}+\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x}}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}})\mathbf{X}_{i}}_{\mathbf{I}=O_{p}(1)}\left(\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}}_{\mathbf{II}=O_{p}(T^{-\tau})}\right)^{-1}\right]+o_{p}(1),$ (2.7)
+where the determinant does not converge to a strictly positive constant, again.
+The reason behind this is that the rank condition is satisfied in $\mathbf{II}$, and by following the steps of Karabiyik et al., (2017) and Westerlund, (2018), it admits (2.6).
+However, since idiosyncratics are stationary, “over-normalization” causes singularity, and we get a new imbalance between $\mathbf{I}$ and $\mathbf{II}$.
 Therefore, IC fundamentally depend not only on stationarity of factors, but also on stationarity of idiosyncratics.
-We confirm this in the additional Monte Carlo simulations presented in the Supplement, where the adjustment of the objective function becomes more harmful with $\tau$.
+We confirm this with the additional simulations in the Supplement, where this adjustment becomes more harmful with $\tau$.
+If $M_{0}\subset M$, (2.5) and normalization by $T^{1+\tau}$ imply that (2.7) is still $\ln\det(\mathbf{I}_{k}+O_{p}(T^{-\tau}C_{N,T}^{-2})\times O_{p}(T^{\tau}))=O_{p}(C_{N,T}^{-2})$, which again means that $p_{N,T}$ needs to have different properties depending on whether $M_{0}\subset M$ or $M\subset M_{0}$.
 Remark 1.
@@
 3 Simulation Study
-The data generating process of the simulation follows Margaritella and Westerlund, (2023):
+The data generating process of the simulation follows MW:
 $\mathbf{y}_{i}$ $=\mathbf{X}_{i}\boldsymbol{\beta}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},\quad\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i},$ (3.1)
-where $\mathbf{X}_{i}=\left[\mathbf{x}_{i,1},...,\mathbf{x}_{i,T}\right]^{\prime}$ is a $T\times k$ matrix of observable regressors, $\boldsymbol{\beta}=\iota^{\prime}0.5$ a $k\times 1$ vector of unobserved parameters set to 0.5.
-$\mathbf{F}$ is a $T\times m$ matrix of $m$ unobserved common factors.
-For simplicity, we assume that $\mathbf{X}_{i}$ and $\mathbf{u}_{i}$ load on $m$ factors.
+where $\mathbf{X}_{i}=\left[\mathbf{x}_{i,1},...,\mathbf{x}_{i,T}\right]^{\prime}$ is a $T\times k$ matrix of observable regressors, $\boldsymbol{\beta}=\iota^{\prime}0.5$ a $k\times 1$ vector of unobserved parameters.
 The factor loadings are generated as $\boldsymbol{\Gamma}_{i}=[\mathbf{I}_{m},\mathbf{0}_{m\times(k-m)}]\psi_{i}$, that is, a matrix $m\times k$, with $\psi_{i}\sim N(1,1)$ and the elements of the vector $1\times m$ $\boldsymbol{\gamma}_{i}$ are drawn from $N(1,1)$.
@@
 In the case of errors correlated weakly across time and cross-sections we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
-$\mathbf{e}_{i,t}=(\epsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}=\mathbf{P}_{\rho}\mathbf{e}_{i,t-1}+\mathbf{z}_{\mathbf{w},i,t},\quad\mathbf{P}_{\rho}=\begin{bmatrix}\rho&\mathbf{0}_{1\times k}\\ \mathbf{0}_{k\times 1}&\rho_{v}\mathbf{I}_{k}\end{bmatrix},\quad\mathbf{z}_{\mathbf{w},i,t}=\begin{bmatrix}\sqrt{\frac{m\left(1-\rho^{2}\right)}{1+2J\kappa^{2}}}\boldsymbol{\epsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa\mathbf{w}_{i}^{\prime}\right)\\ \sqrt{\frac{\left(1-\rho_{v}^{2}\right)}{1+2J_{v}\kappa_{v}^{2}}}\boldsymbol{\Upsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa_{v}\boldsymbol{\tilde{\mathbf{w}}}_{i}^{\prime}\right)\end{bmatrix}$ (3.2)
+[…]
 where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero.
@@
 $\rho_{v}=0$ for the time IID case and $\rho_{v}=0.5$ for weakly time-correlated idiosyncratics.
-In both cases, we allow for weak correlation across cross sections, $\kappa_{v}=0.2$, $J_{v}=5$ and $\tilde{\mathbf{w}}_{i}$ is defined as $\mathbf{w}_{i}$ above.
+In both cases, we allow for weak correlation across cross-sections, $\kappa_{v}=0.2$, $J_{v}=5$ and $\tilde{\mathbf{w}}_{i}$ is defined as $\mathbf{w}_{i}$ above.
 Therefore, we also test correlated idiosyncratics over time, whereas our theory prescribes uncorrelated innovations for $\tau\in(0,1)$.
-The factors are drawn as:
-$\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{Q}T^{-\tau},\quad\mathbf{Q}=diag(q_{1},...,q_{m}),q_{j}\sim U(0,2)$ (3.3)
-$\mathbf{u}_{t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.4)
+The factors are drawn as $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t}$, where
+$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{Q}T^{-\tau},\quad\mathbf{Q}=diag(q_{1},...,q_{m}),q_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.3)
 In the stationary factors case, the factors are drawn as a stable VAR with $\tau=0$.
@@
 Moreover, we present extended tables with extra $N,T=150,200$ for all our experiments.
-Table 3 below presents the average of the number of CAs selected for the case of stationary factors ($\tau=0$), the near stationary case and the close to local-to-unity process.
+Table 6.6.5 below presents the average of the number of CAs selected for the case of stationary factors ($\tau=0$), the near stationary case and the close to local-to-unity process.
 The idiosynchratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units.
+The Supplement presents additional results with weak correlation over time.
 We start with the in the literature established case of stationary factors.
 In the small $T$ cases, $T=10$ and $T=20$, we first observe an overselection of the number of CAs, followed by an underselection.
-This pattern resembles the results in Margaritella and Westerlund, (2023) and is consistent across all values for $N$ and $\tau$ [^4] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$ If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors.
+This pattern resembles the results in MW and is consistent across all values for $N$ and $\tau$ [^4] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$ If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors.
 Once the number of cross-sectional units and time periods passes 50, all four criteria select close to or exactly the correct number of factors.
@@
 Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ uncorrelated over time, but weakly correlated across units, see (3.2).
-To investigate the behavior of the IC over different levels of non-stationarity, we increase $\tau$ in $0.05$ steps from 00 to $0.95$.
-Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$.
+Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$, where we increase $\tau$ in $0.05$ steps from 00 to $0.95$.
 We also allow the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ to be uncorrelated over time ($\rho=\rho_{v}=0$, in blue) and weakly correlated ($\rho=\rho_{v}=0.5$, in red).
@@
 Both are weakly correlated across units.
+4 Illustration: a Discussion of Serlenga and Shin, (2007)
+MW apply $IC_{1}^{MW}$ and $IC_{2}^{MW}$ to revisit the study of Serlenga and Shin, (2007) (SS), who investigate the gravity equation of bilateral trade flows among 15 European countries ($N=91$ trading pairs) over 1960–2001 by the means of the CCE estimator.
+While SS include CAs arbitrarily, MW motivate their usage of IC as a theoretically backed and data-driven selection method.
+In neither of the studies are the time series properties of the variables examined.
+Variables of interest include $\mathbf{z}_{i,t}=(\textit{Trade}_{i,t},\textit{RER}_{i,t},\textit{TGDP}_{i,t},\textit{RLF}_{i,t},\textit{SIM}_{i,t},\textit{EEC}_{i,t},\textit{EMU}_{i,t})^{\prime}$, where $\textit{Trade}_{i,t}$ (the dependent variable) denotes a total bilateral trade flow for a trading pair $i$, $\textit{RER}_{i,t}$ is the log real exchange rate between the pair currencies, $\textit{TGDP}_{i,t}$ represents the economic mass of the partners, $\textit{RLF}_{i,t}$ measures differences in the relative factor endowments.
+Moreover, $\textit{SIM}_{i,t}$ is the logarithm of a relative economic size, while dummies $\textit{EEC}_{i,t}$ and $\textit{EMU}_{i,t}$ represent European Economic Community and the European Monetary Union, respectively.
+Finally, $\textit{RERT}_{t}$ (log real exchange rate between the EU currencies and the dollar), a constant ($1$) and a linear trend ($t$) are treated as observed factors.
+Therefore, $\overline{M}$ constitutes 10 CAs to select from, and MW select $\widehat{M}=\{1,t,\textit{RERT}_{t},\overline{Trade}_{t},\overline{TDGP}_{t},\overline{RER}_{t},\overline{\textit{EMU}}_{t}\}$ with $g=7$ to control for $\mathbf{F}$.
+$\widehat{M}$ minimizes the remaining CS correlation in the CCE residuals (see Table 1 in MW).
+We examine the series in SS further and detect strong evidence of unit roots in $\mathbf{F}$.
+Table 2 presents the results of the panel unit root test of Reese and Westerlund, (2016), which is the CAs equivalent of PANIC of Bai and Ng, (2004).
+Although we find strong evidence in favor of stationary idiosyncratics (at $1\%$ based on the PMSB statistic), we do not reject the null of unit root factors in virtually all series.
+[^5] The dummies $\textit{EEC}_{i,t}$ and $\textit{EMU}_{i,t}$ are omitted as they are unlikely to follow a typical factor structure.
+While MW acknowledge sensitivity of the results, our theory only strengthens the fact that the CCE-estimated coefficients in their Table 1 should be interpreted with some caution, and perhaps alternative combinations of CAs with $g\leq 10$ can provide a clearer picture.
+Series $Trade$ $TDGP$ $RER$ $RLF$ $SIM$ $RERT$
+Unit Roots in Factors
+Errors
+Table 2: PANICCA of Reese and Westerlund, (2016). - null of unit root not rejected; - the null rejected. For each series, the rest are used as regressors.
 5 Conclusion
-As predicted by the theory and our simulation results, the inability of IC to accommodate non-stationary factors is evident.
+Our theory and simulation results show the inability of IC to accommodate non-stationary factors.
 The number of CAs is severely underselected, independent of the number of time periods and units.
-The underselection has severe consequences in terms of our examples.
-For the CCE estimator, since the number of factors is larger than the number of CAs, the rank condition fails and therefore consistency of the CCE estimator as well.
+Underselection has severe consequences in terms of our examples.
+For CCE, the number of factors may be larger than the selected number of CAs, which makes the estimator inconsistent.
 In DCE case, the bootstrap procedure cannot help accommodate different factors in $\mathbf{X}$ and $\mathbf{y}$ anymore.
-In forecasting, incorrectly estimated factors harm the forecast comparison procedures.
+In forecasting, incorrectly estimated factors may harm the forecast comparison procedures.
 6 Appendix
@@
 without further restrictions as $\tau\in(0,1)$. Both of these results will be important when determining the orders of many terms.
-Next, in case of $M\subset M_{0}$, we have that $\mathrm{rk}\left(\frac{1}{T^{1+\tau}}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)=m$ even asymptotically, and hence
+Next, in case of $M\subset M_{0}$, we have that $\mathrm{rank}\left(\frac{1}{T^{1+\tau}}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)=m$ even asymptotically, and hence
 $\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}=\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}+O_{p}(\xi_{N,T,\tau}^{-1}).$ (A.15)
@@
 $\overline{\mathbf{H}}_{M}=\begin{bmatrix}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}&-\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\overline{\boldsymbol{\Gamma}}_{M_{0}^{C}}\\ \mathbf{0}_{(g-m)\times m}&\mathbf{I}_{g-m}\end{bmatrix}=\begin{bmatrix}\overline{\mathbf{H}}_{M_{0}},\overline{\mathbf{H}}_{M_{0}^{C}}\end{bmatrix},$ (A.41)
-with the obvious definitions of $\overline{\mathbf{H}}_{M_{0}}$ and $\overline{\mathbf{H}}_{M_{0}^{C}}$ and $\mathrm{rk}(\overline{\mathbf{H}}_{M})=g$.
+with the obvious definitions of $\overline{\mathbf{H}}_{M_{0}}$ and $\overline{\mathbf{H}}_{M_{0}^{C}}$ and $\mathrm{rank}(\overline{\mathbf{H}}_{M})=g$.
 In what follows, the post-multiplication by this matrix leads to
@@
 Let us use the fact that $\mathbf{q}_{M_{0}}$ gives the minimal set of averages that asymptotically span the space of $\mathbf{F}_{\mathbf{x}}$.
-This means that $\mathrm{rk}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}_{M_{0}})=m$ and this matrix is $m\times m$.
+This means that $\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=\mathrm{rank}(\overline{\boldsymbol{\Gamma}}_{M_{0}})=m$ and this matrix is $m\times m$.
 This means that
@@
 6.6.2 Non-Stationary Idiosyncratics in $\mathbf{X}$
-Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
-$\tau=0$ $\tau=0.1$ $\tau=0.9$
-$N$ $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
-10 10 9.00 8.58 5.12 1.00 8.99 8.65 4.93 1.00 8.98 8.52 2.14 1.02
-20 10 8.99 8.78 7.98 3.18 8.98 8.79 7.94 3.23 8.95 8.34 7.12 1.52
-50 10 8.99 8.99 7.94 6.83 9.00 8.98 7.94 6.71 8.96 8.86 5.96 2.50
-100 10 9.00 9.00 7.93 7.52 9.00 9.00 7.96 7.67 8.98 8.93 5.46 3.41
-150 10 9.00 9.00 7.99 7.97 9.00 9.00 7.99 7.89 8.98 8.97 4.95 3.44
-200 10 9.00 9.00 8.00 7.99 9.00 9.00 8.00 8.00 9.00 9.00 5.56 4.08
-10 20 1.15 1.02 1.00 1.00 1.19 1.03 1.01 1.00 1.20 1.12 1.05 1.01
-20 20 1.23 1.01 1.93 1.02 1.27 1.02 1.97 1.04 1.33 1.15 1.38 1.14
-50 20 2.68 2.04 2.02 1.24 2.73 2.11 2.09 1.32 1.45 1.35 1.32 1.20
-100 20 3.79 3.57 3.29 2.67 3.60 3.41 2.99 2.44 1.53 1.48 1.36 1.30
-150 20 3.25 3.14 2.94 2.76 3.54 3.46 3.33 3.13 1.55 1.51 1.43 1.38
-200 20 3.73 3.63 3.43 3.27 3.95 3.92 3.79 3.70 1.84 1.78 1.62 1.57
-10 50 1.21 1.13 1.05 1.03 1.11 1.05 1.00 1.00 1.25 1.21 1.14 1.09
-20 50 1.09 1.03 1.21 1.01 1.08 1.02 1.28 1.01 1.30 1.24 1.33 1.20
-50 50 3.60 2.80 3.91 2.45 3.90 3.27 3.97 2.95 1.93 1.66 1.91 1.53
-100 50 3.33 3.21 3.30 2.95 3.18 3.05 3.07 2.70 1.59 1.51 1.59 1.47
-150 50 3.99 3.98 3.98 3.95 3.95 3.91 3.93 3.83 1.57 1.51 1.60 1.51
-200 50 4.00 4.00 3.99 3.98 3.98 3.97 3.94 3.92 1.89 1.83 1.65 1.60
-10 100 1.00 1.00 1.26 1.18 1.00 1.00 1.27 1.22 1.16 1.15 1.04 1.03
-20 100 1.66 1.36 2.50 2.15 1.77 1.45 2.61 2.41 1.42 1.37 1.66 1.58
-50 100 4.00 3.99 4.00 4.00 4.00 3.98 4.00 3.99 2.10 1.96 1.99 1.81
-100 100 4.01 4.00 4.00 4.00 4.01 4.00 4.00 4.00 2.24 2.03 2.19 1.86
-150 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.96 2.77 2.97 2.68
-200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.54 2.43 2.54 2.35
-10 150 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.12 1.10 1.10 1.08
-20 150 1.01 1.00 3.09 2.13 1.01 1.01 3.08 2.14 1.16 1.14 1.62 1.52
-50 150 3.27 3.17 3.01 2.78 3.99 3.95 3.95 3.89 1.50 1.42 2.20 2.01
-100 150 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.25 2.08 2.74 2.38
-150 150 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.92 2.70 2.86 2.65
-200 150 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.10 2.94 3.13 2.96
-10 200 1.16 1.08 1.02 1.01 1.28 1.16 1.02 1.01 1.35 1.33 1.24 1.23
-20 200 2.98 2.96 3.97 3.86 2.95 2.90 3.90 3.74 1.64 1.59 1.55 1.50
-50 200 4.07 4.04 4.00 4.00 4.07 4.03 4.00 4.00 1.67 1.59 1.84 1.74
-100 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.45 2.34 2.50 2.25
-150 200 4.00 4.00 3.99 3.98 4.00 4.00 3.95 3.91 2.22 2.11 2.49 2.36
-200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.40 3.21 3.39 3.19
 Table 5: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$.

### v3

@@
 Abstract
-Information criteria (IC) have been widely used in factor models to estimate an unknown number of latent factors.
-It has recently been shown that IC perform well in Common Correlated Effects (CCE) and related setups in selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors.
-As CAs can proxy non-stationary factors, it is tempting to claim such generality of IC, too.
-We show formally and in simulations that IC have a severe underselection issue even under very mild forms of factor non-stationarity, which goes against the sentiment in the literature.
+Information criteria (ICs) have been widely used in factor models to estimate an unknown number of latent factors.
+It has recently been shown that ICs perform well in Common Correlated Effects (CCE) and related settings when selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors.
+As CAs can proxy non-stationary factors, it is tempting to claim an excellent performance of ICs under general factors, too.
+We show formally and in simulations that they remain consistent, but the more persistent factors are, the poorer they perform in small samples, which goes against the sentiment in the CCE literature.
+Abstract
+In this Supplementary Material, we provide the proofs of our auxiliary results and our main result in Proposition 1 (Sections 2 and 3).
+In addition, we demonstrate how the proofs change in the case of the IC of Margaritella and Westerlund, (2023).
+In Section 4, we provide additional simulation evidence, including, e.g., correlated idiosyncratics or eigenvalue ratio method of Juodis, (2022).
 Keywords: Information criteria, factors, CCE, cross-section averages, non-stationary data
@@
 1 Introduction
-Consider the following set of $K$ variables that admits a factor structure for $i=1,\ldots,N$:
+Consider a set of $K$ variables (stacked over time $t=1,\ldots,T$) that admits a factor structure for cross-sectional units $i=1,\ldots,N$:
 $\mathbf{Z}_{i}=\mathbf{F}\mathbf{C}_{i}+\mathbf{U}_{i},$ (1.1)
-where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is the error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ is latent common factors that can be either stationary or not, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix.
-(1.1) nests several popular settings.
-Example 1.
-Common Correlated Effects (CCE).
-Let $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables and $\mathbf{y}_{i}$ is the dependent variable.
-Here, $K=k+1$, and $i$ indexes individuals, such as countries or households.
-Then (1.1) represents a classical CCE setup of Pesaran, (2006) with $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ and $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$.
-Here, $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter vector, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings).
-Both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, and this induces cross-section dependence among individuals and endogeneity as the error in $\mathbf{y}_{i}$ is correlated with $\mathbf{X}_{i}$.
-(1.1) nests this model with $\mathbf{C}_{i}=[\boldsymbol{\Gamma}_{i}\boldsymbol{\beta}+\boldsymbol{\gamma}_{i},\boldsymbol{\Gamma}_{i}]$ and $\mathbf{U}_{i}=[\mathbf{V}_{i}\boldsymbol{\beta}+\boldsymbol{\varepsilon}_{i},\mathbf{V}_{i}]$.
-Example 2.
-Distinct Correlated Effects (DCE).
-We can introduce distinct but correlated factors $\mathbf{F}=[\mathbf{F}_{\mathbf{x}},\mathbf{F}_{\mathbf{y}}]\in\mathbb{R}^{T\times(m_{x}+m_{y})}$ in Example 1, which makes the setup mode applicable: $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}_{\mathbf{y}}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$, $\mathbf{X}_{i}=\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$ as considered in Cui et al., (2022) or Stauskas and De Vos, (2025).
-Here, (1.1) nests such model similarly to Example 1, but $\boldsymbol{\gamma}_{i}$ is modified accordingly, depending on whether $\mathbf{F}_{\mathbf{y}}\subset\mathbf{F}_{\mathbf{x}}$, vice versa or $\mathbf{F}_{\mathbf{x}}\cap\mathbf{F}_{\mathbf{y}}=\emptyset$.
-Example 3.
-Block Structure (FRED-MD-QD , EA-MD-QD).
-The regressors in Examples 1 and 2 can be high-dimensional, but follow a natural structure of $k$ blocks and $m$ underlying factors.
-Let $i_{j}$ index series in block $j$ for $j=1,\ldots,k$.
-Then $\mathbf{Z}_{i}=\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}=[\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{k}}]^{\prime}$.
-Here, $N$ is the number of series per block.
-Monetary, industrial or fiscal blocks of series can be found in FRED-MD (see McCracken and Ng,, 2016), FRED-QD (see McCracken and Ng,, 2020) or EA-MD-QD (see Barigozzi et al.,, 2024) datasets.
-These datasets have been very popular in factor-augmented forecasts (see e.g.
-Karabiyik and Westerlund,, 2021, or Morico and Stauskas,, 2025).
-In all the cases above, we can estimate the factors by taking $K$ cross-section averages (CAs): $\widehat{\mathbf{F}}=\overline{\mathbf{Z}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g.
+where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is an error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ are latent common factors, where $m$ denotes the number of factors, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix.
+The most popular setting nested in (1.1) is that of interactive effects.
+Then $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, and $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ gives a model with unobserved heterogeneity, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables.
+Naturally, $K=k+1$, and $i$ indexes individuals, such as countries or households.
+The unobserved component of $\mathbf{y}_{i}$ is likely to be correlated with $\mathbf{X}_{i}$ (see Bai, 2009a, ).
+If $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, (1.1) represents the Common Correlated Effects (CCE) setting of Pesaran, (2006), where $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter of interest, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings).
+Since both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, we have endogeneity and cross-section dependence among individuals that make the least squares (LS) estimator of $\boldsymbol{\beta}$ inconsistent.
+It is possible to estimate unobserved factors in several ways, including the Principal Components (PC) method (see Bai and Ng,, 2006) or diversified projections (see Fan and Liao,, 2022).
+Pesaran, (2006) suggests a simple and elegant way to proxy $\mathbf{F}$ up to a linear transformation by taking cross-section averages (CAs): $\overline{\mathbf{Z}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]=\widehat{\mathbf{F}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g.
 Pesaran and Tosetti,, 2011).
-In Example 1, $\widehat{\mathbf{F}}$ is used as an additional regressor to merely control for the factors, and subsequently $\boldsymbol{\beta}$ is estimated with the least squares (LS), which constitutes the typical CCE estimator (see e.g.
-Karabiyik et al.,, 2017).
-In Example 2, only $\widehat{\mathbf{F}}_{\mathbf{x}}=\overline{\mathbf{X}}$ is used as a regressor, while $\mathbf{F}_{\mathbf{y}}$ is controlled by bootstrap (see Stauskas and De Vos,, 2025).
-In Example 3, the coefficient of $\widehat{\mathbf{F}}$ itself matters for forecast evaluation (see e.g.
-Massacci and Kapetanios,, 2024).
+The CCE estimator is then given by
+$\widehat{\boldsymbol{\beta}}_{CCE}=\overline{\mathbf{Q}}^{-1}\overline{\mathbf{q}},$ (1.2)
+where $\overline{\mathbf{Q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{X}_{i}$, $\overline{\mathbf{q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{y}_{i}$, $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix and $\mathbf{A}^{+}$ is the Moore-Penrose inverse.
+The core principle of CCE is to project the variables onto the orthogonal complement of their CAs in order to purge the factors asymptotically.
+Note that CAs are applicable beyond the usual interactive effects and CCE.
+Stauskas and De Vos, (2025) consider Distinct Correlated Effects (DCE), where $\mathbf{y}_{i}$ and $\mathbf{X}_{i}$ load on distinct sets of factors, and (1.2) is aided by bootstrap.
+Massacci and Kapetanios, (2024) use $\overline{\mathbf{X}}$ to approximate factors and produce factor-augmented forecasts.
 A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties.
@@
 Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023).
-This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs ($K\in\{k,k+1\}$ in Examples 1-3).
-However, there are reasons to ensure that $m=K$.
-If $m<K$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017).
+This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs.
+However, there are reasons to ensure that $m=k+1$.
+If $m<k+1$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017).
 This situation is common in macro datasets (see De Vos and Stauskas,, 2024).
-In DCE, $\mathbf{F}_{\mathbf{x}}$ estimated with an excess error makes bootstrap inconsistent (see Stauskas and De Vos,, 2025).
-In Example 3, popular tests of forecast accuracy comparison have highly non-standard asymptotic distributions (see Stauskas and Westerlund,, 2022).
 In general, too many CAs might even lead to a model that over-fits the data.
 In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion).
-In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that Information Criteria (IC) of Bai and Ng, (2002) from Principal Components (PC) literature can be applied in the CCE setting in Example 1 under stationary factors to identify an optimal set of CAs.
-De Vos and Stauskas, (2024) (DVS) showed similar results applicable to more general Examples 2 and 3.111Juodis, (2022) considers eigenvalue-based methods of Ahn and Horenstein, (2013)) under stationarity.
-The focus there is the number of factors but not the optimal set of CAs.
-We leave analysis of this approach for the future research.
+Beyond CCE, Stauskas and De Vos, (2025) show that DCE only uses $\overline{\mathbf{X}}$, but $m=k$ must hold.
+In forecasting, the same is needed to avoid only conservative confidence intervals around factor-augmented forecasts (see Karabiyik and Westerlund,, 2021).
+In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that the information criteria (ICs) of Bai and Ng, (2002) and Bai, 2009b from the PC literature can be applied in the CCE setting under stationary factors to identify an optimal set of CAs from $\overline{\mathbf{Z}}$.
+De Vos and Stauskas, (2024) (DVS) show similar results when CAs are selected from $\overline{\mathbf{X}}$, which is more applicable in the DCE and forecasting settings.
+[^1] Even in basic CCE setting, it is advised to omit $\overline{\mathbf{y}}$ and use $\overline{\mathbf{X}}$ only to avoid computational issues (see Karavias et al.,, 2023)., [^2] PC techniques are limited in CCE, as they focus solely on $m$ by detecting the largest eigenvalues of the data matrix.
+They cannot detect sets of CAs as they do not have a natural ordering.
+With IC, we learn $m$ from the cardinality of the selected set.
 MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC.
-The power of CAs to proxy general factor structures makes it natural to evaluate this statement and re-visit IC proposed in MW and DVS.
-We show that, contrary to the sentiment in the CCE/CAs literature, IC crumble under very mild forms of nostationarity.
-We demonstrate formally and in simulations that they exhibit a severe under-selection problem, which is independent of penalties used in the objective functions of IC.
-Moreover, nonstationarity of $\mathbf{F}$ cannot be accommodated in the objective functions a priori, as their very design is based on the idea that properties of factors and idiosyncratics should be similar.
-This study can be seen as a caution to practitioners to transform data and not to extrapolate attractive properties of CAs should they need to choose a number thereof.
+As CAs proxy a general factor structure, it is natural to evaluate this statement and re-visit the ICs proposed in MW and DVS.
+Therefore, this study can be seen as the CAs counterpart of Bai, (2004), where classical PC results of Bai and Ng, (2002) are evaluated against pure unit root factors.
+Instead, we use a mildly integrated process by Magdalinos and Phillips, (2009) to model $\mathbf{F}$, which allows us to experiment with varying degrees of factor persistence.
+We demonstrate formally and in simulations that, while ICs remain consistent as $(N,T)\to\infty$, highly persistent factors negatively affect their small sample performance.
+We discuss differences between our results and those in Bai, (2004), and also explore penalty functions adapted to non-stationary $\mathbf{F}$, as suggested in the latter study.
+Since they render our ICs inconsistent, we argue that practitioners should not mechanically take recommendations from the PC literature if the goal is to select CAs.
+More importantly, the ICs of DVS and MW (with the original penalties) should be applied in relatively large samples if the presence of non-stationary factors is suspected.
 2 Econometric Setup
-To further operationalize our examples, we introduce $M$ which is a set of column indices of $\overline{\mathbf{Z}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{K\times g}$ picks the corresponding $g$ averages in practice.
-For instance, $\overline{\mathbf{Z}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k+1$ CAs in the typical CCE setup.
-Moreover, let $\mathbf{q}_{M}=[\boldsymbol{0}_{g\times 1},\mathbf{q}_{M_{\mathbf{x}}}^{\prime}]^{\prime}$, then $\overline{\mathbf{Z}}\mathbf{q}_{M}=\overline{\mathbf{X}}\mathbf{q}_{M_{\mathbf{x}}}=\widehat{\mathbf{F}}_{M_{\mathbf{x}}}$ selects only from the set of $k$ regressors (or blocks), as required in Examples 2 and 3.222Throughout the remainder of the study, we use $m$ to denote the number of factors without the loss of generality.
-Clearly, this number can vary among our examples.
-Let accordingly $M_{0}$ ($M_{\mathbf{x},0}$) denote the set of averages from $\overline{\mathbf{Z}}$ ($\overline{\mathbf{X}}$) such that $\mathrm{rank}(\overline{\mathbf{C}}\mathbf{q}_{M_{0}})=\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x},0}})=m$, and $|M_{0}|=|M_{\mathbf{x},0}|=m$, where $|M|$ denotes cardinality of an arbitrary set $M$.
-IC under consideration are given by
-$\mathrm{IC}^{MW}(M)=\ln\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}\right)+g\cdot p_{N,T},$ (2.1)
-$\mathrm{IC}^{DVS}(M_{\mathbf{x}})=\ln\det\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}\right)+g\cdot k\cdot p_{N,T},$ (2.2)
-$p_{N,T}\in\left\{\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right)\right\}$ (2.3)
-where $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix onto the orthogonal complement of $\mathbf{A}$, $\mathbf{A}^{+}$ is a Moore-Penrose (MP) inverse, while $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ with $\widehat{\boldsymbol{\beta}}$ is obtained with the CCE estimator under all available $k+1$ CAs.
-Finally, $p_{N,T}$ is a penalty term which goes to 0 with $(N,T)\to\infty$.
-Examples of feasible and most popular penalties are given in (2.3) for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002).
-Note that (2.1) - (2.2) are functions of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
-Let $\overline{M}$ ($\overline{M}_{\mathbf{x}}$) be the set of indices of all available CAs that $|\overline{M}|=k+1$ ($|\overline{M}_{\mathbf{x}}|=k$). Then
-$\widehat{M}\hskip 2.84526pt(\widehat{M}_{\mathbf{x}})=\operatorname*{arg\,min}_{M\subseteq\overline{M}\hskip 2.84526pt(M_{\mathbf{x}}\subseteq\overline{M}_{\mathbf{x}})}\mathrm{IC}(M)\hskip 2.84526pt(\mathrm{IC}(M_{\mathbf{x}})),$ (2.4)
-where $|\widehat{M}|\hskip 2.84526pt(|\widehat{M}_{\mathbf{x}}|)=g$ provides the estimator of the number of factors for both criteria.
-In MW and DVS, it is demonstrated that $\mathbb{P}(\widehat{M}=M_{0})\to 1$ and $\mathbb{P}(\widehat{M}_{\mathbf{x}}=M_{\mathbf{x},0})\to 1$, respectively, as $(N,T)\to\infty$.
-These results theoretically justify the use of IC previously very popular in the PC literature to select the number of factors.
+We focus on the IC of DVS due to its wider applicability.
+While we provide numerical experiments on both DVS and MW, theoretical arguments regarding the latter are nearly identical, and we relegate them to Section 3.3 of the Supplementary material.
+To operationalize our analysis, we introduce $M$ which is a set of column indices of $\overline{\mathbf{X}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ picks the corresponding $g$ averages in practice.
+For instance, $\overline{\mathbf{X}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k$ CAs.
+Consequently, let $M_{0}$ denote the true set of averages from $\overline{\mathbf{X}}$ such that $\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=m$, and $|M_{0}|=m$, where $|M|$ denotes the cardinality of an arbitrary set $M$.
+IC under consideration is given by
+$\mathrm{IC}(M)=\ln\det\left(\overline{\mathbf{Q}}_{M}\right)+g\cdot k\cdot p_{N,T},$ (2.1)
+where $\ln(.)$ denotes the natural logarithm and $\mathrm{det}(\mathbf{A})$ is a determinant of any square matrix $\mathbf{A}$.
+$\overline{\mathbf{Q}}_{M}$ is $\overline{\mathbf{Q}}$ from (1.2) evaluated at $\widehat{\mathbf{F}}_{M}$, and $p_{N,T}$ is a penalty term.
+[^3] In case of MW, $\overline{\mathbf{Q}}_{M}=\ln\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}\right)$, where $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ and $\widehat{\boldsymbol{\beta}}$ is obtained using the CCE estimator under all available $k+1$ CAs.
+Examples of feasible and most popular penalties are given by
+$p_{N,T,1}=\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\hskip 5.69054ptp_{N,T,2}=\frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right)$ (2.2)
+for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002).
+Note that (2.1) is function of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
+Let $\overline{M}$ be the set of indices of all available CAs that, so that $|\overline{M}|=k$. Then
+$\widehat{M}\hskip 2.84526pt=\operatorname*{arg\,min}_{M\subseteq\overline{M}\hskip 2.84526pt}\mathrm{IC}(M),$ (2.3)
+where $|\widehat{M}|=g$ provides the estimator of the number of factors as a consequence.
+In DVS, it is demonstrated that $\mathbb{P}(\widehat{M}=M_{0})\to 1$ as $(N,T)\to\infty$.
+This result theoretically justifies the use of IC previously very popular in the PC literature to select the number of factors.
 These findings are based on stationary $\mathbf{F}$, which has been considerably relaxed in the CCE/CAs literature.
-Therefore, our natural goal is to examine (2.1) and (2.2) under departures from this restrictive assumption.
+Therefore, our natural goal is to examine (2.1) under departures from this restrictive assumption.
 Throughout our analysis, we employ the following set of assumptions.
-Assumption 1.
+Assumption 1. $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009).
   • (a)
-    $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009). In particular,
-    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}+\frac{\mathbf{Q}}{T^{\tau}},\quad\mathbf{Q}<0\hskip 5.69054pt(diagonal),$
+    In particular,
+    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}>0\hskip 5.69054pt(diagonal),$
     where $\mathbf{u}_{f,t}$ is a zero-mean linear process.
   • (b)
-    If $\tau=0$, then $|\lambda_{\mathrm{min}}(\mathbf{Q})|<2$, where $\lambda_{\mathrm{min}}$ is the minimum eigenvalue, in order to ensure that $\mathbf{R}_{f,T}=\mathbf{I}_{m}+\mathbf{Q}$ is inside the unit circle, and $\mathbf{u}_{f,t}$ is such that $\{\mathbf{f}_{t}\}$ has absolute summable autocovariances.
+    If $\tau=0$, then $\{\mathbf{f}_{t}\}$ is a stationary process with summable autocovariances.
 Assumption 2. Let $\mathbf{e}_{i,t}=(\varepsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}\in\mathbb{R}^{k+1}$. Then
@@
     • (i)
-      If $\tau\in(0,1)$, then $\{\mathbf{e}_{i,t}\}$ is uncorrelated over time with $\mathbb{E}(\mathbf{e}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\mathbf{e}_{i,t}\mathbf{e}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{ee,i,t}$ positive definite and $\mathbb{E}(\|\mathbf{e}_{i,t}\|^{4})<\infty$.
+      If $\tau\in(0,1)$, then $\{\mathbf{e}_{i,t}\}$ is a martingale difference sequence, $\mathbb{E}(\mathbf{e}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\mathbf{e}_{i,t}\mathbf{e}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{ee,i,t}$, $\lim_{N,T\to\infty}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\boldsymbol{\Sigma}_{ee,i,t}=\boldsymbol{\Sigma}_{ee}=\mathrm{diag}(\sigma^{2},\boldsymbol{\Sigma}_{\mathbf{v}})$ positive definite and $\mathbb{E}(\|\mathbf{e}_{i,t}\|^{4})<\infty$.
     • (ii)
       If $\tau=0$, then we let $\mathbf{e}_{i,t}=\mathbf{K}_{i}(L)\boldsymbol{\epsilon}_{i,t}=\sum_{j=0}^{\infty}\mathbf{K}_{i,j}\boldsymbol{\epsilon}_{i,t-j}$, where $\boldsymbol{\epsilon}_{i,t}$ is independent across $t$ with $\mathbb{E}(\boldsymbol{\epsilon}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\boldsymbol{\epsilon}_{i,t}\boldsymbol{\epsilon}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{\epsilon\epsilon,i,t}$ positive definite, $\mathbb{E}(\|\boldsymbol{\epsilon}_{i,t}\|^{4})<\infty$, and $\sum_{j=0}^{\infty}j^{1/2}\|\mathbf{K}_{i,j}\|<\infty$.
-Also, $\boldsymbol{\xi}_{t}=\mathrm{vec}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}-\mathbb{E}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right)\right)$ is strong mixing with coefficients of size $-bd(b-d)$ with $b>4$ and $b>d>2$, $\mathbb{E}(\|\boldsymbol{\xi}_{t}\|^{b})<\infty$, and $\lim_{T\to\infty}T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{s}^{\prime})$ is positive definite.
+Also, $\boldsymbol{\xi}_{t}=\mathrm{vec}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}-\mathbb{E}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right)\right)$ is strong mixing with coefficients of size bd(bd)-bd(b-d) with $b>4$ and $b>d>2$, $\mathbb{E}(\|\boldsymbol{\xi}_{t}\|^{b})<\infty$, and $\lim_{T\to\infty}T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{s}^{\prime})$ is positive definite.
 Here and throughout, $\left\|\mathbf{A}\right\|=\sqrt{\mathrm{trace}(\mathbf{A}^{\prime}\mathbf{A})}$ is Frobenius norm.
@@
   • (b)
-    We have $\left\|\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\mathbf{e}_{i,t}\mathbf{e}_{j,t}^{\prime}-\boldsymbol{\Sigma}_{ee}\right\|=o_{p}(1)$ $N\to\infty$ for a positive definite matrix $\boldsymbol{\Sigma}_{ee}$.
+    We have $\left\|\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{e}_{i,t}\mathbf{e}_{j,t}^{\prime}\right)-\boldsymbol{\Sigma}_{ee}\right\|=o(1)$ as $(N,T)\to\infty$ for a positive definite matrix $\boldsymbol{\Sigma}_{ee}$.
 Assumption 3. $\mathbf{f}_{t}$ and $\mathbf{e}_{i,s}$ are independent for all $t$, $s$ and $i$.
 Assumption 4.
-$\mathbf{C}_{i}$ is a matrix, such that $\mathbf{C}_{i}=\mathbf{C}+\boldsymbol{\eta}_{i},\mathrm{vec}(\boldsymbol{\eta}_{i})\sim IID(\boldsymbol{0}_{Km\times 1},\boldsymbol{\Omega})$ for positive semi-definite $\boldsymbol{\Omega}$.
-Also, $\overline{\mathbf{C}}=[\overline{\mathbf{C}}_{m},\overline{\mathbf{C}}_{-m}]$, where $\overline{\mathbf{C}}_{-m}\in\mathbb{R}^{m\times(K-m)}$ and $\overline{\mathbf{C}}_{m}\in\mathbb{R}^{m\times m}$ is full rank for all $N$, including $N\to\infty$, and $\|\mathbf{C}_{i}\|<\infty$.
-If $m=K$, then $\overline{\mathbf{C}}=\overline{\mathbf{C}}_{m}$.
-Assumption 1 (a) treats factors very flexibly and allows us perform comparative statics, as $\tau\to 1$ increases persistence of the process.
-We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allow us vary $\tau$ and give theoretical guarantees, such as $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$, which is a constant positive definite matrix.
+$\mathbf{C}_{i}$ is a deterministic matrix, such that $\|\mathbf{C}_{i}\|<\infty$ and $\frac{1}{N}\sum_{i=1}^{N}\mathbf{C}_{i}\mathbf{C}_{i}^{\prime}\to\boldsymbol{\Sigma}_{\mathbf{C}}$ positive definite.
+Also, $\overline{\mathbf{C}}=[\overline{\mathbf{C}}_{m},\overline{\mathbf{C}}_{-m}]$, where $\overline{\mathbf{C}}_{-m}\in\mathbb{R}^{m\times(k+1-m)}$ and $\overline{\mathbf{C}}_{m}=\overline{\mathbf{C}}\mathbf{q}_{M_{0}}\in\mathbb{R}^{m\times m}$ for a unique $M_{0}$ is full rank for all $N$, including $N\to\infty$.
+If $m=k+1$, then $\overline{\mathbf{C}}=\overline{\mathbf{C}}_{m}$.
+Assumption 1 (a) treats the factors very flexibly and offers comparative statics, as $\tau\to 1$ increases persistence of the process.
+We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allow us to vary $\tau$ and give theoretical guarantees, such as $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$, which is a constant positive definite matrix.
 Part (b) ensures that we return to the usual stationarity conditions under $\tau=0$.
+Specifically, we need the eigenvalues of $\mathbf{I}_{m}-\mathbf{G}$ to be within a unit circle, so that the process can be inverted and admit a representation of $MA(\infty)$.
 Assumption 2 (a) is split into two parts.
-If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}$ converges to zero too slow for our purposes, which would dominate and obscure the very effect of factor persistence on IC, and we would like to avoid this (see Lemma 3.1 in Magdalinos and Phillips,, 2009).
-We present simulations under serial correlation in the Supplement, which does not demonstrate a major negative effect on IC.
-The rest of the assumptions are standard, as they ensure mutual independence of factors and the error terms (Assumption 3), weak cross-section dependence of idiosyncratics (Assumption 2 (b)) and informativeness of the loadings (rank condition in Assumption 4).
-The loadings are random as in e.g.
-Pesaran, (2006), but they can also be treated as deterministic parameters.
-To formulate our key theoretical result, let us denote the objective function component in (2.1) and (2.2) evaluated at some set of CAs $M$ by $\mathcal{V}_{j}$ for $j\in\{MW,DVS\}$.
-Similarly, let $\mathcal{V}_{j}^{0}$ be the same component evaluated at $M_{0}$.
-Proposition 1 below characterizes the asymptotic behavior of $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}$ under $M_{0}\subset M$ (over-specification) and $M\subset M_{0}$ (under-specification).333 Strictly speaking, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$.
+If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, we have that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-\tau/2})$ (see Lemma 3.1 in Magdalinos and Phillips,, 2009), which is too slow to demonstrate that the factor estimation error is negligible.
+In order not to obscure the very effect of factor persistence on IC, we restrict the correlation of idiosyncratics and show that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$ as needed (see our auxiliary results in the Supplementary material, and a comparative rate requirement when $\tau=0$ in e.g.
+Karabiyik et al.,, 2017).
+We present simulations under serial correlation in the Supplement, which do not show any negative effect on IC.
+The rest of the assumptions are standard, as they ensure weak cross-section dependence of the idiosyncratics (Assumption 2 (b)), mutual independence of factors and the error terms (Assumption 3; see also Pesaran,, 2006), and informativeness of the loadings (rank condition in Assumption 4).
+The loadings are deterministic, but they can also admit a random coefficient model (see e.g.
+De Vos and Everaert,, 2021).
+In our assumptions, we cover $k+1$ variables in order to accommodate MW, as well.
+To formulate our instrumental theoretical result, let $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\mathrm{det(\overline{\mathbf{Q}}_{M})}-\ln\mathrm{det(\overline{\mathbf{Q}}_{M_{0}})}$.
+This quantity is important because it allows us to track the asymptotic contribution of factor estimates to the minimization problem under $M_{0}\subset M$ (over-specification) and $M\subset M_{0}$ (under-specification).
+Crucially, the behavior of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ prescribes properties of $p_{N,T}$.
+To illustrate this, Bai, (2004) (in PC setting) shows that under over-specification, (an equivalent of) $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is $O_{p}(1)$ when factors are non-stationary, and so the contribution of the extra estimates to the sum of squared residuals is non-negligible.
+This leads to $p_{N,T}\to\infty$ to heavily penalize redundant estimates.
+However, when factors are stationary, Bai and Ng, (2002) (and DVS) detect $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$, which means that $p_{N,T}\to 0$ at a slower rate in order to penalize lightly.
+Proposition 1 below can be seen as the CAs equivalent of Lemma A3 and Lemma A4 of Bai, (2004).444 Strictly speaking, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$.
 We only analyze the case of $M\subset M_{0}$ for brevity.
 As Margaritella and Westerlund, (2023) pointed out analysis would lead to the same conclusion in all the cases.
-Proposition 1. Under Assumptions 1-4 as $(N,T)\to\infty$, we have that
-$\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=\begin{cases}O_{p}(T^{\tau})\hskip 5.69054pt\textit{under}\hskip 5.69054ptM\subset M_{0},\\ O_{p}(C_{N,T}^{-2})\hskip 5.69054pt\textit{under}\hskip 5.69054ptM_{0}\subset M.\end{cases}$ (2.5)
-Let us begin from the over-specification.
+Proposition 1. Under Assumptions 1-4 as $(N,T)\to\infty$, we have that:
+  • (a)
+    Under $M\subset M_{0}$, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}+\infty$;
+  • (b)
+    Under $M_{0}\subset M$, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$.
+Proof: Section 3 of Supplementary material.
+Let us discuss Proposition 1 and start with (b).
 In this case, the CAs are able to approximate all $m$ factors under Assumption 4.
-Also, for illustration purposes let $j=DVS$.
-Then $\mathcal{V}_{DVS}$ (and $\mathcal{V}_{DVS}^{0})$ is exactly (log-determinant of) the denominator of the CCE estimator.
-It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x}}})^{+}+o_{p}(1)$, and by inserting this into $\mathbf{X}_{i}$, we can show that
-$\mathcal{V}_{DVS}=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}\right)+o_{p}(1)=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}\right)+O_{p}(T^{-1}),$ (2.6)
-whose argument is consistent for a positive definite matrix under general factors (see Westerlund,, 2018), and a similar logic applies to the objective function of MW.
-Thus, it is natural that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}$ is negligible.
-Moreover, the rate is already detected in Bai and Ng, (2002) under stationary $\mathbf{F}$ for PC, and it remains identical here.
+Then $\overline{\mathbf{Q}}_{M}$ is exactly (log-determinant of) the denominator of the CCE estimator.
+It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}+o_{p}(1)$ for fixed $T$, and by inserting this into $\mathbf{X}_{i}$, we can show that
+$\ln\mathrm{det}(\overline{\mathbf{Q}}_{M})=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}\right)+o_{p}(1)=\ln\mathrm{det}\left(\boldsymbol{\Sigma}_{\mathbf{v}}\right)+o_{p}(1),$ (2.4)
+whose argument is a positive definite matrix under general factors (see also Westerlund,, 2018).
+Given that $\ln\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}})$ admits the same asymptotic representation, it is natural that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is negligible.
+Moreover, the rate is already detected by DVS and Bai and Ng, (2002) (for PC) under stationary $\mathbf{F}$, and it remains identical here.
 This rate is instructive and determines the properties of penalty $p_{N,T}$, which is the key motivation behind the functional forms in (2.2).
-Indeed, if $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$ as $(N,T)\to\infty$, Proposition 1 is the key ingredient to show that $\mathbb{P}(\mathrm{IC}^{DVS}(M_{\mathbf{x}})-\mathrm{IC}^{DVS}(M_{\mathbf{x},0})>0)\to 1$, and the same statement holds for MW.
-This means that the probability that IC are minimized at some other set containing $M_{0}$ goes to 0.
-In general, under $M_{0}\subset M$, the usual CCE logic prevails deceptively.
-Under $M\subset M_{0}$, the rank condition in Assumption 4 is not satisfied.
-This means that $m-g$ factors are not asymptotically projected out, and the representation in (2.6) is invalid.
-Given that $\mathcal{V}_{j}$ is a quadratic form for both $j\in\{MW,DVS\}$, their second empirical moment is of the order $O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a).
-However, the objective function offers normalization by $T$ only.
-Exactly this imbalance induces divergence, and this implies that the probability that IC are minimized at some other set contained in $M_{0}$ does not necessarily go to 0.
-Under stationary $\mathbf{F}$, however, it is clear that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=O_{p}(1)$ and, in fact, $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}\to_{p}c>0$ as $(N,T)\to\infty$ as required (see the analysis of MW, DVS and Bai and Ng,, 2002).
-Our Proposition 1 naturally provides the same result under $\tau=0$.
-Note that the failure of both IC is purely driven by the objective function, and it complicates the role of $p_{N,T}$.
-If $M_{0}\subset M$, $p_{N,T}$ should simply vanish slower than $C_{N,T}^{-2}$, but under $M\subset M_{0}$ the penalty has to counter divergence.
-This discussion suggests that accounting for non-stationarity of $\mathbf{F}$ in the objective function should solve the issue, but this is not the case.
-Assuming that $\tau$ is known, we normalize the objective function by $T^{1+\tau}$ as Assumption 1 (a) suggests.
-In the Supplement, we show that under $M\subset M_{0}$:
-$\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}=\ln\mathrm{det}\left[\mathbf{I}_{k}+\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x}}}}-\mathbf{M}_{\widehat{\mathbf{F}}_{\mathbf{M}_{\mathbf{x},0}}})\mathbf{X}_{i}}_{\mathbf{I}=O_{p}(1)}\left(\underbrace{\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}}_{\mathbf{II}=O_{p}(T^{-\tau})}\right)^{-1}\right]+o_{p}(1),$ (2.7)
-where the determinant does not converge to a strictly positive constant, again.
-The reason behind this is that the rank condition is satisfied in $\mathbf{II}$, and by following the steps of Karabiyik et al., (2017) and Westerlund, (2018), it admits (2.6).
-However, since idiosyncratics are stationary, “over-normalization” causes singularity, and we get a new imbalance between $\mathbf{I}$ and $\mathbf{II}$.
-Therefore, IC fundamentally depend not only on stationarity of factors, but also on stationarity of idiosyncratics.
-We confirm this with the additional simulations in the Supplement, where this adjustment becomes more harmful with $\tau$.
-If $M_{0}\subset M$, (2.5) and normalization by $T^{1+\tau}$ imply that (2.7) is still $\ln\det(\mathbf{I}_{k}+O_{p}(T^{-\tau}C_{N,T}^{-2})\times O_{p}(T^{\tau}))=O_{p}(C_{N,T}^{-2})$, which again means that $p_{N,T}$ needs to have different properties depending on whether $M_{0}\subset M$ or $M\subset M_{0}$.
+Specifically, because redundant gmg-m CAs have an asymptotically negligible contribution to the sum of squared residuals, over-specification does not need to be penalized heavy, and so it is sufficient that $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$ as $(N,T)\to\infty$.
+Note that this stands in sharp contrast to findings in Lemma A4 of Bai, (2004), where excess factor estimates contribute non-trivially.
+Overall, the usual consistency of the CCE estimator in (1.2) under general unknown factors prevails when $M_{0}\subset M$ (see also Westerlund,, 2018, or Stauskas,, 2023).
+Under $M\subset M_{0}$ in (a), the rank condition in Assumption 4 is not satisfied.
+This means that mgm-g factors are not asymptotically projected out, and the representation in (2.4) is invalid.
+Note that $\overline{\mathbf{Q}}_{M}$ is a quadratic form in $\mathbf{F}$, and we know that $\left\|\mathbf{F}^{\prime}\mathbf{F}\right\|=O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a).
+However, the objective function offers normalization by $T$ only as the integration order is unknown.
+Exactly this imbalance causes divergence of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$.
+A similar situation arises in Lemma A3 of Bai, (2004), where the central step is to detect the sign of divergence, and we follow this path.
+Note that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\det\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)$, where $\boldsymbol{\Omega}_{M,M_{0}}:=T^{-\tau}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}\geq 0$ (not zero) as $(N,T)\to\infty$, which we demonstrate in the Supplement.
+Using this, the difference can be simplified to
+$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\tau b\ln(T)+\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}})\right)+R_{N,T},$ (2.5)
+where $|R_{N,T}|=o_{p}(1)$ is the remainder, $\lambda_{j}(\mathbf{A})$ is the $j$-th eigenvalue of $\mathbf{A}$, while $b$ is a number of strictly positive eigenvalues.
+Because $b$ is finite ($b\leq k$), it is clear that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to+\infty$ at a rate of $O_{p}(\ln(T))$.
+Two comments are in order.
+Firstly, in the Supplement we show that $|R_{N,T}|=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$, and it is not necessarily positive.
+This means that for $\tau\approx 1$ and a small $N$, the remainder vanishes slowly as $T\to\infty$.
+Hence, it may be necessary to rely on large $N,T$ combinations to detect the minimum of IC outside of $M\subset M_{0}$ region if $\mathbf{F}$ is very persistent.
+We explore (and confirm) such risks in Monte Carlo simulations in Section 3.
+Secondly, while this result is similar to Lemma A3 in Bai, (2004), there the divergence rate is $O_{p}(T/\ln\ln(T))$, since the law of iterated logarithm is used to arrive at the expression analogous to (2.5) when factors have unit root.
+Theorem 1 leads to our consistency result.
+Theorem 1. Under conditions of Proposition 1 with $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$, we have
+$\mathbb{P}\left(\widehat{M}=M_{0}\right)\to 1$
+as $(N,T)\to\infty$.
+Proof. Follows directly from Proposition 1, because $\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})>0\right)\to 1$ under (a)-(b) and the penalty properties.
+The key message of Theorem 1 is that the IC of DVS (and MW) is consistent, but exhibits hybrid asymptotic properties characteristic to stationary and non-stationary factors.
+Consequently, two forces are in effect: $p_{N,T}$ should still be negligible, but the factor estimation error can be “large” if $M\subset M_{0}$ as indicated in (2.5).
+To balance these properties in practice, it may be tempting to further experiment with the penalty.
+In an analogous theorem of Bai, (2004), it is required that $p_{N,T}\to\infty$ and $p_{N,T}/\ln(T)\to 0$ if the criterion is based on a logarithmic-valued sum of squared residuals as is the case here (see footnote 3 in that study).
+Importantly, this condition does not interfere with consistency in our Theorem 1, because clearly $p_{N,T}=o(\ln(T))$ still holds.
+Therefore, we compare both types of penalties in the simulations to illustrate the outcome if practitioners simply extrapolate recommendations from the PC literature.
+We let $\widetilde{p}_{N,T}=\ln(T)p_{N,T}$, which follows the suggestion in (12) in Bai, (2004), and obeys $\widetilde{p}_{N,T}\to\infty$ and $\widetilde{p}_{N,T}/\ln(T)\to 0$.
 Remark 1.
-Throughout, we assume that $M_{0}$ ($M_{0,\mathbf{x}}$) is unique.
+Note that under stationarity, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}c>0$ when $M\subset M_{0}$, according to the results in DVS.
+Our Proposition 1 naturally accommodates this result under $\tau=0$, because, by following the same approximation steps, (2.5) becomes
+$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\sum_{j:\lambda>0}\ln\left(1+\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}})\right)+R_{N,T}\to_{p}c>0$
+as $(N,T)\to\infty$ because each summand is greater than 1 w.p.
+1.
+Also, $|R_{N,T}|=O_{p}(N^{-1})+O_{p}((NT)^{-1/2})$, which is the same as in DVS under $\tau=0$, as expected from the discussion of (2.5).
+Remark 2.
+Throughout, we assume that $M_{0}$ is unique.
 However, because CAs do not have a natural ordering, the problem is only set-identified.
-It can be shown that both IC select the set that minimizes the asymptotic mean squared error of all the sets that satisfy Assumption 4 (see e.g.
+It can be shown that both ICs select the set that minimizes the asymptotic mean squared error out of all the sets that satisfy Assumption 4 (see e.g.
 Corollary 3.1 in Margaritella and Westerlund,, 2023).
 In this study, the first order issue is ability of IC to select any $\widehat{M}$ with the cardinality of $m$.
+Remark 3.
+Juodis, (2022) considers eigenvalue-based method of Ahn and Horenstein, (2013) under stationarity, where the focus is learning $m$, but not the set of CAs.
+We leave analysis of this approach for the future research, but provide comparative simulation evidence under mildly integrated $\mathbf{F}$ in the Supplementary material.
 3 Simulation Study
@@
 In the case of time- and space-independent errors, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$.
-In the case of errors correlated weakly across time and cross-sections we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
-[…]
+In the case of errors correlated weakly across time and cross-section we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
+$\mathbf{e}_{i,t}=(\epsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}=\mathbf{P}_{\rho}\mathbf{e}_{i,t-1}+\mathbf{z}_{\mathbf{w},i,t},\hskip 5.69054pt\mathbf{P}_{\rho}=\begin{bmatrix}\rho&\mathbf{0}_{1\times k}\\ \mathbf{0}_{k\times 1}&\rho_{v}\mathbf{I}_{k}\end{bmatrix},\hskip 5.69054pt\mathbf{z}_{\mathbf{w},i,t}=\begin{bmatrix}\sqrt{\frac{m\left(1-\rho^{2}\right)}{1+2J\kappa^{2}}}\boldsymbol{\epsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa\mathbf{w}_{i}^{\prime}\right)\\ \sqrt{\frac{\left(1-\rho_{v}^{2}\right)}{1+2J_{v}\kappa_{v}^{2}}}\boldsymbol{\Upsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa_{v}\boldsymbol{\tilde{\mathbf{w}}}_{i}^{\prime}\right)\end{bmatrix}$ (3.2)
 where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero.
@@
 The factors are drawn as $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t}$, where
-$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{Q}T^{-\tau},\quad\mathbf{Q}=diag(q_{1},...,q_{m}),q_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.3)
+$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=diag(g_{1},...,g_{m}),g_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.3)
 In the stationary factors case, the factors are drawn as a stable VAR with $\tau=0$.
-Two cases for non-stationary factors are considered: the near stationary case of $\tau=0.1$ and the case close to local-to-unity process $\tau=0.9$.
-All simulations have 1000 repetitions.
+Two cases for non-stationary factors are considered: the non-stationary case of $\tau=0.4$ and the case close to local-to-unity process $\tau=0.9$.
+All simulations have 1000 repetitions and we compare DVO and MW.
 In the Supplement, we provide additional simulation evidence of the eigenvalue growth ratio (ER) based on Juodis, (2022) for comparison.
 The focus of this method is the number of factors ($m$), and it does not select the sufficient set of CAs.
-Nevertheless, its performance crumbles with $\tau\to 1$, as well.
-Moreover, we present extended tables with extra $N,T=150,200$ for all our experiments.
-Table 6.6.5 below presents the average of the number of CAs selected for the case of stationary factors ($\tau=0$), the near stationary case and the close to local-to-unity process.
-The idiosynchratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units.
+Nevertheless, its performance crumbles with $\tau\to 1$ and does not improve as $(N,T)\to\infty$.
+Table 3 below presents the frequency of the correctly selected number of CAs, $g=4$, with the default penalties, $p_{N,T,1}$ and $p_{N,T,2}$ from 2.2, and the penalty from Bai, (2004), $\widetilde{p}_{N,T}$, for the case of stationary factors ($\tau=0$), the mildly non-stationary case ($\tau=0.4$) and the close to local-to-unity process.
+The idiosyncratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units.
 The Supplement presents additional results with weak correlation over time.
-We start with the in the literature established case of stationary factors.
-In the small $T$ cases, $T=10$ and $T=20$, we first observe an overselection of the number of CAs, followed by an underselection.
-This pattern resembles the results in MW and is consistent across all values for $N$ and $\tau$ [^4] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$ If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors.
-Once the number of cross-sectional units and time periods passes 50, all four criteria select close to or exactly the correct number of factors.
-Moving to the near-stationary case, the pattern remains the same.
-However, once moving to the near unit root process, we observe for all four criteria, overselection of the number of CAs for the small time series case of $T=10$, followed by underselection if $T=20$.
-The underselection persists even if $N$ and $T$ increase to 100.
-This showcases the collapse of all information criteria in the case of non-stationary factors.
+We start with case of stationary factors established in the literature and the default penalties.
+In the small $T$ and $N$ case, $T=N=50$, and for stationary factors, we observe a frequency between $16\%$ and $36\%$, a pattern which resembles the results in MW.
+[^5] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$.
+If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors.
+The selection frequency decreases to zero in the case of non-stationary factors.
+As expected, consistency, illustrated here by a selection frequency of 100%, is quickly achieved when $N$ and $T$ increase.
+In the non-stationary cases, consistency requires much larger combinations of $(N,T)$.
+For the extreme case of a process close to local-to-unit, combinations of $(N,T)>300$ are required.
+Bai, (2004) proposes to adjust the penalty by $\ln(T)$ in the presence of non-stationary factors.
+The lower panel of Table 3 presents the frequency of the correctly selected number of CAs with a penalty of $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(\frac{NT}{N+T})$ for $\mathrm{IC}_{1}$ and $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(C^{2}_{N,T})$ for $\mathrm{IC}_{2}$.
+We observe for all four ICs a selection frequency of 0 for almost all combinations of $N$ and $T$ and $\tau$.
+Only in the large sample cases of $N=500$ and $T>300$ and only under stationary factors the selection frequency reaches 100%.
+This strongly suggests that prescriptions from the PC literature do not directly apply to CAs, as ICs are clearly inconsistent, and over-selection should not be heavily penalized.
+Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$ (”small sample”) and $N=T=500$ (”large sample”), where we increase $\tau$ in $0.01$ steps from 0 to $0.95$.
+As in Table 3 the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ are uncorrelated over time, but weakly correlated across units.
+For the small sample, the share of misselected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint.
+Then, it increases to almost $100\%$ with swings showcasing that the criteria become more unstable as $\tau$ increases.
+As expected, the degree of non-stationarity is much less of a problem for the large sample, and we observe misselection and instability of the criteria only for a high degree of non-stationarity.
+To shed more light on the importance of the sample size, we present in Figure 2 the share of misselected number of CAs with varying number of time periods from 25 to 500, $\tau=0.9$ and $N=[100,500]$.
+The share of misselection remains high for $N=100$ for both ICs, but declines for $N=500$.
+Again, this presents further evidence of the consistency of the ICs, but large samples are required to use them effectively.
 Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
-$\tau=0$ $\tau=0.1$ $\tau=0.9$
-$N$ $T$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$
-10 10 9.00 8.58 5.12 1.00 8.99 8.65 4.92 1.00 8.99 8.51 2.11 1.02
-20 10 8.99 8.78 7.98 3.18 8.98 8.78 7.94 3.19 8.95 8.31 7.12 1.52
-50 10 8.99 8.99 7.94 6.83 9.00 8.98 7.94 6.68 8.96 8.87 5.97 2.47
-100 10 9.00 9.00 7.93 7.52 9.00 9.00 7.97 7.69 8.98 8.94 5.42 3.42
-10 20 1.15 1.02 1.00 1.00 1.19 1.03 1.01 1.00 1.21 1.12 1.05 1.01
-20 20 1.23 1.01 1.93 1.02 1.27 1.02 1.97 1.04 1.33 1.16 1.37 1.13
-50 20 2.68 2.04 2.02 1.24 2.72 2.10 2.08 1.33 1.45 1.35 1.32 1.20
-100 20 3.79 3.57 3.29 2.67 3.58 3.40 2.98 2.45 1.51 1.46 1.35 1.30
-10 50 1.21 1.13 1.05 1.03 1.11 1.05 1.00 1.00 1.25 1.21 1.14 1.09
-20 50 1.09 1.03 1.21 1.01 1.08 1.02 1.29 1.01 1.29 1.23 1.32 1.20
-50 50 3.60 2.80 3.91 2.45 3.90 3.24 3.97 2.94 1.95 1.68 1.92 1.55
-100 50 3.33 3.21 3.30 2.95 3.19 3.06 3.08 2.71 1.59 1.51 1.59 1.48
-10 100 1.00 1.00 1.26 1.18 1.00 1.00 1.28 1.22 1.15 1.14 1.04 1.03
-20 100 1.66 1.36 2.50 2.15 1.76 1.43 2.61 2.41 1.43 1.36 1.66 1.58
-50 100 4.00 3.99 4.00 4.00 4.00 3.98 4.00 3.99 2.11 1.97 1.97 1.80
-100 100 4.01 4.00 4.00 4.00 4.01 4.00 4.00 4.00 2.25 2.04 2.17 1.86
-Table 1: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.2) and (2.1).
-For $IC^{DVS}_{1}$ and $IC^{MW}_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $IC^{DVS}_{2}$ and $IC^{MW}_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$
+Penalty $p$
+50 50 31.00 15.90 36.40 16.10 7.30 3.60 3.50 0.30 0.10 0.00 0.00 0.00
+100 50 97.30 95.00 99.20 96.40 91.50 82.80 95.10 85.90 11.50 8.60 13.40 7.50
+200 50 95.50 94.30 99.70 98.80 48.60 45.50 77.70 69.20 0.20 0.20 5.50 4.10
+300 50 100.00 100.00 99.60 99.50 81.30 78.20 69.90 67.70 4.30 3.70 1.80 1.60
+500 50 100.00 100.00 100.00 100.00 99.60 99.40 96.30 95.10 21.20 20.10 11.50 10.10
+50 100 97.10 89.30 100.00 100.00 68.60 55.30 99.30 94.50 7.00 5.90 18.20 10.90
+100 100 100.00 99.90 99.90 99.00 62.50 49.40 56.30 39.70 2.70 1.30 1.70 0.60
+200 100 100.00 100.00 100.00 100.00 99.90 99.60 99.90 99.70 24.50 19.70 26.30 18.80
+300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 39.10 34.30 46.60 39.30
+500 100 100.00 100.00 100.00 100.00 99.70 99.70 95.90 94.80 16.30 15.30 8.50 7.60
+50 200 99.70 99.90 99.90 99.80 97.70 96.70 89.00 79.90 9.80 8.70 5.10 3.50
+100 200 100.00 100.00 100.00 100.00 93.00 88.80 99.50 98.60 5.30 4.70 12.80 7.70
+200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 40.80 32.70 60.20 44.50
+300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 22.40 18.90 81.40 73.20
+500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 54.10 49.30 80.80 75.50
+50 300 100.00 100.00 100.00 100.00 88.60 86.40 99.90 99.40 7.40 6.20 15.20 11.00
+100 300 98.90 99.80 100.00 100.00 98.70 99.40 100.00 100.00 21.30 18.60 27.20 22.40
+200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 56.00 49.10 52.60 45.60
+300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 89.30 82.90 94.40 89.10
+500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 80.10 75.70 93.70 91.90
+50 500 100.00 100.00 100.00 100.00 90.10 89.00 100.00 100.00 3.40 3.10 11.20 10.20
+100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 38.50 37.10 53.70 50.80
+200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 84.10 81.30 97.30 95.80
+300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 88.30 85.90 97.00 95.60
+500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 92.70 88.20 99.50 98.70
+Penalty by Bai, (2004), $\widetilde{p}=\ln(T)p$
+50 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+300 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+500 50 0.00 0.00 0.00 0.00 0.10 0.10 0.00 0.00 0.00 0.00 0.00 0.00
+50 100 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 100 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 100 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+300 100 0.00 0.00 0.00 0.00 0.10 0.10 0.00 0.00 0.00 0.00 0.00 0.00
+500 100 0.10 0.00 0.00 0.00 1.20 0.40 0.00 0.00 0.10 0.10 0.00 0.00
+50 200 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 200 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 200 0.00 0.00 0.00 0.00 0.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+300 200 2.40 0.00 0.00 0.00 8.80 2.70 0.00 0.00 0.10 0.00 0.00 0.00
+500 200 26.10 4.70 0.00 0.00 12.90 7.20 0.00 0.00 0.40 0.10 0.00 0.00
+50 300 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 300 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 300 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.10 0.10 0.00 0.00
+300 300 19.30 0.00 0.00 0.00 16.00 1.90 0.00 0.00 0.90 0.70 0.00 0.00
+500 300 99.90 92.90 0.00 0.00 72.70 53.60 1.20 0.00 2.70 1.30 0.00 0.00
+50 500 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 500 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 500 0.00 0.00 0.00 0.00 0.30 0.00 0.00 0.00 0.80 0.50 0.00 0.00
+300 500 15.80 0.10 0.00 0.00 14.90 3.10 0.00 0.00 0.90 0.70 0.00 0.00
+500 500 100.00 100.00 77.40 0.20 94.50 81.20 31.50 2.00 2.50 1.70 0.30 0.00
+Table 1: Correct Selection Frequency for $g$ with $m=4$ and $K\in\{8,9\}$.
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1).
+In the upper part, for $IC^{DVS}_{1}$ and $IC^{MW}_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $IC^{DVS}_{2}$ and $IC^{MW}_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
+In the lower part the penalty from Bai, (2004) is used and defined as $\widetilde{p}_{N,T}=\ln(T)p_{N,T}$.
 Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ uncorrelated over time, but weakly correlated across units, see (3.2).
-Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$, where we increase $\tau$ in $0.05$ steps from 00 to $0.95$.
-We also allow the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ to be uncorrelated over time ($\rho=\rho_{v}=0$, in blue) and weakly correlated ($\rho=\rho_{v}=0.5$, in red).
-The share of misselected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint.
-Neither of IC is strongly affected by weak dependence over time in the idiosyncratics.
-However, as shown in the Supplement, if the idiosynchratics in $\mathbf{X}_{i}$ are non-stationary as well, the IC equally fail to identify the correct number of CAs, with and without the adjustment of $T^{1+\tau}$.
-This again stresses the requirement for balanced stationarity in $\mathbf{F}$ and $\mathbf{V}_{i}$.
 (a) $IC^{DVS}_{1}$
@@
 Figure 1: Share of incorrect selected cross-sectional averages with increasing $\tau$.
-$N=T=100$.
 DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023).
-Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time ($\rho=\rho_{v}=0$) and weakly correlated if $\rho=\rho_{v}=0.5$.
-Both are weakly correlated across units.
-4 Illustration: a Discussion of Serlenga and Shin, (2007)
-MW apply $IC_{1}^{MW}$ and $IC_{2}^{MW}$ to revisit the study of Serlenga and Shin, (2007) (SS), who investigate the gravity equation of bilateral trade flows among 15 European countries ($N=91$ trading pairs) over 1960–2001 by the means of the CCE estimator.
-While SS include CAs arbitrarily, MW motivate their usage of IC as a theoretically backed and data-driven selection method.
-In neither of the studies are the time series properties of the variables examined.
-Variables of interest include $\mathbf{z}_{i,t}=(\textit{Trade}_{i,t},\textit{RER}_{i,t},\textit{TGDP}_{i,t},\textit{RLF}_{i,t},\textit{SIM}_{i,t},\textit{EEC}_{i,t},\textit{EMU}_{i,t})^{\prime}$, where $\textit{Trade}_{i,t}$ (the dependent variable) denotes a total bilateral trade flow for a trading pair $i$, $\textit{RER}_{i,t}$ is the log real exchange rate between the pair currencies, $\textit{TGDP}_{i,t}$ represents the economic mass of the partners, $\textit{RLF}_{i,t}$ measures differences in the relative factor endowments.
-Moreover, $\textit{SIM}_{i,t}$ is the logarithm of a relative economic size, while dummies $\textit{EEC}_{i,t}$ and $\textit{EMU}_{i,t}$ represent European Economic Community and the European Monetary Union, respectively.
-Finally, $\textit{RERT}_{t}$ (log real exchange rate between the EU currencies and the dollar), a constant ($1$) and a linear trend ($t$) are treated as observed factors.
-Therefore, $\overline{M}$ constitutes 10 CAs to select from, and MW select $\widehat{M}=\{1,t,\textit{RERT}_{t},\overline{Trade}_{t},\overline{TDGP}_{t},\overline{RER}_{t},\overline{\textit{EMU}}_{t}\}$ with $g=7$ to control for $\mathbf{F}$.
-$\widehat{M}$ minimizes the remaining CS correlation in the CCE residuals (see Table 1 in MW).
-We examine the series in SS further and detect strong evidence of unit roots in $\mathbf{F}$.
-Table 2 presents the results of the panel unit root test of Reese and Westerlund, (2016), which is the CAs equivalent of PANIC of Bai and Ng, (2004).
-Although we find strong evidence in favor of stationary idiosyncratics (at $1\%$ based on the PMSB statistic), we do not reject the null of unit root factors in virtually all series.
-[^5] The dummies $\textit{EEC}_{i,t}$ and $\textit{EMU}_{i,t}$ are omitted as they are unlikely to follow a typical factor structure.
-While MW acknowledge sensitivity of the results, our theory only strengthens the fact that the CCE-estimated coefficients in their Table 1 should be interpreted with some caution, and perhaps alternative combinations of CAs with $g\leq 10$ can provide a clearer picture.
-Series $Trade$ $TDGP$ $RER$ $RLF$ $SIM$ $RERT$
-Unit Roots in Factors
-Errors
-Table 2: PANICCA of Reese and Westerlund, (2016). - null of unit root not rejected; - the null rejected. For each series, the rest are used as regressors.
+Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time but weakly correlated across units.
+(a) $IC^{DVS}_{1}$
+(b) $IC^{MW}_{1}$
+Figure 2: Share of incorrect selected cross-sectional averages with increasing $T$, $\tau=0.9$.
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023).
+Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time but weakly correlated across units.
 4 Conclusion
-Our theory and simulation results show the inability of IC to accommodate non-stationary factors.
-The number of CAs is severely underselected, independent of the number of time periods and units.
-Underselection has severe consequences in terms of our examples.
-For CCE, the number of factors may be larger than the selected number of CAs, which makes the estimator inconsistent.
-In DCE case, the bootstrap procedure cannot help accommodate different factors in $\mathbf{X}$ and $\mathbf{y}$ anymore.
-In forecasting, incorrectly estimated factors may harm the forecast comparison procedures.
-6 Appendix
+This study is the first to examine selection of an optimal set of CAs in CCE and related settings by ICs inspired by Bai and Ng, (2002) when latent factors are non-stationary.
+In particular, we use mild integration in order to explore varying degrees of factor persistence, and demonstrate that ICs remain consistent without any modifications.
+However, the more persistent common factors are, the worse small sample performance becomes, and ICs regain selection consistency only in very large samples (i.e.
+$(N,T)>300$, according to our experiments).
+Importantly, a diverging penalty suggested by Bai, (2004) under non-stationary factors for PC renders ICs inconsistent in the case of CAs.
+Therefore, our recommendation for CCE practitioners is not to take prescriptions from the PC literature automatically, and interpret IC results with caution in presence of highly persistent data, unless $N,T$ are substantial.
+5 Supplement
+5.1 Assumptions
+Throughout our analysis, we employ the following set of assumptions.
+Assumption 1.
+  • (a)
+    $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009). In particular,
+    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\frac{\mathbf{G}}{T^{\tau}},\quad\mathbf{G}>0\hskip 5.69054pt(diagonal),$
+    where $\mathbf{u}_{f,t}$ is a zero-mean linear process.
+  • (b)
+    If $\tau=0$, then $\{\mathbf{f}_{t}\}$ is a stationary process with summable autocovariances.
+Assumption 2. Let $\mathbf{e}_{i,t}=(\varepsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}\in\mathbb{R}^{k+1}$. Then
+  • (a)
+    • (i)
+      If $\tau\in(0,1)$, then $\{\mathbf{e}_{i,t}\}$ is a martingale difference sequence with $\mathbb{E}(\mathbf{e}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\mathbf{e}_{i,t}\mathbf{e}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{ee,i,t}$ with $\lim_{N,T\to\infty}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\boldsymbol{\Sigma}_{ee,i,t}=\boldsymbol{\Sigma}_{ee}=\mathrm{diag}(\sigma^{2},\boldsymbol{\Sigma}_{\mathbf{v}})$ positive definite and $\mathbb{E}(\|\mathbf{e}_{i,t}\|^{4})<\infty$.
+    • (ii)
+      If $\tau=0$, then we let $\mathbf{e}_{i,t}=\mathbf{K}_{i}(L)\boldsymbol{\epsilon}_{i,t}=\sum_{j=0}^{\infty}\mathbf{K}_{i,j}\boldsymbol{\epsilon}_{i,t-j}$, where $\boldsymbol{\epsilon}_{i,t}$ is independent across $t$ with $\mathbb{E}(\boldsymbol{\epsilon}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\boldsymbol{\epsilon}_{i,t}\boldsymbol{\epsilon}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{\epsilon\epsilon,i,t}$ positive definite, $\mathbb{E}(\|\boldsymbol{\epsilon}_{i,t}\|^{4})<\infty$, and $\sum_{j=0}^{\infty}j^{1/2}\|\mathbf{K}_{i,j}\|<\infty$.
+Also, $\boldsymbol{\xi}_{t}=\mathrm{vec}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}-\mathbb{E}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right)\right)$ is strong mixing with coefficients of size bd(bd)-bd(b-d) with $b>4$ and $b>d>2$, $\mathbb{E}(\|\boldsymbol{\xi}_{t}\|^{b})<\infty$, and $\lim_{T\to\infty}T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{s}^{\prime})$ is positive definite.
+Here and throughout, $\left\|\mathbf{A}\right\|=\sqrt{\mathrm{trace}(\mathbf{A}^{\prime}\mathbf{A})}$ is Frobenius norm.
+    • (iii)
+      In both cases, $\varepsilon_{i,t}$ and $\mathbf{v}_{j,s}$ are independent for all $i,j,t,s$.
+  • (b)
+    We have $\left\|\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{e}_{i,t}\mathbf{e}_{j,t}^{\prime}\right)-\boldsymbol{\Sigma}_{ee}\right\|=o(1)$ as $(N,T)\to\infty$ for a positive definite matrix $\boldsymbol{\Sigma}_{ee}$.
+Assumption 3. $\mathbf{f}_{t}$ and $\mathbf{e}_{i,s}$ are independent for all $t$, $s$ and $i$.
+Assumption 4.
+$\mathbf{C}_{i}$ is a deterministic matrix, such that $\|\mathbf{C}_{i}\|<\infty$ and $\frac{1}{N}\sum_{i=1}^{N}\mathbf{C}_{i}\mathbf{C}_{i}^{\prime}\to\boldsymbol{\Sigma}_{\mathbf{C}}$ positive definite.
+Also, $\overline{\mathbf{C}}=[\overline{\mathbf{C}}_{m},\overline{\mathbf{C}}_{-m}]$, where $\overline{\mathbf{C}}_{-m}\in\mathbb{R}^{m\times(K-m)}$ and $\overline{\mathbf{C}}_{m}=\overline{\mathbf{C}}\mathbf{q}_{M_{0}}\in\mathbb{R}^{m\times m}$ for a unique $M_{0}$ is full rank for all $N$, including $N\to\infty$.
+If $m=K$, then $\overline{\mathbf{C}}=\overline{\mathbf{C}}_{m}$.
 5.2 Auxiliary Results
@@
 $=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\right)\right]\right)$
-$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\right)\right]\right)$
+$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\right)\right]\right)$
+$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}|\mathcal{F}_{(t-1)\lor(s-1)}\right)\right)\right]\right)$
 $=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{t}\right)\right]\right)$
@@
 Under Assumptions 1 - 4, but $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ are uncorrelated over time, we have that
-$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{U}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$
+$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$
 as $T\to\infty$
@@
 $=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\right)\right]\right)$
-$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\right)\right]\right)$
+$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\right)\right]\right)$
+$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}|\mathcal{F}_{(t-1)\lor(s-1)}\right)\right)\right]\right)$
 $=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,t}\right)\right]\right)$
@@
 $=\ln\mathrm{det}\left(\mathbf{I}_{k}+\overline{\mathbf{Q}}_{M}\overline{\mathbf{Q}}_{M_{0}}^{-1}-\overline{\mathbf{Q}}_{M_{0}}\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
-$=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+k(g-m)p_{N,T},$ (A.8)
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\left[T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T},$ (A.8)
 where we used $\ln(a)-\ln(b)=\ln\left(\frac{a}{b}\right)$ and $\mathrm{det}(\mathbf{A}\mathbf{B})=\mathrm{det}(\mathbf{A})\mathrm{det}(\mathbf{B})$.
-This implies that we will examine the asymptotic behavior of $\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$.
+This implies that we will examine the asymptotic behavior of $T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)$.
 Its behavior is determined by whether $M_{0}\subset M$ (over-specification) or $M\subset M_{0}$ (under-specification).
@@
 We only analyze the case of $M\subset M_{0}$ as in Margaritella and Westerlund, (2023) or De Vos and Stauskas, (2024), because other cases will lead to the same asymptotic conclusions.
-6.3 Part $M\subset M_{0}$
+5.3.1 Case of $M\subset M_{0}$
 By following De Vos and Stauskas, (2024), we have
-$\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$ $=\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]+\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
+$T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)$ $=T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-T^{-\tau}\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
+$+T^{-\tau}\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
 $=\mathbf{I}-\mathbf{II}+\mathbf{III}.$ (A.9)
@@
 We begin with
-$\mathbf{I}$ $=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
-$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
-$+\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}$
+$\mathbf{I}$ $=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
+$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}$
 $=\mathbf{I}_{a}+\mathbf{I}_{b}+\mathbf{I}_{c}+\mathbf{I}_{d}.$ (A.10)
@@
 $=O_{p}(T^{-\tau})+O_{p}(\sqrt{N}T^{-1/2}T^{-\tau/2})=O_{p}(1)$ (A.13)
-without further restrictions (but negligible under $TN^{-1}=O(1)$. Also,
+without further restrictions (but negligible under $TN^{-1}=O(1)$. Also, let
 $T^{\tau}\times\xi_{N,T,\tau}^{-1}$ $=T^{\tau}\times(O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(1+\tau)/2}))$
-$=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})=o_{p}(1)$ (A.14)
-without further restrictions as $\tau\in(0,1)$. Both of these results will be important when determining the orders of many terms.
+$=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$
+$=o_{p}(1))=R_{N,T}$ (A.14)
+without further restrictions as $\tau\in(0,1)$.
+Note that the latter will be the dominant component in the further analysis.
+Both of these results will be important when determining the orders of many terms.
 Next, in case of $M\subset M_{0}$, we have that $\mathrm{rank}\left(\frac{1}{T^{1+\tau}}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)=m$ even asymptotically, and hence
-$\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}=\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}+O_{p}(\xi_{N,T,\tau}^{-1}).$ (A.15)
+$\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}=\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}+O_{p}(\xi_{N,T,\tau}^{-1}),$ (A.15)
+because
+$\left\|\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}-\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}\right\|$
+$=\left\|\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}-T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}\right\|$
+$\leq\left\|\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}\right\|\left\|T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}-\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|\left\|\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}\right\|$
+$=O_{p}(\xi_{N,T,\tau}^{-1}).$ (A.16)
 This leads to the expansion
@@
 which now takes into account the mildly integrated factors. Then, we can continue with
-$\left\|\mathbf{I}_{a}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
-$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
-$\leq\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}}_{O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(2\tau+1)/2})}$
-$+2\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(N^{-1}T^{-(1+\tau)/2})+O_{p}(N^{-1/2}T^{-(2+\tau)/2})}$
-$+\underbrace{T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|^{2}}_{O_{p}(T^{-1}\xi_{N,T,\tau}^{-1})}$
-$=O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(2\tau+1)/2}),$ (A.17)
-which is the dominating order. Next, we have
-$\left\|\mathbf{I}_{b}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$\leq\underbrace{N^{-1}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}((NT)^{-1})}$
-$+\underbrace{2T^{(\tau-1)/2}N^{-1/2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{(\tau-1)/2}N^{-1/2})}$
-$+\underbrace{T^{\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{x}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})}$
-$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.18)
-which is a slowly decaying term dominated by $T^{(\tau-1)/2}=o(1)$, since $\tau\in(0,1)$. We let $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$, and we move on to
-$\left\|\mathbf{I}_{c}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$\leq\underbrace{T^{-(1+\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{o_{p}(C_{N,T}^{-2})}$
-$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(C_{N,T}^{-2})}$
-$+\underbrace{N^{-1/2}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1})}$
-$+\underbrace{T^{(\tau-1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{T^{(\tau-1)/2}\xi_{N,T,\tau}^{-1})}$
-×$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
-$=O_{p}(C_{N,T}^{-2}),$ (A.19)
+$\left\|\mathbf{I}_{a}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
+$\leq\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
+$\leq\underbrace{T^{-2\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}}_{O_{p}(N^{-1}T^{-2\tau})+O_{p}(N^{-1/2}T^{-(4\tau+1)/2})}$
+$+2\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(N^{-1}T^{-(1+2\tau)/2})+O_{p}(N^{-1/2}T^{-(2+3\tau)/2})}$
+$+\underbrace{T^{-1-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|^{2}}_{O_{p}(T^{-(1+\tau)}\xi_{N,T,\tau}^{-1})}$
+$=o_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.18)
+where each of the term is of the lower order than $O_{p}(\xi_{N,T,\tau}^{-1})$. Next, we have
+$\left\|\mathbf{I}_{b}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$\leq\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$\leq\underbrace{N^{-1}T^{-1-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}(N^{-1}T^{-1-\tau})}$
+$+\underbrace{2T^{-(\tau+1)/2}N^{-1/2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{-(\tau+1)/2}N^{-1/2})}$
+$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}(\xi_{N,T,\tau}^{-1})}$
+$=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.19)
+which remains a slowly decaying term even if scaled by $T^{\tau}$, because it will be dominated by $T^{(\tau-1)/2}=o(1)$ stemming from the second term, since $\tau\in(0,1)$.
+We let $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$, and we move on to
+$\left\|\mathbf{I}_{c}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$\leq\underbrace{T^{-(1+3\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{O_{p}(T^{-(1+3\tau)/2}N^{-1/2}C_{N,T}^{-2})}$
+$+\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{-\tau}C_{N,T}^{-2})}$
+$+\underbrace{N^{-1/2}T^{-(1+\tau)}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1-\tau})}$
+$+\underbrace{T^{-(\tau+1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(T^{-(\tau+1)/2}\xi_{N,T,\tau}^{-1})}$
+$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
+$=O_{p}(T^{-\tau}C_{N,T}^{-2})+O_{p}(N^{-1/2}T^{-1-\tau})=o_{p}(T^{-\tau}),$ (A.20)
 which is negligible. Ultimately,
-$\left\|\mathbf{I}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
-$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
-$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
-$\leq\underbrace{T^{-(1+\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{o_{p}(C_{N,T}^{-2})}$
-$+\underbrace{N^{-1/2}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1})}$
-$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(C_{N,T}^{-2})}$
-$+\underbrace{T^{(\tau-1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(T^{\tau/2}C_{N,T^{\tau}}^{-2})}$
-×$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
-$=O_{p}(C_{N,T}^{-2}),$ (A.20)
+$\left\|\mathbf{I}_{d}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
+$\leq\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
+$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
+$\leq\underbrace{T^{-(1+3\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{O_{p}(T^{-(1+3\tau)/2}N^{-1/2}C_{N,T}^{-2})}$
+$+\underbrace{N^{-1/2}T^{-1-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1-\tau})}$
+$+\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{-\tau}C_{N,T}^{-2})}$
+$+\underbrace{T^{-(\tau+1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(T^{-(\tau+1)/2}\xi_{N,T,\tau}^{-1})}$
+$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
+$=O_{p}(T^{-\tau}C_{N,T}^{-2})+O_{p}(N^{-1/2}T^{-1-\tau})=o_{p}(T^{-\tau}),$ (A.21)
 which has an identical asymptotic behavior to the one of $\mathbf{I}_{c}$. Hence, overall,
-$\left\|\mathbf{I}\right\|=\left\|\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right\|=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.21)
+$\left\|\mathbf{I}\right\|=\left\|T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right)\right\|=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.22)
 which is the slowest decaying term. Importantly, we also have that
-$\left\|\mathbf{I}\mathbf{I}\right\|=\left\|\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right\|=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.22)
+$\left\|\mathbf{I}\mathbf{I}\right\|=\left\|T^{-\tau}\left(\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right)\right\|=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.23)
 as in this case $m=g$, and therefore the asymptotic behavior of this term will be identical.
@@
 By following this, we obtain
-$\mathbf{III}$ $=\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}$
-$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})\mathbf{X}_{i}$
-$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
-$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}+\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
-$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{X}_{i}$
-$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
-$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
+$\mathbf{III}$ $=T^{-\tau}\left(\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right)$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})\mathbf{X}_{i}$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}+\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}-\mathbf{P}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{X}_{i}$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
+$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
 $=\mathbf{III}_{a}+\mathbf{III}_{b}+\mathbf{III}_{c}+\mathbf{III}_{d},$ (A.24)
-where the term $\mathbf{III}_{d}$ is particularly problematic. By using random loadings assumption, it explicitly can be written as
-$\mathbf{III}_{d}$ $=\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
-$=T^{-1}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}$
-$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}$
-$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}$
-$+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}$
-$=\mathbf{III}^{1}_{d}+\mathbf{III}^{2}_{d}+\mathbf{III}_{d}^{3}+\mathbf{III}^{4}_{d}.$ (A.24)
-Note that the key component in all $\mathbf{III}^{1}_{d}-\mathbf{III}^{4}_{d}$ is $\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$, where
+where we will start the term $\mathbf{III}_{d}$. Note that it explicitly looks like
+$\mathbf{III}_{d}$ $=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$ (A.25)
+with the key component $\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$, where
 $T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$ $=T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}-T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}$
@@
 $\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}}-\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\boldsymbol{\Sigma}_{\mathbf{F}_{\mathbf{x}}},$ (A.26)
-as $T\to\infty$, which is positive definite. Now, moving on to $\mathbf{III}^{1}_{d}$, we obtain
-$\mathbf{III}^{1}_{d}$ $=T^{-1}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$
-$=T^{\tau}\boldsymbol{\Gamma}^{\prime}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}\underbrace{(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}})}_{O_{p}(1)}$
-$=O_{p}(T^{\tau}),$ (A.26)
-which means that the whole term is divergent, and the same applies to $\mathbf{III}^{1}_{d}$. In particular, $\mathbf{III}^{4}_{d}$ has the same order, since
-$\left\|\mathbf{III}^{4}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|$
-$\leq T^{\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|^{2}$
-×$\times\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}T^{-(1+\tau)}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
-$=O_{p}(T^{\tau}),$ (A.27)
-while
-$\left\|\mathbf{III}^{2}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|$
-$\leq T^{\tau}\left\|\boldsymbol{\Gamma}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
-×$\times\underbrace{\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|}_{O_{p}(N^{-1/2})}$
-$=O_{p}(T^{\tau}N^{-1/2}),$ (A.28)
-which is negligible only if $\tau\in(0,0.5)$ under $TN^{-1}=O(1)$ (to relate this term to (6.3) and the subsequent similar terms), and the same condition applies to $\mathbf{III}^{3}_{d}$:
-$\left\|\mathbf{III}^{3}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}\right\|$
-$\leq T^{\tau}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}\right\|$
-×$\times\underbrace{\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|}_{O_{p}(N^{-1/2})}$
-$=O_{p}(T^{\tau}N^{-1/2}),$ (A.29)
-as welll. Next, we move on to $\mathbf{III}_{a}$, and for this we firstly analyze
+as $T\to\infty$, which is positive definite constant matrix (see a similar result in A.
+51 in the Supplement of Margaritella and Westerlund,, 2023).
+Therefore, clearly $\mathbf{III}_{d}$ converges to a positive semi-definite matrix, because it is a matrix-valued form.
+Importantly, it is not a zero matrix, because, according to Exercise 8.13 (b) in Abadir and Magnus, (2005), $\mathbf{III}_{d}=\mathbf{0}_{k\times k}$ is equivalent to $\mathrm{tr}(\mathbf{III}_{d})=0$ for a positive semi-definite matrix.
+Note that
+$\mathrm{tr}\left(\mathbf{III}_{d}\right)=\mathrm{tr}\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right)$
+$=\mathrm{tr}\left(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\left[\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}\right]\right)$
+$=\mathrm{tr}\left(\left[\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}\right]T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right)$
+$\geq\lambda_{\min}\left(\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}\right)$
+$\times\mathrm{tr}\left(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right)$
+$\to_{p}c>0$ (A.27)
+as $(N,T)\to\infty$, for a positive constant $c$ by the results in Fang et al., (1994).
+This follows from the fact that second positive semi-definite matrix in the product is non-zero by Exercise 8.26 (c) in Abadir and Magnus, (2005).
+Particularly, for a positive definite $\mathbf{A}\in\mathbb{R}^{p\times p}$ and $\mathbf{B}\in\mathbb{R}^{p\times q}$, $\mathbf{B}^{\prime}\mathbf{A}\mathbf{B}=\mathbf{0}_{q\times q}$ if and only if $\mathbf{B}=\mathbf{0}_{p\times q}$, which means that $\mathrm{rk}(\mathbf{B})=0$.
+In our case, $\mathbf{B}=T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$ and we know that $\mathrm{rk}\left(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right)=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}{\mathbf{q}}_{M^{C}})=m_{x}-g>0$ even in the limit.
+Also, $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}$ is an $m_{x}\times m_{x}$ positive definite matrix by our assumptions.
+Next in our analysis, we move on to $\mathbf{III}_{a}$, and for this we firstly analyze
 $\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$ $\leq\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|$
@@
 $=O_{p}(1).$ (A.28)
-Next,
-$\left\|\mathbf{III}_{a}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
-$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
-$\leq T^{-1}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|$
-$=O_{p}(T^{-1}),$ (A.31)
-which would not be negligible under correlated idiosyncratics (see Lemma 2). The remaining two terms, are slowly vanishing. In particular,
-$\left\|\mathbf{III}_{b}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
-$\leq T^{(\tau-1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
-×$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
-$=O_{p}(T^{(\tau-1)/2}),$ (A.32)
-and
-$\left\|\mathbf{III}_{c}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
-$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
-$\leq T^{(\tau-1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
-×$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
-$=O_{p}(T^{(\tau-1)/2}),$ (A.33)
+This gives us
+$\left\|\mathbf{III}_{a}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
+$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
+$\leq T^{-1-\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|$
+$=O_{p}(T^{-1-\tau})=o_{p}(T^{-\tau}),$ (A.29)
+which would not be negligible under correlated idiosyncratics (see Lemma 2). The remaining two terms are vanishing, as well. In particular,
+$\left\|\mathbf{III}_{b}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
+$\leq T^{-(\tau+1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
+$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
+$=O_{p}(T^{-(\tau+1)/2})=o_{p}(T^{-\tau}),$ (A.30)
+for $\tau\in(0,1)$, and
+$\left\|\mathbf{III}_{c}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
+$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
+$\leq T^{-(\tau+1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
+$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
+$=O_{p}(T^{-(\tau+1)/2})=o_{p}(T^{-\tau}),$ (A.31)
 which is just a transpose of $\mathbf{III}_{b}$. Note again, that both of the terms would slowly diverge under the conditions of Lemma 2. Therefore, by combining our interim results, we obtain
-$\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$ $=\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]+\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
-$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})+O_{p}(T^{\tau})+O_{p}(T^{(\tau-1)/2})$
-$=O_{p}(T^{\tau}),$ (A.34)
-which is the dominant order. This implies that
-$T^{-\tau}(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}})$ $=T^{-\tau}\mathbf{III}_{d}+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-(\tau+1)/2})$
-$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
-$+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-(\tau+1)/2}),$ (A.35)
-which converges to a positive semi-definite matrix.
-However, note how by using the fact that $\overline{\mathbf{Q}}_{M_{0}}=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})$, we have that
-$T^{-\tau}\overline{\mathbf{Q}}_{M_{0}}$ $=T^{-\tau}\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+O_{p}(\xi_{N,T,\tau}^{-1})$
-$=T^{-\tau}\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}}+O_{p}(\xi_{N,T,\tau}^{-1})$
-$=\frac{1}{T^{1+\tau}}\frac{1}{N}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{X}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
-$=\frac{1}{T^{1+\tau}}\frac{1}{N}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{V}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
-$=\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}-T^{-1}\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
-$=O_{p}(T^{-\tau})+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-1})$
-$=o_{p}(1),$ (A.36)
-because the idiosyncratics are stationary. Combining (6.3) and (6.3), we get
-$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
-$=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right](T^{-\tau}\overline{\mathbf{Q}}_{M_{0}})^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+o(1)$
-↛$\not\to_{p}c>0,$ (A.37)
-as in total $T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right](T^{-\tau}\overline{\mathbf{Q}}_{M_{0}})^{-1}$ does not converge to a positive semi-definite matrix. Therefore
-$\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0\right)\not\to 0,$ (A.38)
-and the under-specification risk is not eliminated.
-6.4 Part $M_{0}\subset M$
+$T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)$ $=T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-T^{-\tau}\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
+$+T^{-\tau}\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
+$=\mathbf{III}_{d}+o_{p}(T^{-\tau}),$ (A.32)
+which is a positive semi-definite matrix and $o_{p}(T^{-\tau})=O_{p}(\xi_{N,T,\tau}^{-1})$.
+Further, note how by using the fact that $\overline{\mathbf{Q}}_{M_{0}}=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+o_{p}(1)$ from (A.23), we have that
+$\overline{\mathbf{Q}}_{M_{0}}$ $=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+o_{p}(1)$
+$=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}}+o_{p}(1)$
+$=\frac{1}{T}\frac{1}{N}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{X}_{i}+o_{p}(1)$
+$=\frac{1}{T}\frac{1}{N}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{V}_{i}+o_{p}(1)$
+$=\frac{1}{N}\sum_{i=1}^{N}T^{-1}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}-T^{-1}\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}+o_{p}(1)$
+$=\frac{1}{N}\sum_{i=1}^{N}T^{-1}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}+o_{p}(1)$
+$\to_{p}\boldsymbol{\Sigma}_{\mathbf{v}},$ (A.33)
+because the idiosyncratics are stationary. This is a positive definite matrix. In what follows, we use Theorem 2 and Exercise 1 in Section 4 of Magnus and Neudecker, (2019). In particular,
+$\ln\mathrm{det}(\mathbf{A}+\mathbf{H})=\ln\mathrm{det}(\mathbf{A})+\mathrm{tr}(\mathbf{A}^{-1}\mathbf{H})+O\left(\left\|\mathbf{A}^{-1}\mathbf{H}\right\|^{2}\right),$ (A.34)
+for a positive definite matrix $\mathbf{A}$ and a symmetric perturbation $\mathbf{H}$.
+Further, to simplify notation, let $\boldsymbol{\Omega}_{M,M_{0}}:=T^{-\tau}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}$ and $\boldsymbol{\Omega}_{M,M_{0}}^{0}:=\mathbf{III}_{d}\times\overline{\mathbf{Q}}_{M_{0}}^{-1}$, which we know is a positive semi-definite matrix.
+Next, let $\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}=\boldsymbol{\Omega}_{M,M_{0}}-\boldsymbol{\Omega}_{M,M_{0}}^{0}$, where we know that $\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau})$.
+In our setting, $\mathbf{A}=\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}$ and $\mathbf{H}=\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}$.
+Therefore,
+$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)$ $=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}+\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)$
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)$
+$+\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}+T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)$
+$-\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)$
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)+\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]$
+$+O_{p}\left(\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|^{2}\right)$
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}).$ (A.35)
+Here, we used the following important facts.
+Let $\lambda_{NT,j}\geq 0$ be an eigenvalue of $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ for $j=1,\ldots,k$ with the probability limit $\lambda>0$ or 0 immediately.
+As this positive semi-definite matrix is nonzero even in the limit, we know that not all eigenvalues are zero.
+Then we know that the eigenvalues of $\mathbf{B}=\mathbf{I}_{k}+c\mathbf{A}$ are given by $1+c\lambda_{j}$ for a constant $c$, where $\lambda_{j}$ is an eigenvalue of generic square matrix $\mathbf{A}$.
+In particular, for the $j$-th eigenvector of $\mathbf{A}$ ($\boldsymbol{\nu}_{j}$), we have
+$\mathbf{B}\boldsymbol{\nu}_{j}=(\mathbf{I}_{k}+c\mathbf{A})\boldsymbol{\nu}_{j}=\boldsymbol{\nu}_{j}+c\mathbf{A}\boldsymbol{\nu}_{j}=1\times\boldsymbol{\nu}+c\lambda_{j}\boldsymbol{\nu}_{j}=(1+c\boldsymbol{\lambda}_{j})\boldsymbol{\nu}_{j}.$ (A.36)
+Next, we know that eigenvalues of $\mathbf{B}^{-1}=(\mathbf{I}_{k}+c\mathbf{A})^{-1}$ are given by $\frac{1}{1+c\lambda_{j}}$.
+Which means that eigenvalues of $\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}$ are given by $\frac{1}{1+T^{\tau}\lambda_{NT,j}}$.
+Note that we can diagonalize this matrix and obtain
+$\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}=\boldsymbol{\Upsilon}\left(\underbrace{\mathrm{diag}\left[\frac{1}{1+T^{\tau}\lambda_{NT,1}},\ldots,\frac{1}{1+T^{\tau}\lambda_{NT,k}}\right]}_{=\boldsymbol{\Lambda}_{N,T}}\right)\boldsymbol{\Upsilon}^{\prime},$ (A.37)
+where $\boldsymbol{\Upsilon}$ is an orthonormal vector.
+Let $b$ denote the number of eigenvalues of $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ that is strictly positive.
+Then kbk-b is zero eigenvalues, which means that we will have kbk-b eigenvalues of $\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}$ that are equal to 1.
+Then by using the trace properties, we obtain
+$\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|=\sqrt{\mathrm{tr}\left(\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}\boldsymbol{\Upsilon}^{\prime}\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}\boldsymbol{\Upsilon}^{\prime}\right)}$ $=\sqrt{\mathrm{tr}\left(\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}^{2}\boldsymbol{\Upsilon}^{\prime}\right)}$
+$=\sqrt{\mathrm{tr}\left(\boldsymbol{\Upsilon}^{\prime}\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}^{2}\right)}$
+$=\sqrt{\sum_{j=1}^{k}\frac{1}{(1+T^{\tau}\lambda_{NT,j})^{2}}}$
+$\to_{p}\sqrt{k-b},$ (A.38)
+because $\lambda_{NT,j}=0$ or $\lambda_{NT,j}\to_{p}\lambda_{j}>0$, where the latter case produces $b$ zero entries into the total sum, and the former case gives kbk-b units.
+Therefore, $\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|=O_{p}(1)$.
+Further,
+$\left\|\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]\right\|$ $\leq\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|T^{\tau}\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|$
+$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$
+$=R_{N,T}.$ (A.39)
+Therefore, combining (5.3.1) and (5.3.1), we get
+$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)+k(g-m)p_{N,T}$
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\left(\boldsymbol{\Omega}_{M,M_{0}}^{0}+\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)\right)+k(g-m)p_{N,T}$
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}+T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)+k(g-m)p_{N,T}$
+$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)+k(g-m)p_{N,T}+R_{N,T}.$ (A.40)
+Clearly, the difference diverges to $\infty$, because $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ converges to a nonzero positive semi-definite matrix.
+To complete the proof, it is useful to derive the rate at which the divergence occurs.
+Again, recall that $\lambda_{NT,j}\geq 0$ is an eigenvalue of $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ for $j=1,\ldots,k$ with the probability limit $\lambda>0$ or zero immediately.
+As this positive semi-definite matrix is nonzero even in the limit, we know that not all eigenvalues are zero.
+Recall that $b$ is the number of positive eigenvalues.
+Therefore, using $\ln(1+x)=x+O(x^{2})$ for a small $x$ and $p_{N,T}/\ln(T)=o(1)$, we obtain the following:
+$(\mathrm{IC}^{DVS}(M)$ $-\mathrm{IC}^{DVS}(M_{0}))/\ln(T)=\left[\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)\right]/\ln(T)+o_{p}(1)$
+$=\left[\ln\prod_{j:\lambda>0}(1+T^{\tau}\lambda_{NT,j})\right]/\ln(T)+o_{p}(1)$
+$=\left[\sum_{j:\lambda>0}\ln(1+T^{\tau}\lambda_{NT,j})\right]/\ln(T)+o_{p}(1)$
+$=\left[\sum_{j:\lambda>0}\ln(T^{\tau}\lambda_{NT,j})\right]/\ln(T)+\left[\sum_{j:\lambda>0}[\ln(1+T^{\tau}\lambda_{NT,j})-\ln(T^{\tau}\lambda_{NT,j})]\right]/\ln(T)+o_{p}(1)$
+$=\frac{\tau b\ln(T)}{\ln(T)}+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\ln(\lambda_{NT,j})+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\ln\left(1+\frac{1}{T^{\tau}\lambda_{NT,j}}\right)+o_{p}(1)$
+$=\tau b+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\ln(\lambda_{NT,j})+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\left(\frac{1}{T^{\tau}\lambda_{NT,j}}+O(T^{-2\tau})\right)+o_{p}(1)$
+$\to_{p}\tau b>0,$ (A.41)
+because $\lambda_{NT,j}\to_{p}\lambda_{j}>0$ as $(N,T)\to\infty$. This overall implies that
+$\mathbb{P}\left(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})<0\right)=\mathbb{P}\left(\frac{\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})}{\ln(T)}<0\right)\to 0$ (A.42)
+as $(N,T)\to\infty$.
+5.3.2 Case of $M_{0}\subset M$
 We now move to the case when $M_{0}\subset M$ (over-specification).
 To analyze this case we will introduce additional notation.
-We decompose the selector matrix $\mathbf{q}_{M}=[\mathbf{q}_{M_{0}},\mathbf{q}_{M_{0}^{C}}]$, where $\mathbf{q}_{M_{0}}\in\mathbb{R}^{k\times m}$ and $\mathbf{q}_{M_{0}^{C}}\in\mathbb{R}^{k\times(g-m)}$, where the latter selector corresponds to $g-m$ excess (over-selected) averages.
+We decompose the selector matrix $\mathbf{q}_{M}=[\mathbf{q}_{M_{0}},\mathbf{q}_{M_{0}^{C}}]$, where $\mathbf{q}_{M_{0}}\in\mathbb{R}^{k\times m}$ and $\mathbf{q}_{M_{0}^{C}}\in\mathbb{R}^{k\times(g-m)}$, where the latter selector corresponds to gmg-m excess (over-selected) averages.
 By using this, we can further decompose
@@
 In what follows, the post-multiplication by this matrix leads to
-$\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\overline{\mathbf{H}}_{M}=\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M}+\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M}=\begin{bmatrix}\mathbf{F}_{\mathbf{x}},\mathbf{0}_{T\times(g-m)}\end{bmatrix}+\begin{bmatrix}\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}},\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}^{C}}\end{bmatrix}.$ (A.42)
+$\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\overline{\mathbf{H}}_{M}=\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M}+\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M}=\begin{bmatrix}\mathbf{F}_{\mathbf{x}},\mathbf{0}_{T\times(g-m)}\end{bmatrix}+\begin{bmatrix}\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}},\overline{\mathbf{V}}\mathbf{q}_{M}\overline{\mathbf{H}}_{M_{0}^{C}}\end{bmatrix}$.
+(A.46)
 Finally, we introduce $\mathbf{D}_{M}=\mathrm{diag}(\mathbf{I}_{m},\sqrt{N}\mathbf{I}_{g-m})$, such that
@@
 $=\mathbf{F}_{\mathbf{x}}^{0}+\overline{\mathbf{V}}_{M}^{0},$ (A.47)
-where $\left\|\overline{\mathbf{V}}_{M_{1}}^{0}\right\|=O_{p}(N^{-1/2})$, but $\left\|\overline{\mathbf{V}}_{M_{2}}^{0}\right\|=O_{p}(1)$, such that the last $g-m$ columns are non-degenerate.
+where $\left\|\overline{\mathbf{V}}_{M_{1}}^{0}\right\|=O_{p}(N^{-1/2})$, but $\left\|\overline{\mathbf{V}}_{M_{2}}^{0}\right\|=O_{p}(1)$, such that the last gmg-m columns are non-degenerate.
 Further, let us introduce $\mathbf{G}_{T}=\mathrm{diag}(T^{(1+\tau)/2}\mathbf{I}_{m},T^{1/2}\mathbf{I}_{g-m})$.
@@
 $=\mathbf{S}+O_{p}(N^{-1/2}T^{-\tau/2})+O_{p}(T^{-1/2}),$ (A.48)
-where $\mathbf{S}=\mathrm{diag}(T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}},T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})\in\mathbb{R}^{g\times g}$.
-The rate comes from the leading, but vanishing components of the last three terms on the right-hand side of (6.4).
+where $\mathbf{S}=\mathrm{diag}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}},T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})\in\mathbb{R}^{g\times g}$.
+The rate comes from the leading, but vanishing components of the last three terms on the right-hand side of (5.3.2).
 Specifically,
@@
 $\left|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\leq N^{-1/2}T^{-(1+\tau)/2}\underbrace{\left\|\sqrt{N}T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\right\|}_{O_{p}(1)}\left\|\mathbf{q}_{M}\right\|\left\|\overline{\mathbf{H}}_{M_{0}^{C}}\right\|=O_{p}(N^{-1/2}T^{-(1+\tau)/2})$ (A.52)
-by Corollary (1). This implies that
-$\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|=O_{p}(N^{-1/2}T^{-\tau/2})+O_{p}(T^{-1/2}),$ (A.49)
+by Corollary (1). By following the approach in Bai, (2004), we see that this implies that
+$\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|$ $=\left\|\mathbf{S}^{+}(\mathbf{S}-\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}\right\|$
+$\leq\left\|\mathbf{S}^{+}\right\|\left\|\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}-\mathbf{S}\right\|\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}\right\|$
+$=O_{p}(N^{-1/2}T^{-\tau/2})+O_{p}(T^{-1/2}),$ (A.53)
 so the rate result applies to the MP inverses, as well.
@@
 $+\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}((\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+})\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}.$ (A.71)
-Hence, if we insert (6.4) into $\mathbf{II}$, we obtain $\mathbf{II}=\mathbf{II}^{1}-\mathbf{II}^{2}-\mathbf{II}^{3}+\mathbf{II}^{4}$, and so
+Hence, if we insert (5.3.2) into $\mathbf{II}$, we obtain $\mathbf{II}=\mathbf{II}^{1}-\mathbf{II}^{2}-\mathbf{II}^{3}+\mathbf{II}^{4}$, and so
 $\left\|\mathbf{II}^{1}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})\mathbf{V}_{i}\right\|$
@@
 $+2\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}_{M_{1}}^{0}\right\|\left\|(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
-$+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1/2}\mathbf{V}_{i}^{\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|^{2}\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|}_{=o_{p}(C_{N,T})^{-2}}$
+$+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1/2}\mathbf{V}_{i}^{\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|^{2}\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|}_{=o_{p}(C_{N,T}^{-2})}$
 $=O_{p}(N^{-1})+O_{p}(T^{-1})$
@@
 $=O_{p}(N^{-1/2})+O_{p}(T^{-1/2})=O_{p}(C_{N,T}^{-1}),$ (A.73)
-and so the last term in (6.4) is $o_{p}(C_{N,T}^{-2})$. The other terms are clearly dominated. Next, we go to
+and so the last term in (5.3.2) is $o_{p}(C_{N,T}^{-2})$. The other terms are clearly dominated. Next, we go to
 $\left\|\mathbf{II}^{2}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1\prime}\overline{\mathbf{V}}_{M_{0}}^{0\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})\mathbf{V}_{i}\right\|$
@@
 $=O_{p}(N^{-1/2}),$ (A.75)
-and so the overall order is dominated by the first and third terms in (6.4). Further, note that $\mathbf{II}^{3}$ is just a transpose of $\mathbf{II}^{2}$ and so
+and so the overall order is dominated by the first and third terms in (5.3.2). Further, note that $\mathbf{II}^{3}$ is just a transpose of $\mathbf{II}^{2}$ and so
 $\left\|\mathbf{II}^{3}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\mathbf{F}_{\mathbf{x}}^{0}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}})\overline{\mathbf{V}}_{M_{0}}^{0}\overline{\boldsymbol{\Gamma}}_{M_{0}}^{-1}\boldsymbol{\Gamma}_{i}\right\|=O_{p}(C_{N,T}^{-2}).$ (A.76)
@@
 Then
-$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+k(g-m)p_{N,T}$
+$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
 $=\ln\mathrm{det}\left(\mathbf{I}_{k}+O_{p}(C_{N,T}^{-2})\right)+k(g-m)p_{N,T}$
 $=\ln\mathrm{det}(\mathbf{I}_{k})+O_{p}(C_{N,T}^{-2})+k(g-m)p_{N,T}$
-$\to_{p}0,$ (A.75)
+$=O_{p}(C_{N,T}^{-2})+k(g-m)p_{N,T},$ (A.79)
 where the approximation comes from p. 119 in Paulsen, (1984):
 $\ln\mathrm{det}(\mathbf{I}+O_{p}(n^{-1}))=O_{p}(n^{-1})$
-for some integer $n$. However, given that $p_{N,T}O_{p}(C_{N,T}^{-2})\to\infty$, then we have that
+for some integer $n$. However, given that $p_{N,T}^{-1}O_{p}(C_{N,T}^{-2})=o(1)$, then we have that
 $p_{N,T}^{-1}(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0}))=k(g-m)+o_{p}(1)$ (A.80)
 and so
-$\mathbb{P}\left(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0}))<0)\right)\to 0,$ (A.77)
+$\mathbb{P}\left(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0}))<0\right)\to 0,$ (A.81)
 and so the overspecification risk is eliminated.
-6.5 Proof Changes for IC of Margaritella and Westerlund, (2023)
-Note that IC of MW is based on the residuals from the first stage CCE regression under the full set of available $k+1$ CAs. In particular, let us assume homogeneous $\boldsymbol{\beta}$. Then
+5.3.3 Proof Changes for IC of Margaritella and Westerlund, (2023)
+5.3.3.1  Homogeneous $\boldsymbol{\beta}$
+Note that IC of MW is based on the residuals from the first stage CCE regression under the full set of available $k+1$ CAs (note that $|M|=k+1$ now).
+In this case, we have $\widehat{\mathbf{F}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]=\mathbf{F}\overline{\mathbf{C}}+\overline{\mathbf{U}}$, where now $\overline{\mathbf{C}}$ is a loading matrix with rank of $m$ even when $N\to\infty$.
+In particular, let us assume homogeneous $\boldsymbol{\beta}$.
+Then
 $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}.$ (A.82)
 We then obtain
-$\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})$ $=\underbrace{\ln\left[1+\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}\right)\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}\right)^{-1}\right]}_{\mathcal{V}_{MW}-\mathcal{V}_{MW}^{0}}$
-$+(g-m)\cdot p_{N,T},$ (A.79)
+$\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})$ $=\ln\left[1+\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}\right)\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}\right)^{-1}\right]$
+$+(g-m)\cdot p_{N,T}$
+$=\ln\left[1+T^{\tau}\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}\right)\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}\right)^{-1}\right]$
+$+(g-m)\cdot p_{N,T}$ (A.83)
 where, by inserting (A.82) we further analyze
-$\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}$ $=\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
-$-\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
-$-\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
-$+\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}).$ (A.80)
+$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}$ $=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
+$-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
+$-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
+$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}).$ (A.84)
 Clearly,
-$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq(NT)^{-1}\underbrace{\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|}_{O_{p}(T^{\tau})}$
-=$=\begin{cases}O_{p}(N^{-1}T^{\tau-1})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.81)
-because the rate of $\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$ was already derived for the IC of DVS under both cases.
+$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ (A.85)
+$\leq(NT)^{-1}\underbrace{\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$
+$=\begin{cases}O_{p}((NT)^{-1})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.86)
+because the rate of $\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$ was already derived for the IC of DVS under both cases (either $O_{p}(1)$ or vanishing), and its behavior will not change when we use the full set of averages $\widehat{\mathbf{F}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]$.
 Also, $\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see Westerlund,, 2018), and thus the $\sqrt{NT}$-consistency will hold under our assumptions, as well.
 Similarly,
-$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|$ $\leq(NT)^{-1/2}\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$
-×$\times\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|}_{O_{p}(T^{\tau})}$
-=$=\begin{cases}O_{p}(N^{-1/2}T^{\tau-1/2}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.82)
+$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|$ $\leq(NT)^{-1/2}\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$
+$\times\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|$
+$=\begin{cases}O_{p}((NT)^{-1/2}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.87)
 because asymptotically $\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ behaves identically to $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, and thus the dominating component will have the same rate.
-Eventually,
-[…]
-which will be the dominating term.
+Eventually, by following the same analysis of $\mathbf{I},\mathbf{II}$ and $\mathbf{III}$ in (5.3.1), we have
+$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|=\begin{cases}O_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M.\end{cases}$ (A.88)
+In fact, we can demonstrate that in analogy to the analysis of (5.3.1), we can decompose and simplify the term:
+$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
+$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})+o_{p}(T^{-\tau})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})+o_{p}(T^{-\tau})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\gamma}_{i}^{\prime}\mathbf{F}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}\boldsymbol{\gamma}_{i}+o_{p}(T^{-\tau})$
+$=c_{NT}+o_{p}(T^{-\tau})$ (A.89)
+for a strictly positive $c_{NT}$ with $c_{NT}\to c>0$ under $M\subset M_{0}$.
+To complete, we will assume that $\tau\in(0,0.5)$, in order to isolate $o_{p}(T^{-\tau})$ terms. Then we have
+$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}=c_{NT}+o_{p}(T^{-\tau}),$ (A.90)
+and, because we can show that $b_{NT}:=\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}$ converges to a positive scalar quantity, we can analogously show that
+$(\mathrm{IC}^{MW}(M)$ $-\mathrm{IC}^{MW}(M_{0}))/\ln(T)=\ln(1+T^{\tau}(c_{NT}\times b_{NT}^{-1}+o_{p}(T^{-\tau}))/\ln(T)+o(1)$
+$=\ln(1+T^{\tau}c_{NT}\times b_{NT}^{-1})/\ln(T)+o_{p}(1)$
+$=\ln(T^{\tau}c\times b_{NT}^{-1})/\ln(T)+\left(\ln(1+T^{\tau}c_{NT}\times b_{NT}^{-1})-\ln(T^{\tau}c_{NT}\times b_{NT}^{-1}))\right)/\ln(T)$
+$+o_{p}(1)$
+$=\ln(T^{\tau}c_{NT}\times b_{NT}^{-1})/\ln(T)+\frac{1}{\ln(T)}\ln\left(1+\frac{b_{NT}}{cT^{\tau}}\right)+o_{p}(1)$
+$=\tau+\ln(c_{NT}\times b_{NT}^{-1})/\ln(T)+\frac{1}{\ln(T)}\left(\frac{b_{NT}}{c_{NT}T^{\tau}}+O(T^{-2\tau})\right)+o_{p}(1)$
+$\to_{p}\tau>0$ (A.91)
+as $(N,T)\to\infty$. Therefore,
+$\mathbb{P}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0)=\mathbb{P}\left(\frac{\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})}{\ln(T)}<0\right)\to 0$ (A.92)
+as $(N,T)\to\infty$.
+Under $M_{0}\subset M$, the proof goes exactly the same as in Margaritella and Westerlund, (2023),and so given the rates declared above and the fact that $p_{N,T}^{-1}O_{p}(C_{N,T}^{-2})=o(1)$, we obtain
+$p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))=(g-m)+o_{p}(1)$ (A.93)
+and so
+$\mathbb{P}\left(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))<0)\right)\to 0,$ (A.94)
+and the overspecification risk is eliminated.
+5.3.3.2  Heterogeneous $\boldsymbol{\beta}$
 If the parameter vector is heterogeneous, e.g. $\boldsymbol{\beta}_{i}=\boldsymbol{\beta}+\boldsymbol{\upsilon}_{i}$, where $\boldsymbol{\upsilon}_{i}$ is IID and mean-zero, then
-$\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},$ (A.84)
-where now $\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see e.g. Stauskas,, 2023). This means that now, for example,
-$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq N^{-1}\underbrace{\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|}_{O_{p}(T^{\tau})}$
-=$=\begin{cases}O_{p}(N^{-1}T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.85)
-and other terms will behave similarly, while
-$\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\upsilon}_{i}^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}\right\|=\begin{cases}O_{p}(T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.86)
-[…]
-as before, whereas the cross-terms are of lower order that could be made negligible by imposing restrictions on $N$ and $T$ divergence rate.
+$\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ $=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$
+$=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}\boldsymbol{\Gamma}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{V}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$
+$=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}(\boldsymbol{\Gamma}_{i}\boldsymbol{\upsilon}_{i}+\boldsymbol{\gamma}_{i})+\mathbf{V}_{i}\boldsymbol{\upsilon}_{i}+\boldsymbol{\varepsilon}_{i}$
+$=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*},$ (A.95)
+with the obvious definitions of $\boldsymbol{\gamma}_{i}^{*}$ and $\boldsymbol{\varepsilon}_{i}^{*}$.
+Note that now $\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see e.g.
+Stauskas,, 2023), which is a considerably lower rate.
+This means that now, for example,
+$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq N^{-1}\underbrace{\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$
+$=\begin{cases}O_{p}(N^{-1}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.96)
+and other terms will behave similarly:
+$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|$
+$\leq N^{-1/2}\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|$
+$=\begin{cases}O_{p}(N^{-1/2}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M.\end{cases}$ (A.97)
+Finally,
+$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|=\begin{cases}O_{p}(T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.98)
+and, in particular, under $M\subset M_{0}$:
+$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})+o_{p}(T^{-\tau})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})+o_{p}(T^{-\tau})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\gamma}_{i}^{*\prime}\mathbf{F}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+o_{p}(T^{-\tau})$
+$=d_{NT}+o_{p}(T^{-\tau})$ (A.99)
+with $d_{NT}\to_{p}d>0$ in analogy to the analysis in (5.3.3.1).
+Now, we need to assume $T^{\tau}N^{-1/2}\to 0$ in addition to $\tau\in(0,0.5)$ to isolate the $o_{p}(T^{-\tau})$ terms.
+Then in the under-specification case, we have
+$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}=d_{NT}+o_{p}(T^{-\tau}),$ (A.100)
+and, since again we can show that $e_{NT}:=\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}$ converges to a positive scalar quantity, we can demonstrate that
+$(\mathrm{IC}^{MW}(M)$ $-\mathrm{IC}^{MW}(M_{0}))/\ln(T)=\ln(1+T^{\tau}(d_{NT}\times e_{NT}^{-1}+o_{p}(T^{-\tau}))/\ln(T)+o(1)$
+$=\ln(1+T^{\tau}d_{NT}\times e_{NT}^{-1})/\ln(T)+o_{p}(1)$
+$\to_{p}\tau>0$ (A.101)
+as $(N,T)\to\infty$. Therefore,
+$\mathbb{P}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0)=\mathbb{P}\left(\frac{\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})}{\ln(T)}<0\right)\to 0$ (A.102)
+as $(N,T)\to\infty$.
+Under $M_{0}\subset M$, the proof once again the same as in Margaritella and Westerlund, (2023),and so given the rates declared above and the fact that $p_{N,T}^{-1}O_{p}(C_{N,T}^{-2})=o(1)$, we get
+$p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))=(g-m)+o_{p}(1)$ (A.103)
+and so
+$\mathbb{P}\left(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))<0)\right)\to 0,$ (A.104)
+as expected.
 5.4 Additional Monte Carlo Results
 5.4.1 Independent Idiosyncratics
+Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$
+Correct Selection Frequency for $g$
+50 50 31.00 15.90 36.40 16.10 7.30 3.60 3.50 0.30 0.10 0.00 0.00 0.00
+100 50 97.30 95.00 99.20 96.40 91.50 82.80 95.10 85.90 11.50 8.60 13.40 7.50
+200 50 95.50 94.30 99.70 98.80 48.60 45.50 77.70 69.20 0.20 0.20 5.50 4.10
+300 50 100.00 100.00 99.60 99.50 81.30 78.20 69.90 67.70 4.30 3.70 1.80 1.60
+500 50 100.00 100.00 100.00 100.00 99.60 99.40 96.30 95.10 21.20 20.10 11.50 10.10
+50 100 97.10 89.30 100.00 100.00 68.60 55.30 99.30 94.50 7.00 5.90 18.20 10.90
+100 100 100.00 99.90 99.90 99.00 62.50 49.40 56.30 39.70 2.70 1.30 1.70 0.60
+200 100 100.00 100.00 100.00 100.00 99.90 99.60 99.90 99.70 24.50 19.70 26.30 18.80
+300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 39.10 34.30 46.60 39.30
+500 100 100.00 100.00 100.00 100.00 99.70 99.70 95.90 94.80 16.30 15.30 8.50 7.60
+50 200 99.70 99.90 99.90 99.80 97.70 96.70 89.00 79.90 9.80 8.70 5.10 3.50
+100 200 100.00 100.00 100.00 100.00 93.00 88.80 99.50 98.60 5.30 4.70 12.80 7.70
+200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 40.80 32.70 60.20 44.50
+300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 22.40 18.90 81.40 73.20
+500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 54.10 49.30 80.80 75.50
+50 300 100.00 100.00 100.00 100.00 88.60 86.40 99.90 99.40 7.40 6.20 15.20 11.00
+100 300 98.90 99.80 100.00 100.00 98.70 99.40 100.00 100.00 21.30 18.60 27.20 22.40
+200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 56.00 49.10 52.60 45.60
+300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 89.30 82.90 94.40 89.10
+500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 80.10 75.70 93.70 91.90
+50 500 100.00 100.00 100.00 100.00 90.10 89.00 100.00 100.00 3.40 3.10 11.20 10.20
+100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 38.50 37.10 53.70 50.80
+200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 84.10 81.30 97.30 95.80
+300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 88.30 85.90 97.00 95.60
+500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 92.70 88.20 99.50 98.70
+Average of $g$
+50 50 3.16 2.48 3.14 1.96 2.41 2.02 2.27 1.68 1.51 1.38 1.40 1.23
+100 50 3.97 3.94 3.99 3.96 3.88 3.71 3.95 3.82 2.11 1.96 2.23 1.96
+200 50 3.96 3.94 3.99 3.98 3.43 3.38 3.56 3.38 1.86 1.82 1.61 1.54
+300 50 4.00 4.00 4.00 4.00 3.80 3.77 3.70 3.67 2.21 2.17 2.14 2.07
+500 50 4.00 4.00 4.00 4.00 4.00 3.99 3.96 3.95 2.65 2.61 2.37 2.32
+50 100 3.96 3.86 4.00 4.00 3.65 3.34 3.99 3.92 2.08 1.95 2.31 2.03
+100 100 4.00 4.00 4.00 3.99 3.45 3.19 3.51 3.21 1.45 1.33 1.73 1.52
+200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.69 2.54 2.62 2.41
+300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.14 3.05 3.21 3.06
+500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.95 2.63 2.58 2.76 2.70
+50 200 4.00 4.00 4.00 4.00 3.98 3.96 3.84 3.70 1.93 1.86 1.58 1.50
+100 200 4.00 4.00 4.00 4.00 3.92 3.87 3.99 3.98 2.09 2.02 1.95 1.77
+200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.16 2.98 3.36 3.04
+300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.01 2.92 3.61 3.44
+500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.21 3.11 3.69 3.56
+50 300 4.00 4.00 4.00 4.00 3.89 3.86 4.00 3.99 2.14 2.09 2.24 2.11
+100 300 4.01 4.00 4.00 4.00 4.01 4.00 4.00 4.00 2.83 2.75 2.69 2.57
+200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.50 3.38 3.48 3.38
+300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.85 3.74 3.94 3.87
+500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.80 3.75 3.93 3.91
+50 500 4.00 4.00 4.00 4.00 3.90 3.89 4.00 4.00 1.92 1.89 1.98 1.94
+100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.08 3.04 3.29 3.22
+200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.81 3.77 3.97 3.95
+300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.87 3.83 3.97 3.96
+500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.90 3.83 3.99 3.98
 Table 2: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.2) and (2.1).
-For $IC^{DVS}_{1}$ and $IC^{MW}_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $IC^{DVS}_{2}$ and $IC^{MW}_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text.
+For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
 Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
+Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
+Correct Selection Frequency for $g$
+50 50 82.60 74.70 37.60 16.30 26.30 17.90 2.60 0.40 0.10 0.00 0.00 0.00
+100 50 100.00 100.00 99.20 95.20 99.90 99.70 94.10 83.60 42.00 33.80 7.20 4.30
+200 50 100.00 100.00 99.60 98.70 99.90 99.90 75.70 67.40 28.20 24.80 2.30 1.40
+300 50 100.00 100.00 99.60 99.50 93.90 92.80 69.70 67.00 8.80 7.40 1.50 1.10
+500 50 100.00 100.00 100.00 100.00 100.00 100.00 96.20 94.90 49.80 47.70 9.70 8.50
+50 100 100.00 100.00 100.00 99.40 100.00 100.00 97.20 89.80 42.40 32.10 7.50 3.90
+100 100 100.00 100.00 99.90 98.40 90.50 79.90 47.90 31.80 2.80 1.70 0.40 0.20
+200 100 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 78.90 72.30 19.30 12.90
+300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 85.50 83.20 41.80 36.10
+500 100 100.00 100.00 100.00 100.00 100.00 99.80 96.00 94.70 27.10 25.20 7.40 6.80
+50 200 100.00 100.00 100.00 99.60 100.00 100.00 85.50 73.00 59.30 53.30 0.60 0.50
+100 200 100.00 100.00 100.00 100.00 100.00 100.00 99.50 98.40 49.80 42.90 5.50 3.80
+200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 95.20 92.20 49.70 36.30
+300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 98.80 96.90 77.10 67.70
+500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.70 76.40 69.70
+50 300 100.00 100.00 100.00 100.00 100.00 100.00 99.50 98.90 58.50 55.40 4.80 4.00
+100 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 85.50 82.80 16.70 12.50
+200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 91.50 88.50 48.10 40.50
+300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 91.40 85.60
+500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 93.10 90.90
+50 500 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.80 17.40 16.00 3.00 2.30
+100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 95.70 95.00 40.70 37.10
+200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 95.70 94.20
+300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 95.80 94.20
+500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.30 97.80
+Average of $g$
+50 50 3.83 3.74 3.07 1.87 3.15 2.97 2.19 1.63 1.85 1.71 1.32 1.20
+100 50 4.00 4.00 3.99 3.94 4.00 4.00 3.93 3.78 3.11 2.90 2.02 1.80
+200 50 4.00 4.00 3.99 3.97 4.00 4.00 3.52 3.35 2.48 2.38 1.58 1.52
+300 50 4.00 4.00 4.00 4.00 3.94 3.93 3.70 3.67 2.82 2.77 2.09 2.03
+500 50 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.95 3.38 3.35 2.33 2.27
+50 100 4.00 4.00 4.00 3.99 4.00 4.00 3.97 3.85 3.02 2.77 2.07 1.86
+100 100 4.00 4.00 4.00 3.98 3.90 3.79 3.40 3.07 1.85 1.70 1.45 1.33
+200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.69 3.58 2.48 2.27
+300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.84 3.81 3.15 3.01
+500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.95 3.20 3.17 2.69 2.63
+50 200 4.00 4.00 4.00 4.00 4.00 4.00 3.77 3.55 3.20 3.08 1.45 1.41
+100 200 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.98 3.10 2.93 1.82 1.71
+200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.94 3.89 3.19 2.90
+300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.94 3.54 3.35
+500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.62 3.48
+50 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.40 3.33 1.96 1.90
+100 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.78 3.73 2.53 2.42
+200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.88 3.43 3.32
+300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.91 3.83
+500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.89
+50 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.42 2.38 1.75 1.71
+100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.95 3.94 3.04 2.95
+200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.95 3.93
+300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.94
+500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.97
 Table 3: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2).
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text.
 For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
@@
 5.4.2 Non-Stationary Idiosyncratics in $\mathbf{X}$
+Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
+Correct Selection Frequency for $g$
+50 50 31.00 15.90 36.40 16.10 7.40 3.60 3.50 0.30 0.10 0.00 0.00 0.00
+100 50 97.30 95.00 99.20 96.40 91.50 82.40 95.10 86.20 12.20 9.10 13.90 7.60
+200 50 95.50 94.30 99.70 98.80 48.70 45.60 78.30 69.70 0.20 0.20 5.80 4.30
+300 50 100.00 100.00 99.60 99.50 81.20 78.20 69.70 67.50 3.90 3.40 1.60 1.40
+500 50 100.00 100.00 100.00 100.00 99.60 99.40 96.00 94.60 21.80 20.50 11.60 9.90
+50 100 97.10 89.30 100.00 100.00 68.20 54.70 99.40 94.40 7.00 5.70 18.30 10.80
+100 100 100.00 99.90 99.90 99.00 62.90 49.40 55.80 38.80 2.50 1.20 1.60 0.40
+200 100 100.00 100.00 100.00 100.00 99.90 99.60 99.90 99.70 24.20 19.20 25.80 18.60
+300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 39.50 34.70 46.90 39.00
+500 100 100.00 100.00 100.00 100.00 99.70 99.70 95.50 94.50 16.40 15.20 8.40 7.40
+50 200 99.70 99.90 99.90 99.80 97.80 96.70 88.70 79.60 9.90 8.80 5.50 3.70
+100 200 100.00 100.00 100.00 100.00 93.10 88.70 99.60 98.80 5.10 4.60 12.70 7.70
+200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 41.00 33.00 60.40 44.50
+300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 22.50 19.20 81.00 72.40
+500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 54.10 49.20 80.70 75.40
+50 300 100.00 100.00 100.00 100.00 88.90 86.70 100.00 99.50 7.10 6.00 14.10 10.20
+100 300 98.90 99.80 100.00 100.00 98.80 99.50 100.00 100.00 21.10 18.40 26.90 22.10
+200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 55.90 49.00 52.10 45.30
+300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 88.50 82.10 94.20 88.90
+500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 79.90 75.40 94.20 92.20
+50 500 100.00 100.00 100.00 100.00 89.70 88.70 100.00 100.00 3.40 3.00 10.60 9.40
+100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 39.30 37.80 53.00 50.00
+200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 84.00 81.00 97.40 96.00
+300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 89.00 86.50 97.20 95.90
+500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 92.50 87.90 99.40 98.40
+Average of $g$
+50 50 3.16 2.48 3.14 1.96 2.43 2.04 2.28 1.70 1.50 1.37 1.39 1.23
+100 50 3.97 3.94 3.99 3.96 3.88 3.70 3.95 3.82 2.13 1.96 2.25 1.96
+200 50 3.96 3.94 3.99 3.98 3.44 3.38 3.57 3.39 1.88 1.84 1.62 1.55
+300 50 4.00 4.00 4.00 4.00 3.80 3.77 3.70 3.67 2.19 2.15 2.12 2.06
+500 50 4.00 4.00 4.00 4.00 4.00 3.99 3.96 3.94 2.65 2.61 2.38 2.32
+50 100 3.96 3.86 4.00 4.00 3.64 3.33 3.99 3.92 2.06 1.94 2.30 2.02
+100 100 4.00 4.00 4.00 3.99 3.46 3.19 3.50 3.20 1.44 1.32 1.73 1.52
+200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.69 2.54 2.62 2.42
+300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.15 3.06 3.22 3.06
+500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.94 2.63 2.58 2.75 2.69
+50 200 4.00 4.00 4.00 4.00 3.98 3.96 3.84 3.69 1.94 1.86 1.60 1.51
+100 200 4.00 4.00 4.00 4.00 3.92 3.87 3.99 3.98 2.09 2.01 1.95 1.76
+200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.16 2.98 3.37 3.05
+300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.01 2.92 3.60 3.42
+500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.22 3.11 3.69 3.56
+50 300 4.00 4.00 4.00 4.00 3.89 3.87 4.00 3.99 2.12 2.07 2.21 2.09
+100 300 4.01 4.00 4.00 4.00 4.01 4.00 4.00 4.00 2.83 2.75 2.68 2.56
+200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.49 3.38 3.48 3.38
+300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.83 3.73 3.94 3.87
+500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.79 3.75 3.93 3.91
+50 500 4.00 4.00 4.00 4.00 3.90 3.88 4.00 4.00 1.93 1.89 1.97 1.92
+100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.09 3.05 3.29 3.22
+200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.81 3.77 3.97 3.95
+300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.87 3.84 3.97 3.96
+500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.90 3.83 3.99 3.98
 Table 4: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2).
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text.
 For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
@@
 $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
+Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
+Correct Selection Frequency for $g$
+50 50 82.60 74.70 37.60 16.30 26.70 18.10 2.60 0.40 0.10 0.00 0.00 0.00
+100 50 100.00 100.00 99.20 95.20 99.90 99.70 93.90 83.80 42.20 33.70 7.40 4.20
+200 50 100.00 100.00 99.60 98.70 99.90 99.90 76.20 68.00 29.00 25.90 2.60 1.60
+300 50 100.00 100.00 99.60 99.50 93.40 92.50 69.50 66.80 8.60 7.30 1.30 1.00
+500 50 100.00 100.00 100.00 100.00 100.00 100.00 95.80 94.50 50.70 48.70 9.80 8.50
+50 100 100.00 100.00 100.00 99.40 100.00 100.00 97.20 89.40 42.50 32.60 7.70 3.90
+100 100 100.00 100.00 99.90 98.40 90.30 80.00 47.40 31.50 2.50 1.40 0.20 0.10
+200 100 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 79.20 72.40 18.80 12.80
+300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 85.30 83.00 42.00 36.00
+500 100 100.00 100.00 100.00 100.00 100.00 99.80 95.60 94.40 27.30 25.30 7.20 6.80
+50 200 100.00 100.00 100.00 99.60 100.00 100.00 85.20 72.80 59.80 53.80 0.60 0.50
+100 200 100.00 100.00 100.00 100.00 100.00 100.00 99.60 98.50 50.20 43.00 5.40 3.80
+200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 94.90 92.00 49.50 36.40
+300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 98.70 97.00 76.60 66.70
+500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.70 76.10 69.60
+50 300 100.00 100.00 100.00 100.00 100.00 100.00 99.60 99.00 57.90 54.80 4.60 3.90
+100 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 84.50 81.80 16.30 12.00
+200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 91.50 88.60 47.70 40.60
+300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.60 91.40 85.30
+500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.60 93.60 91.10
+50 500 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.80 17.00 15.60 2.70 2.10
+100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 95.70 95.00 39.80 36.10
+200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 95.90 94.40
+300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 96.20 94.70
+500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.20 97.70
+Average of $g$
+50 50 3.83 3.74 3.07 1.87 3.15 2.98 2.20 1.64 1.85 1.70 1.31 1.20
+100 50 4.00 4.00 3.99 3.94 4.00 4.00 3.93 3.78 3.12 2.90 2.03 1.79
+200 50 4.00 4.00 3.99 3.97 4.00 4.00 3.53 3.36 2.49 2.41 1.60 1.53
+300 50 4.00 4.00 4.00 4.00 3.93 3.92 3.69 3.66 2.82 2.77 2.08 2.02
+500 50 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.94 3.39 3.35 2.34 2.28
+50 100 4.00 4.00 4.00 3.99 4.00 4.00 3.97 3.85 3.02 2.77 2.07 1.85
+100 100 4.00 4.00 4.00 3.98 3.90 3.79 3.40 3.07 1.85 1.69 1.44 1.32
+200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.70 3.59 2.48 2.27
+300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.84 3.81 3.16 3.02
+500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.94 3.21 3.17 2.68 2.62
+50 200 4.00 4.00 4.00 4.00 4.00 4.00 3.76 3.55 3.21 3.09 1.45 1.41
+100 200 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.98 3.11 2.94 1.82 1.72
+200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.93 3.89 3.19 2.91
+300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.94 3.53 3.33
+500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.61 3.48
+50 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.38 3.32 1.95 1.89
+100 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.76 3.72 2.52 2.40
+200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.88 3.43 3.32
+300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.91 3.83
+500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.93 3.89
+50 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.40 2.36 1.74 1.71
+100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.95 3.94 3.03 2.94
+200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.95 3.93
+300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.95
+500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.97
 Table 5: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2).
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text.
 For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
@@
 $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
-6.6.3 Objective Function Adjustment
-Table 7: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ DVS criteria from De Vos and Stauskas, (2024), see (2.2).
-IC objective function adjusted to $\mathrm{IC}^{DVS}(M_{\mathbf{x}})=\ln\det\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}\right)+g\cdot k\cdot p_{N,T}$.
-For $DVS_{1}$: $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
-Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
+5.4.3 Penalty by Bai, (2004)
+Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$
+Correct Selection Frequency for $g$
+50 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+300 50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+500 50 0.00 0.00 0.00 0.00 0.10 0.10 0.00 0.00 0.00 0.00 0.00 0.00
+50 100 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 100 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 100 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+300 100 0.00 0.00 0.00 0.00 0.10 0.10 0.00 0.00 0.00 0.00 0.00 0.00
+500 100 0.10 0.00 0.00 0.00 1.20 0.40 0.00 0.00 0.10 0.10 0.00 0.00
+50 200 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 200 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 200 0.00 0.00 0.00 0.00 0.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+300 200 2.40 0.00 0.00 0.00 8.80 2.70 0.00 0.00 0.10 0.00 0.00 0.00
+500 200 26.10 4.70 0.00 0.00 12.90 7.20 0.00 0.00 0.40 0.10 0.00 0.00
+50 300 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 300 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 300 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.10 0.10 0.00 0.00
+300 300 19.30 0.00 0.00 0.00 16.00 1.90 0.00 0.00 0.90 0.70 0.00 0.00
+500 300 99.90 92.90 0.00 0.00 72.70 53.60 1.20 0.00 2.70 1.30 0.00 0.00
+50 500 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+100 500 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+200 500 0.00 0.00 0.00 0.00 0.30 0.00 0.00 0.00 0.80 0.50 0.00 0.00
+300 500 15.80 0.10 0.00 0.00 14.90 3.10 0.00 0.00 0.90 0.70 0.00 0.00
+500 500 100.00 100.00 77.40 0.20 94.50 81.20 31.50 2.00 2.50 1.70 0.30 0.00
+Average of $g$
+50 50 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.02 1.01 1.00 1.00
+100 50 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.07 1.05 1.01 1.00
+200 50 1.00 1.00 1.00 1.00 1.03 1.02 1.00 1.00 1.09 1.08 1.00 1.00
+300 50 1.00 1.00 1.00 1.00 1.05 1.04 1.00 1.00 1.12 1.10 1.02 1.02
+500 50 1.00 1.00 1.00 1.00 1.03 1.02 1.00 1.00 1.20 1.18 1.04 1.04
+50 100 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.01 1.01 1.00 1.00
+100 100 1.00 1.00 1.00 1.00 1.07 1.04 1.00 1.00 1.01 1.00 1.00 1.00
+200 100 1.00 1.00 1.00 1.00 1.09 1.06 1.00 1.00 1.12 1.10 1.01 1.00
+300 100 1.00 1.00 1.00 1.00 1.05 1.03 1.00 1.00 1.21 1.18 1.06 1.04
+500 100 1.01 1.01 1.00 1.00 1.26 1.21 1.01 1.00 1.16 1.15 1.06 1.05
+50 200 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.01 1.01 1.00 1.00
+100 200 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.07 1.05 1.00 1.00
+200 200 1.00 1.00 1.00 1.00 1.23 1.11 1.00 1.00 1.23 1.17 1.08 1.04
+300 200 1.07 1.00 1.00 1.00 1.47 1.17 1.01 1.00 1.31 1.25 1.07 1.05
+500 200 1.80 1.15 1.00 1.00 1.80 1.55 1.16 1.11 1.25 1.23 1.06 1.06
+50 300 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.01 1.01 1.00 1.00
+100 300 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.07 1.05 1.01 1.00
+200 300 1.00 1.00 1.00 1.00 1.04 1.02 1.00 1.00 1.26 1.19 1.04 1.02
+300 300 1.61 1.00 1.00 1.00 1.82 1.24 1.02 1.00 1.39 1.31 1.07 1.04
+500 300 4.00 3.80 1.02 1.00 3.58 3.05 1.21 1.03 1.73 1.61 1.31 1.23
+50 500 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.01 1.01 1.00 1.00
+100 500 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.08 1.07 1.00 1.00
+200 500 1.09 1.06 1.00 1.00 1.01 1.00 1.00 1.00 1.30 1.25 1.08 1.06
+300 500 1.58 1.05 1.00 1.00 1.78 1.22 1.05 1.01 1.47 1.40 1.18 1.13
+500 500 4.00 4.00 3.65 1.06 3.90 3.60 2.52 1.52 1.67 1.56 1.28 1.19
+Table 6: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$.
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text.
+For $IC^{DVS}_{1}$ and $IC^{MW}_{1}$, $\tilde{p}_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})\ln(T)$; for $IC^{DVS}_{2}$ and $IC^{MW}_{2}$, $\widetilde{p}_{N,T}=\ln(T)p_{N,T}$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
+Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
 5.4.4 Correlated Idiosyncratics
-Table 8: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2).
+Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
+Correct Selection Frequency for $g$
+50 50 74.10 28.50 98.70 72.70 11.70 6.40 12.30 4.30 0.60 0.10 0.00 0.00
+100 50 94.80 93.30 99.40 97.10 94.00 89.00 99.30 97.20 23.50 15.90 43.20 26.90
+200 50 96.80 95.20 99.90 99.50 64.50 61.40 97.90 95.30 1.60 1.30 24.30 19.20
+300 50 100.00 100.00 100.00 100.00 90.00 88.30 91.00 89.60 10.30 8.80 12.30 10.20
+500 50 100.00 100.00 100.00 100.00 99.80 99.80 99.70 99.40 47.10 45.10 50.10 47.10
+50 100 77.30 86.90 98.80 96.10 69.10 58.00 99.90 98.70 7.70 5.80 33.00 21.60
+100 100 98.90 94.90 99.60 96.80 69.20 56.10 78.20 59.70 3.00 1.20 4.00 1.10
+200 100 99.70 100.00 100.00 100.00 99.90 99.70 99.90 99.90 34.70 27.80 52.90 40.20
+300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 54.60 48.80 77.30 70.90
+500 100 100.00 100.00 100.00 100.00 99.90 99.90 99.20 98.90 30.40 27.40 23.50 20.50
+50 200 99.30 99.00 100.00 100.00 94.50 94.10 95.70 89.80 9.80 8.50 8.70 5.60
+100 200 95.90 96.40 100.00 100.00 92.30 88.30 99.80 99.70 5.40 4.60 21.40 13.70
+200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 45.40 36.40 76.80 63.60
+300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 27.60 22.60 94.70 88.30
+500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 66.00 59.80 95.40 91.50
+50 300 99.10 99.60 100.00 100.00 85.40 82.60 100.00 99.90 6.70 5.60 18.70 15.70
+100 300 100.00 100.00 100.00 100.00 95.50 98.50 100.00 100.00 19.10 17.50 36.50 29.90
+200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 56.80 50.80 63.10 55.10
+300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 90.50 84.10 98.30 95.10
+500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 85.00 79.90 98.00 96.70
+50 500 100.00 100.00 100.00 100.00 85.10 83.10 100.00 100.00 2.80 2.30 12.30 11.30
+100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 36.80 34.40 60.60 55.60
+200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 83.10 80.80 98.60 97.80
+300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 88.60 84.70 99.10 98.10
+500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 93.60 88.70 99.90 99.70
+Average of $g$
+50 50 3.57 2.11 3.99 3.53 2.65 2.21 2.76 2.24 1.63 1.47 1.65 1.43
+100 50 4.00 3.91 3.99 3.95 3.94 3.84 3.99 3.97 2.52 2.26 3.10 2.71
+200 50 3.98 3.96 4.00 3.99 3.62 3.58 3.96 3.91 2.19 2.12 2.30 2.14
+300 50 4.00 4.00 4.00 4.00 3.90 3.88 3.91 3.90 2.62 2.55 2.81 2.72
+500 50 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.27 3.23 3.36 3.30
+50 100 4.22 4.01 3.99 3.95 3.68 3.39 4.00 3.98 2.17 2.00 2.72 2.40
+100 100 3.99 3.92 4.00 3.96 3.60 3.33 3.78 3.54 1.51 1.35 1.99 1.73
+200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.98 2.80 3.26 2.98
+300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.43 3.33 3.74 3.64
+500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.99 3.03 2.94 3.15 3.10
+50 200 4.00 3.99 4.00 4.00 3.98 3.95 3.94 3.86 1.93 1.85 1.74 1.62
+100 200 4.00 3.97 4.00 4.00 3.92 3.87 4.00 4.00 2.11 2.03 2.21 1.98
+200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.25 3.05 3.66 3.42
+300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.11 3.01 3.89 3.75
+500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.43 3.32 3.93 3.88
+50 300 4.01 4.00 4.00 4.00 3.85 3.83 4.00 4.00 2.08 2.02 2.36 2.26
+100 300 4.00 4.00 4.00 4.00 4.04 4.01 4.00 4.00 2.81 2.73 2.89 2.74
+200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.51 3.41 3.62 3.51
+300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.87 3.77 3.98 3.95
+500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.85 3.79 3.98 3.96
+50 500 4.00 4.00 4.00 4.00 3.84 3.82 4.00 4.00 1.85 1.81 2.04 1.99
+100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.03 2.97 3.43 3.33
+200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.80 3.76 3.98 3.97
+300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.87 3.82 3.99 3.98
+500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.84 4.00 4.00
+Table 7: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$.
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text.
 For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
 Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are weakly correlated time and units units.
+Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
+Correct Selection Frequency for $g$
+50 50 100.00 99.80 97.20 61.60 36.40 26.90 13.90 4.60 0.50 0.10 0.00 0.00
+100 50 100.00 99.80 99.30 96.40 99.80 99.70 99.10 96.30 63.70 53.20 38.00 23.50
+200 50 100.00 100.00 99.80 99.40 100.00 100.00 98.50 95.80 56.70 50.60 24.50 18.70
+300 50 100.00 100.00 100.00 100.00 97.20 96.70 91.80 89.70 29.60 26.10 12.50 10.70
+500 50 100.00 100.00 100.00 100.00 100.00 100.00 99.60 99.60 77.20 74.90 50.40 46.50
+50 100 100.00 100.00 97.10 90.80 100.00 99.90 99.70 97.10 50.10 38.70 19.60 11.20
+100 100 100.00 100.00 99.50 96.50 93.10 85.10 72.10 54.20 4.60 2.40 0.70 0.30
+200 100 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 89.00 84.60 49.20 36.50
+300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 93.80 91.10 77.30 70.00
+500 100 100.00 100.00 100.00 100.00 100.00 100.00 99.20 98.90 48.90 44.30 23.30 20.40
+50 200 100.00 100.00 100.00 99.90 100.00 100.00 94.00 87.50 57.70 52.20 2.40 1.80
+100 200 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.80 53.50 45.20 12.40 7.70
+200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 97.40 93.40 72.50 56.60
+300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.40 98.70 93.10 87.20
+500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.70 99.70 94.90 91.00
+50 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.60 54.70 50.90 7.60 6.20
+100 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 84.30 80.70 26.10 19.80
+200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 92.20 89.30 61.30 52.60
+300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 97.20 93.60
+500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.80 97.90 96.30
+50 500 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 14.70 13.60 4.00 3.50
+100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 93.90 93.20 47.60 43.90
+200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 98.20 96.60
+300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 98.50 97.00
+500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.50
+Average of $g$
+50 50 4.00 4.00 3.97 3.33 3.29 3.11 2.77 2.23 2.02 1.83 1.59 1.37
+100 50 4.00 4.00 3.99 3.94 4.00 4.00 3.99 3.96 3.51 3.32 3.00 2.61
+200 50 4.00 4.00 4.00 3.99 4.00 4.00 3.97 3.93 3.17 3.03 2.41 2.23
+300 50 4.00 4.00 4.00 4.00 3.97 3.97 3.92 3.90 3.22 3.17 2.83 2.75
+500 50 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.75 3.72 3.37 3.30
+50 100 4.00 4.00 3.97 3.88 4.00 4.00 4.00 3.97 3.18 2.93 2.50 2.20
+100 100 4.00 4.00 4.00 3.96 3.93 3.85 3.71 3.49 2.03 1.80 1.70 1.50
+200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.86 3.79 3.21 2.94
+300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.93 3.90 3.74 3.63
+500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.99 3.48 3.42 3.15 3.09
+50 200 4.00 4.00 4.00 4.00 4.00 4.00 3.91 3.81 3.17 3.04 1.56 1.51
+100 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.18 2.99 2.10 1.92
+200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.97 3.91 3.60 3.31
+300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.98 3.87 3.75
+500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.94 3.88
+50 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.32 3.24 2.08 2.01
+100 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.76 3.70 2.75 2.61
+200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.89 3.60 3.49
+300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.97 3.93
+500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.96
+50 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.31 2.27 1.81 1.78
+100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.93 3.92 3.20 3.12
+200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.96
+300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.97
+500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99
 Table 8: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2).
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text.
 For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$.
@@
 Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
-$\tau=0$ $\tau=0.1$ $\tau=0.9$
-$N$ $T$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$
-10 10 6.80 7.12 7.87 5.98 6.79 7.05 7.67 5.90 4.33 6.17 5.93 5.31
-20 10 7.21 7.41 8.18 5.95 7.14 7.42 8.11 5.81 5.43 6.94 7.05 5.75
-50 10 6.83 7.41 8.05 4.72 6.73 7.40 7.99 4.63 4.57 6.85 6.20 4.85
-100 10 6.71 7.41 7.62 3.88 6.69 7.40 7.73 3.86 3.98 6.34 5.67 3.76
-150 10 6.33 7.12 7.60 4.15 6.28 7.09 7.54 4.26 3.48 5.61 5.60 4.12
-200 10 6.02 6.68 7.73 4.11 5.97 6.62 7.65 4.07 3.09 4.75 5.27 3.78
-10 20 3.14 6.95 2.19 4.46 3.11 6.84 2.12 4.37 2.92 4.76 1.90 3.47
-20 20 6.01 7.62 3.41 3.00 5.94 7.62 3.58 2.99 3.48 6.87 2.38 3.07
-50 20 4.96 7.64 2.52 3.69 4.83 7.63 2.75 3.67 2.54 6.46 1.86 2.75
-100 20 4.14 7.33 3.35 4.00 4.05 7.37 3.24 3.89 1.57 5.73 1.48 3.19
-150 20 3.81 6.02 3.43 3.78 3.95 6.09 3.38 3.76 1.54 3.97 1.54 3.55
-200 20 3.99 4.92 3.69 3.92 4.02 4.90 3.82 3.97 1.43 3.06 1.68 3.40
-10 50 4.46 7.84 1.20 2.88 4.32 7.84 1.24 2.85 2.61 6.26 1.33 2.82
-20 50 6.90 7.94 1.66 2.92 6.88 7.93 1.70 2.88 4.16 7.35 1.73 2.89
-50 50 3.97 7.82 2.64 3.60 4.00 7.83 2.81 3.52 1.95 7.36 1.53 3.15
-100 50 3.69 5.26 2.66 4.51 3.55 5.55 2.58 4.35 1.63 4.66 1.39 2.96
-150 50 4.00 4.59 3.19 4.13 3.95 4.85 3.01 4.16 1.43 3.46 1.34 2.71
-200 50 4.00 4.02 3.28 2.92 4.00 4.03 3.23 2.97 1.29 2.50 1.37 2.80
-10 100 3.00 7.87 3.39 3.95 2.98 7.87 3.28 3.81 2.65 6.09 1.51 3.42
-20 100 6.03 7.99 1.19 3.71 6.27 7.99 1.23 3.53 3.91 7.38 1.53 2.84
-50 100 3.92 7.98 1.55 3.38 4.00 7.98 1.78 3.41 2.42 7.42 1.42 3.05
-100 100 4.00 5.80 2.44 3.00 4.00 5.37 2.36 2.95 1.67 5.22 1.49 2.66
-150 100 4.00 4.00 2.64 5.50 4.00 4.00 2.64 5.42 1.67 3.40 1.72 2.46
-200 100 4.00 4.00 4.08 4.76 4.00 4.00 4.08 4.24 1.54 2.70 1.59 2.23
-10 150 4.00 6.42 3.18 3.24 4.00 6.92 3.08 3.38 2.21 5.25 1.47 2.60
-20 150 6.42 8.00 1.05 4.77 6.49 8.00 1.12 4.57 5.17 7.41 1.66 3.23
-50 150 3.60 7.91 2.47 2.98 3.96 7.90 2.41 2.97 3.29 6.54 1.37 2.84
-100 150 4.00 4.52 2.87 4.89 4.00 4.29 2.73 4.92 1.66 5.51 1.67 2.36
-150 150 4.00 4.00 2.78 5.53 4.00 4.00 2.57 5.45 1.78 3.75 1.82 2.57
-200 150 4.00 4.00 4.51 5.24 4.00 4.00 4.47 5.23 1.79 2.88 1.69 2.34
-10 200 4.00 7.98 3.52 3.46 4.00 7.97 3.54 3.37 2.22 6.99 1.83 2.91
-20 200 4.06 8.00 2.69 3.77 4.12 8.00 2.79 3.65 4.38 7.42 1.90 2.81
-50 200 4.00 8.00 2.05 4.32 4.00 8.00 2.05 3.97 3.77 7.62 1.59 2.77
-100 200 4.00 4.24 2.05 5.25 4.00 4.23 1.94 5.30 1.93 5.17 1.63 2.49
-150 200 3.99 3.95 2.25 4.33 3.97 3.89 2.36 3.79 1.77 4.24 1.49 1.79
-200 200 4.00 4.00 4.26 5.50 4.00 4.00 4.29 5.46 1.78 2.64 1.76 2.48
+$\tau=0$ $\tau=0.4$ $\tau=0.9$
+[…]
+Correct Selection Frequency for $g$
+50 50 66.80 66.80 14.20 14.20 5.60 5.60 3.10 3.10 0.00 0.00 0.00 0.00
+100 50 99.80 99.80 51.10 51.10 95.80 95.80 37.80 37.80 0.80 0.80 1.20 1.20
+200 50 97.40 97.40 24.20 24.20 52.10 52.10 11.00 11.00 0.20 0.20 0.10 0.10
+300 50 100.00 100.00 17.40 17.40 80.90 80.90 14.90 14.90 1.70 1.70 0.90 0.90
+500 50 99.20 99.20 5.90 5.90 83.10 83.10 7.80 7.80 7.90 7.90 3.30 3.30
+50 100 100.00 100.00 52.70 52.70 98.70 98.70 37.40 37.40 1.00 1.00 1.10 1.10
+100 100 99.60 99.60 48.80 48.80 24.50 24.50 7.70 7.70 0.00 0.00 0.00 0.00
+200 100 100.00 100.00 42.10 42.10 98.00 98.00 25.70 25.70 2.10 2.10 2.00 2.00
+300 100 100.00 100.00 8.40 8.40 97.70 97.70 14.40 14.40 15.30 15.30 4.70 4.70
+500 100 99.70 99.70 0.40 0.40 68.20 68.20 4.80 4.80 3.10 3.10 1.50 1.50
+50 200 100.00 100.00 32.70 32.70 93.60 93.60 27.30 27.30 0.00 0.00 0.00 0.00
+100 200 100.00 100.00 21.20 21.20 94.50 94.50 25.80 25.80 0.00 0.00 0.00 0.00
+200 200 100.00 100.00 38.50 38.50 98.70 98.70 25.70 25.70 5.30 5.30 2.70 2.70
+300 200 100.00 100.00 1.70 1.70 93.50 93.50 7.60 7.60 7.00 7.00 2.20 2.20
+500 200 100.00 100.00 0.00 0.00 80.40 80.40 2.20 2.20 7.70 7.70 2.30 2.30
+50 300 100.00 100.00 40.70 40.70 99.60 99.60 32.30 32.30 0.00 0.00 0.90 0.90
+100 300 100.00 100.00 34.20 34.20 99.70 99.70 31.00 31.00 2.40 2.40 1.40 1.40
+200 300 100.00 100.00 29.70 29.70 100.00 100.00 26.40 26.40 19.40 19.40 5.40 5.40
+300 300 100.00 100.00 2.70 2.70 99.80 99.80 10.70 10.70 17.80 17.80 3.80 3.80
+500 300 100.00 100.00 0.00 0.00 97.50 97.50 0.30 0.30 26.90 26.90 5.90 5.90
+50 500 100.00 100.00 82.80 82.80 100.00 100.00 46.40 46.40 0.10 0.10 0.30 0.30
+100 500 100.00 100.00 28.30 28.30 100.00 100.00 36.20 36.20 6.80 6.80 3.80 3.80
+200 500 99.10 99.10 16.60 16.60 100.00 100.00 26.10 26.10 37.40 37.40 7.40 7.40
+300 500 100.00 100.00 0.70 0.70 100.00 100.00 6.30 6.30 32.80 32.80 8.00 8.00
+500 500 100.00 100.00 0.00 0.00 94.70 94.70 0.20 0.20 17.80 17.80 3.10 3.10
+Average of $g$
+50 50 3.78 3.78 1.97 1.97 3.02 3.02 2.05 2.05 2.64 2.64 1.45 1.45
+100 50 4.00 4.00 2.98 2.98 3.94 3.94 2.92 2.92 1.49 1.49 1.65 1.65
+200 50 3.97 3.97 3.20 3.20 3.40 3.40 2.49 2.49 1.49 1.49 1.48 1.48
+300 50 4.00 4.00 4.73 4.73 3.92 3.92 4.23 4.23 1.50 1.50 1.58 1.58
+500 50 4.01 4.01 4.94 4.94 4.16 4.16 4.89 4.89 1.72 1.72 1.74 1.74
+50 100 4.00 4.00 2.69 2.69 3.99 3.99 2.71 2.71 1.98 1.98 1.66 1.66
+100 100 4.00 4.00 3.08 3.08 2.30 2.30 2.16 2.16 2.05 2.05 1.38 1.38
+200 100 4.00 4.00 3.06 3.06 3.98 3.98 3.22 3.22 1.50 1.50 1.54 1.54
+300 100 4.00 4.00 4.89 4.89 4.02 4.02 4.73 4.73 1.90 1.90 1.81 1.81
+500 100 4.00 4.00 5.00 5.00 4.21 4.21 4.91 4.91 1.59 1.59 1.54 1.54
+50 200 4.00 4.00 1.98 1.98 3.94 3.94 2.26 2.26 3.40 3.40 1.54 1.54
+100 200 4.00 4.00 1.64 1.64 3.91 3.91 2.21 2.21 2.04 2.04 1.45 1.45
+200 200 4.00 4.00 3.44 3.44 4.00 4.00 3.81 3.81 1.54 1.54 1.66 1.66
+300 200 4.00 4.00 4.98 4.98 4.03 4.03 4.90 4.90 1.97 1.97 1.72 1.72
+500 200 4.00 4.00 5.00 5.00 4.18 4.18 4.97 4.97 1.77 1.77 1.70 1.70
+50 300 4.00 4.00 2.25 2.25 3.99 3.99 2.40 2.40 2.56 2.56 1.62 1.62
+100 300 4.00 4.00 2.03 2.03 4.00 4.00 2.23 2.23 1.91 1.91 1.75 1.75
+200 300 4.00 4.00 3.12 3.12 4.00 4.00 3.56 3.56 2.03 2.03 1.78 1.78
+300 300 4.00 4.00 4.97 4.97 4.00 4.00 4.88 4.88 2.18 2.18 1.96 1.96
+500 300 4.00 4.00 5.00 5.00 4.03 4.03 5.00 5.00 3.00 3.00 2.93 2.93
+50 500 4.00 4.00 3.48 3.48 4.00 4.00 2.90 2.90 2.90 2.90 1.60 1.60
+100 500 4.00 4.00 1.85 1.85 4.00 4.00 2.12 2.12 2.00 2.00 1.81 1.81
+200 500 3.99 3.99 4.25 4.25 4.00 4.00 4.36 4.36 2.59 2.59 2.03 2.03
+300 500 4.00 4.00 4.99 4.99 4.00 4.00 4.94 4.94 2.68 2.68 2.40 2.40
+500 500 4.00 4.00 5.00 5.00 4.05 4.05 5.00 5.00 2.42 2.42 2.25 2.25
 Table 9: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$.

### v4

@@
 As CAs can proxy non-stationary factors, it is tempting to claim an excellent performance of ICs under general factors, too.
-We show formally and in simulations that they remain consistent, but the more persistent factors are, the poorer they perform in small samples, which goes against the sentiment in the CCE literature.
+We show formally and in simulations that they remain consistent, but the more persistent factors are, the poorer they perform in small samples, which goes against the sentiment in the CCE/CAs literature.
 Abstract
@@
 where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is an error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ are latent common factors, where $m$ denotes the number of factors, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix.
-The most popular setting nested in (1.1) is that of interactive effects.
-Then $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, and $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ gives a model with unobserved heterogeneity, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables.
-Naturally, $K=k+1$, and $i$ indexes individuals, such as countries or households.
-The unobserved component of $\mathbf{y}_{i}$ is likely to be correlated with $\mathbf{X}_{i}$ (see Bai, 2009a, ).
-If $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, (1.1) represents the Common Correlated Effects (CCE) setting of Pesaran, (2006), where $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter of interest, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings).
-Since both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, we have endogeneity and cross-section dependence among individuals that make the least squares (LS) estimator of $\boldsymbol{\beta}$ inconsistent.
+(1.1) nests several settings.
+The most popular one is that of interactive effects (see Bai, 2009a, ).
+Then $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]\in\mathbb{R}^{T\times(k+1)}$, where $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ gives a model with unobserved heterogeneity.
+If $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}\in\mathbb{R}^{T\times k}$, we have the Common Correlated Effects (CCE) setting of Pesaran, (2006) with $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i},\boldsymbol{\gamma}_{i}$) representing idiosyncratics (loadings).
+(1.1) goes beyond interactive effects and CCE.
+Stauskas and De Vos, (2025) consider Distinct Correlated Effects (DCE), where $\mathbf{y}_{i}$ and $\mathbf{X}_{i}$ load on distinct sets of factors.
+Massacci and Kapetanios, (2024) use $\mathbf{Z}_{i}=\mathbf{X}_{i}$ to model a set of co-moving regressors in a forecasting equation.
 It is possible to estimate unobserved factors in several ways, including the Principal Components (PC) method (see Bai and Ng,, 2006) or diversified projections (see Fan and Liao,, 2022).
-Pesaran, (2006) suggests a simple and elegant way to proxy $\mathbf{F}$ up to a linear transformation by taking cross-section averages (CAs): $\overline{\mathbf{Z}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]=\widehat{\mathbf{F}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g.
+Pesaran, (2006) suggests a simple and elegant way to proxy $\mathbf{F}$ up to a linear transformation by taking cross-section averages (CAs): $\overline{\mathbf{Z}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}=\widehat{\mathbf{F}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g.
 Pesaran and Tosetti,, 2011).
-The CCE estimator is then given by
-$\widehat{\boldsymbol{\beta}}_{CCE}=\overline{\mathbf{Q}}^{-1}\overline{\mathbf{q}},$ (1.2)
-where $\overline{\mathbf{Q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{X}_{i}$, $\overline{\mathbf{q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{y}_{i}$, $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix and $\mathbf{A}^{+}$ is the Moore-Penrose inverse.
-The core principle of CCE is to project the variables onto the orthogonal complement of their CAs in order to purge the factors asymptotically.
-Note that CAs are applicable beyond the usual interactive effects and CCE.
-Stauskas and De Vos, (2025) consider Distinct Correlated Effects (DCE), where $\mathbf{y}_{i}$ and $\mathbf{X}_{i}$ load on distinct sets of factors, and (1.2) is aided by bootstrap.
-Massacci and Kapetanios, (2024) use $\overline{\mathbf{X}}$ to approximate factors and produce factor-augmented forecasts.
+This is enough to estimate $\mathbf{F}$ in DCE and forecasting settings.
+In interactive effects, the CCE estimator of $\boldsymbol{\beta}$ is then simply the least squares (LS) estimator augmented with $\overline{\mathbf{Z}}=\widehat{\mathbf{F}}$ as an additional regressor.
 A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties.
@@
 This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs.
-However, there are reasons to ensure that $m=k+1$.
-If $m<k+1$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017).
+However, there are reasons to ensure that the number of CAs matches $m$.
+In CCE, if $m<k+1$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017).
 This situation is common in macro datasets (see De Vos and Stauskas,, 2024).
@@
 In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion).
-Beyond CCE, Stauskas and De Vos, (2025) show that DCE only uses $\overline{\mathbf{X}}$, but $m=k$ must hold.
+Beyond CCE, Stauskas and De Vos, (2025) show that DCE only uses $\overline{\mathbf{Z}}=\overline{\mathbf{X}}$, but $m=k$ must hold.
 In forecasting, the same is needed to avoid only conservative confidence intervals around factor-augmented forecasts (see Karabiyik and Westerlund,, 2021).
-In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that the information criteria (ICs) of Bai and Ng, (2002) and Bai, 2009b from the PC literature can be applied in the CCE setting under stationary factors to identify an optimal set of CAs from $\overline{\mathbf{Z}}$.
-De Vos and Stauskas, (2024) (DVS) show similar results when CAs are selected from $\overline{\mathbf{X}}$, which is more applicable in the DCE and forecasting settings.
-[^1] Even in basic CCE setting, it is advised to omit $\overline{\mathbf{y}}$ and use $\overline{\mathbf{X}}$ only to avoid computational issues (see Karavias et al.,, 2023)., [^2] PC techniques are limited in CCE, as they focus solely on $m$ by detecting the largest eigenvalues of the data matrix.
+In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that the information criteria (ICs) of Bai and Ng, (2002) and Bai, 2009b from the PC literature can be applied in the pure CCE setting under stationary factors to identify an optimal set of CAs from $\overline{\mathbf{Z}}$.
+De Vos and Stauskas, (2024) (DVS) show similar results when CAs are selected from $\overline{\mathbf{X}}$, which is more applicable in the DCE and forecasting settings, but the procedure has a similar theoretical basis.
+[^1] Even in basic CCE setting, it is advised to omit $\overline{\mathbf{y}}$ and use $\overline{\mathbf{X}}$ only to avoid computational issues (see Karavias et al.,, 2023)., [^2] PC techniques are limited in CAs setting, as they focus on $m$ by detecting the largest eigenvalues of the data matrix.
 They cannot detect sets of CAs as they do not have a natural ordering.
@@
 MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC.
-As CAs proxy a general factor structure, it is natural to evaluate this statement and re-visit the ICs proposed in MW and DVS.
+As CAs proxy a general factor structure, it is natural to evaluate this statement and re-visit the ICs proposed in both MW and DVS due to their wide applicability and similar theoretical foundation.
+The subtlety arises here because IC is minimized by grid-search at different combinations of CAs.
+Inevitably, there is a need to understand the asymptotic behavior of IC evaluated at a combination inconsistent for $\mathbf{F}$ that is non-stationary, which is a new undertaking in the CAs literature.
 Therefore, this study can be seen as the CAs counterpart of Bai, (2004), where classical PC results of Bai and Ng, (2002) are evaluated against pure unit root factors.
@@
 We discuss differences between our results and those in Bai, (2004), and also explore penalty functions adapted to non-stationary $\mathbf{F}$, as suggested in the latter study.
-Since they render our ICs inconsistent, we argue that practitioners should not mechanically take recommendations from the PC literature if the goal is to select CAs.
+Since they make the performance of our ICs even worse, we argue that practitioners should not mechanically take recommendations from the PC literature if the goal is to select CAs.
 More importantly, the ICs of DVS and MW (with the original penalties) should be applied in relatively large samples if the presence of non-stationary factors is suspected.
 2 Econometric Setup
-We focus on the IC of DVS due to its wider applicability.
-While we provide numerical experiments on both DVS and MW, theoretical arguments regarding the latter are nearly identical, and we relegate them to Section 3.3 of the Supplementary material.
+We focus on the IC of DVS to analyze our theoretical results, however, the conclusions apply to MW, as well.
+While we provide numerical experiments on both DVS and MW for comparison, theoretical arguments regarding the latter are similar, and we relegate them to Section 3.3 of the Supplementary material.
 To operationalize our analysis, we introduce $M$ which is a set of column indices of $\overline{\mathbf{X}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ picks the corresponding $g$ averages in practice.
-For instance, $\overline{\mathbf{X}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k$ CAs.
+That is, $\overline{\mathbf{X}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ defines a selection of $g$ out of $k$ CAs.
 Consequently, let $M_{0}$ denote the true set of averages from $\overline{\mathbf{X}}$ such that $\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=m$, and $|M_{0}|=m$, where $|M|$ denotes the cardinality of an arbitrary set $M$.
@@
 where $\ln(.)$ denotes the natural logarithm and $\mathrm{det}(\mathbf{A})$ is a determinant of any square matrix $\mathbf{A}$.
-$\overline{\mathbf{Q}}_{M}$ is $\overline{\mathbf{Q}}$ from (1.2) evaluated at $\widehat{\mathbf{F}}_{M}$, and $p_{N,T}$ is a penalty term.
-[^3] In case of MW, $\overline{\mathbf{Q}}_{M}=\ln\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}\right)$, where $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ and $\widehat{\boldsymbol{\beta}}$ is obtained using the CCE estimator under all available $k+1$ CAs.
+$\overline{\mathbf{Q}}_{M}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\mathbf{X}_{i}$, $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix, $\mathbf{A}^{+}$ is the Moore-Penrose inverse, and $p_{N,T}$ is a penalty term.
+[^3] In case of MW, $\overline{\mathbf{Q}}_{M}=\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}$ (scalar), where $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ and $\widehat{\boldsymbol{\beta}}$ is obtained using the CCE estimator under all available $k+1$ CAs.
+Then $\mathrm{IC}(M)=\ln(\overline{\mathbf{Q}}_{M})+g\cdot p_{N,T}$.
 Examples of feasible and most popular penalties are given by
@@
 for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002).
-Note that (2.1) is function of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
+Note that (2.1) is function of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors as long as CAs are (rotationally) consistent for $\mathbf{F}$ (see Theorem 1 in Westerlund,, 2018).
 Let $\overline{M}$ be the set of indices of all available CAs that, so that $|\overline{M}|=k$. Then
@@
 where $|\widehat{M}|=g$ provides the estimator of the number of factors as a consequence.
-In DVS, it is demonstrated that $\mathbb{P}(\widehat{M}=M_{0})\to 1$ as $(N,T)\to\infty$.
-This result theoretically justifies the use of IC previously very popular in the PC literature to select the number of factors.
+In DVS (and MW), it is demonstrated that $\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0\right)\to 0$ (therefore, $\mathbb{P}(\widehat{M}=M_{0})\to 1$) as $(N,T)\to\infty$.
+This result means that some other $M$ does not minimize IC asymptotically and theoretically justifies its use in the case of CAs.
 These findings are based on stationary $\mathbf{F}$, which has been considerably relaxed in the CCE/CAs literature.
-Therefore, our natural goal is to examine (2.1) under departures from this restrictive assumption.
+Therefore, our natural goal is to examine (2.1) in a way that departs from this restrictive assumption.
 Throughout our analysis, we employ the following set of assumptions.
-Assumption 1. $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009).
-  • (a)
-    In particular,
-    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}>0\hskip 5.69054pt(diagonal),$
+Assumption 1. $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009), such that
+$\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=\mathrm{diag}(g_{1},\ldots,g_{m}),\hskip 5.69054ptg_{j}\in(0,2),$
 where $\mathbf{u}_{f,t}$ is a zero-mean linear process.
-  • (b)
-    If $\tau=0$, then $\{\mathbf{f}_{t}\}$ is a stationary process with summable autocovariances.
 Assumption 2. Let $\mathbf{e}_{i,t}=(\varepsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}\in\mathbb{R}^{k+1}$. Then
@@
 We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allow us to vary $\tau$ and give theoretical guarantees, such as $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$, which is a constant positive definite matrix.
-Part (b) ensures that we return to the usual stationarity conditions under $\tau=0$.
-Specifically, we need the eigenvalues of $\mathbf{I}_{m}-\mathbf{G}$ to be within a unit circle, so that the process can be inverted and admit a representation of $MA(\infty)$.
+The current specification ensures that we return to the usual stationarity conditions under $\tau=0$, as the eigenvalues of $\mathbf{I}_{m}-\mathbf{G}$ lie within a unit circle, so that the process can be inverted and admit a representation of $MA(\infty)$.
 Assumption 2 (a) is split into two parts.
@@
 In our assumptions, we cover $k+1$ variables in order to accommodate MW, as well.
-To formulate our instrumental theoretical result, let $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\mathrm{det(\overline{\mathbf{Q}}_{M})}-\ln\mathrm{det(\overline{\mathbf{Q}}_{M_{0}})}$.
-This quantity is important because it allows us to track the asymptotic contribution of factor estimates to the minimization problem under $M_{0}\subset M$ (over-specification) and $M\subset M_{0}$ (under-specification).
-Crucially, the behavior of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ prescribes properties of $p_{N,T}$.
-To illustrate this, Bai, (2004) (in PC setting) shows that under over-specification, (an equivalent of) $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is $O_{p}(1)$ when factors are non-stationary, and so the contribution of the extra estimates to the sum of squared residuals is non-negligible.
-This leads to $p_{N,T}\to\infty$ to heavily penalize redundant estimates.
-However, when factors are stationary, Bai and Ng, (2002) (and DVS) detect $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$, which means that $p_{N,T}\to 0$ at a slower rate in order to penalize lightly.
-Proposition 1 below can be seen as the CAs equivalent of Lemma A3 and Lemma A4 of Bai, (2004).444 Strictly speaking, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$.
+Let $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\mathrm{det}(\overline{\mathbf{Q}}_{M})-\ln\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}})=\ln\det\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)$, where $\boldsymbol{\Omega}_{M,M_{0}}:=T^{-\tau}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}$.
+$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is important since it helps describing the behavior of $\mathrm{IC}(M)$ in the neighborhood of $\mathrm{IC}(M_{0})$ under $M_{0}\subset M$ (over-specification) and $M\subset M_{0}$ (under-specification).
+In addition, it prescribes properties of $p_{N,T}$, which can be seen from
+$\mathrm{IC}(M)-\mathrm{IC}(M_{0})=\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}+k\cdot(g-m)p_{N,T}.$ (2.4)
+To illustrate, Bai, (2004) (in PC setting) shows that under $M_{0}\subset M$, (an equivalent of) $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is $O_{p}(1)$ when factors are non-stationary, and so the contribution of the extra estimates to the sum of squared residuals is non-negligible.
+This leads to $p_{N,T}\to\infty$ to heavily penalize redundant estimates so that $\mathrm{IC}(M)-\mathrm{IC}(M_{0})>0$.
+However, when factors are stationary, Bai and Ng, (2002) (and DVS) detect $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$, which means that $p_{N,T}\to 0$ at a slower rate to penalize lightly to ensure $\mathrm{IC}(M)-\mathrm{IC}(M_{0})>0$ asymptotically.
+Proposition 1 below can be seen as the CAs equivalent of Lemma A3 and A4 of Bai, (2004).
+[^4] Strictly speaking, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$.
 We only analyze the case of $M\subset M_{0}$ for brevity.
-As Margaritella and Westerlund, (2023) pointed out analysis would lead to the same conclusion in all the cases.
+As Margaritella and Westerlund, (2023) pointed out, analysis would lead to the same conclusion in all cases.
 Proposition 1. Under Assumptions 1-4 as $(N,T)\to\infty$, we have that:
-  • (a)
-    Under $M\subset M_{0}$, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}+\infty$;
-  • (b)
-    Under $M_{0}\subset M$, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2})$.
-Proof: Section 3 of Supplementary material.
-Let us discuss Proposition 1 and start with (b).
-In this case, the CAs are able to approximate all $m$ factors under Assumption 4.
+$\textit{(a)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM_{0}\subset M,\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2});$ (2.5)
+$\textit{(b)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM\subset M_{0},\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\tau b\ln(T)+\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)+R_{N,T},$ (2.6)
+where $\boldsymbol{\Omega}_{M.M_{0}}^{0}=\operatorname*{plim}_{(N,T)\to\infty}\boldsymbol{\Omega}_{M.M_{0}}$ (positive semi-definite, non-zero), $|R_{N,T}|=o_{p}(1)$ is the remainder, $\lambda_{j}(\mathbf{A})$ is the $j$-th eigenvalue of $\mathbf{A}$, while $b$ is a number of strictly positive eigenvalues.
+Proof: Section 3 of the Supplementary material.
+In (a), the CAs are able to approximate all $m$ factors under Assumption 4.
 Then $\overline{\mathbf{Q}}_{M}$ is exactly (log-determinant of) the denominator of the CCE estimator.
-It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}+o_{p}(1)$ for fixed $T$, and by inserting this into $\mathbf{X}_{i}$, we can show that
-$\ln\mathrm{det}(\overline{\mathbf{Q}}_{M})=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}\right)+o_{p}(1)=\ln\mathrm{det}\left(\boldsymbol{\Sigma}_{\mathbf{v}}\right)+o_{p}(1),$ (2.4)
-whose argument is a positive definite matrix under general factors (see also Westerlund,, 2018).
+It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}+o_{p}(1)$ for fixed $T$, and by inserting this into $\mathbf{X}_{i}$, we can show that $\ln\mathrm{det}(\overline{\mathbf{Q}}_{M})=\ln\mathrm{det}\left(\boldsymbol{\Sigma}_{\mathbf{v}}\right)+o_{p}(1)$, whose argument is a positive definite matrix under general factors (see also Westerlund,, 2018).
 Given that $\ln\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}})$ admits the same asymptotic representation, it is natural that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is negligible.
-Moreover, the rate is already detected by DVS and Bai and Ng, (2002) (for PC) under stationary $\mathbf{F}$, and it remains identical here.
+Moreover, the rate is already detected by DVS, MW and Bai and Ng, (2002) (for PC) under stationary $\mathbf{F}$, and it remains identical here.
 This rate is instructive and determines the properties of penalty $p_{N,T}$, which is the key motivation behind the functional forms in (2.2).
-Specifically, because redundant gmg-m CAs have an asymptotically negligible contribution to the sum of squared residuals, over-specification does not need to be penalized heavy, and so it is sufficient that $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$ as $(N,T)\to\infty$.
+Specifically, because redundant gmg-m CAs have an asymptotically negligible contribution to the sum of squared residuals, over-specification does not need to be heavily penalized, and so (2.4) becomes
+$\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})\right)/p_{N,T}\to_{p}k\cdot(g-m)>0$ (2.7)
+as $(N,T)\to\infty$ if $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$.
 Note that this stands in sharp contrast to findings in Lemma A4 of Bai, (2004), where excess factor estimates contribute non-trivially.
-Overall, the usual consistency of the CCE estimator in (1.2) under general unknown factors prevails when $M_{0}\subset M$ (see also Westerlund,, 2018, or Stauskas,, 2023).
-Under $M\subset M_{0}$ in (a), the rank condition in Assumption 4 is not satisfied.
-This means that mgm-g factors are not asymptotically projected out, and the representation in (2.4) is invalid.
-Note that $\overline{\mathbf{Q}}_{M}$ is a quadratic form in $\mathbf{F}$, and we know that $\left\|\mathbf{F}^{\prime}\mathbf{F}\right\|=O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a).
+Overall, the usual consistency of CAs under general unknown factors prevails when $M_{0}\subset M$ (see also Westerlund,, 2018, or Stauskas,, 2023).
+However, the result is more nuanced under $M\subset M_{0}$.
+In (b), the rank condition in Assumption 4 is not satisfied.
+This means that CAs are inconsistent for the mgm-g factors.
+Note that $\overline{\mathbf{Q}}_{M}$ is a quadratic form in $\mathbf{F}$, and we know that $\left\|\mathbf{F}^{\prime}\mathbf{F}\right\|=O_{p}(T^{1+\tau})$ as implied by Assumption 1.
 However, the objective function offers normalization by $T$ only as the integration order is unknown.
-Exactly this imbalance causes divergence of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$.
-A similar situation arises in Lemma A3 of Bai, (2004), where the central step is to detect the sign of divergence, and we follow this path.
-Note that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\det\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)$, where $\boldsymbol{\Omega}_{M,M_{0}}:=T^{-\tau}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}\geq 0$ (not zero) as $(N,T)\to\infty$, which we demonstrate in the Supplement.
-Using this, the difference can be simplified to
-$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\tau b\ln(T)+\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}})\right)+R_{N,T},$ (2.5)
-where $|R_{N,T}|=o_{p}(1)$ is the remainder, $\lambda_{j}(\mathbf{A})$ is the $j$-th eigenvalue of $\mathbf{A}$, while $b$ is a number of strictly positive eigenvalues.
-Because $b$ is finite ($b\leq k$), it is clear that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to+\infty$ at a rate of $O_{p}(\ln(T))$.
+Exactly this imbalance causes the divergence of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$.
+A similar situation arises in Lemma A3 of Bai, (2004), where the central step is to detect the sign of divergence, and we follow this route.
+In (A.9), $b$ is finite ($b<k$), so $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}+\infty$ and therefore
+$\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})\right)/\ln(T)\to_{p}\tau b>0,$ (2.8)
+as $(N,T)\to\infty$, as required.
 Two comments are in order.
-Firstly, in the Supplement we show that $|R_{N,T}|=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$, and it is not necessarily positive.
-This means that for $\tau\approx 1$ and a small $N$, the remainder vanishes slowly as $T\to\infty$.
-Hence, it may be necessary to rely on large $N,T$ combinations to detect the minimum of IC outside of $M\subset M_{0}$ region if $\mathbf{F}$ is very persistent.
-We explore (and confirm) such risks in Monte Carlo simulations in Section 3.
-Secondly, while this result is similar to Lemma A3 in Bai, (2004), there the divergence rate is $O_{p}(T/\ln\ln(T))$, since the law of iterated logarithm is used to arrive at the expression analogous to (2.5) when factors have unit root.
+Firstly, neither $\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)$ nor $R_{N,T}$ are guaranteed to be positive.
+In addition, in the Supplement, we show that $|R_{N,T}|=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$, which means that for $\tau\approx 1$ and a small $N$, it vanishes slowly as $T\to\infty$.
+This can induce $\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0$ in small samples.
+Hence, it may be necessary to rely on large $N,T$ combinations so that $\tau b\ln(T)$ dominates to detect the minimum of IC outside of $M\subset M_{0}$ region if $\mathbf{F}$ is very persistent.
+We explore and confirm such risks in Monte Carlo simulations in Section 3, where we see that both MW and DVS misselect (in fact, underselect), unless both $N$ and $T$ are large.
+Secondly, while this result is similar to Lemma A3 in Bai, (2004), there the divergence rate is $O_{p}(T/\ln\ln(T))$, since the law of iterated logarithm is used to arrive at the expression analogous to (A.9) when factors have unit root.
 Theorem 1 leads to our consistency result.
-Theorem 1. Under conditions of Proposition 1 with $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$, we have
-$\mathbb{P}\left(\widehat{M}=M_{0}\right)\to 1$
-as $(N,T)\to\infty$.
-Proof. Follows directly from Proposition 1, because $\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})>0\right)\to 1$ under (a)-(b) and the penalty properties.
+Theorem 1. Under conditions of Proposition 1 with $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$, we have as $(N,T)\to\infty$
+$\mathbb{P}\left(\widehat{M}=M_{0}\right)\to 1.$
+Proof.
+Follows from Proposition 1, because $\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0\right)=\mathbb{P}\left(\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})\right)/p_{N,T}<0\right)\to 0$ under (a) and a similar statement holds under (b).
 The key message of Theorem 1 is that the IC of DVS (and MW) is consistent, but exhibits hybrid asymptotic properties characteristic to stationary and non-stationary factors.
-Consequently, two forces are in effect: $p_{N,T}$ should still be negligible, but the factor estimation error can be “large” if $M\subset M_{0}$ as indicated in (2.5).
-To balance these properties in practice, it may be tempting to further experiment with the penalty.
-In an analogous theorem of Bai, (2004), it is required that $p_{N,T}\to\infty$ and $p_{N,T}/\ln(T)\to 0$ if the criterion is based on a logarithmic-valued sum of squared residuals as is the case here (see footnote 3 in that study).
-Importantly, this condition does not interfere with consistency in our Theorem 1, because clearly $p_{N,T}=o(\ln(T))$ still holds.
+Consequently, two forces are in effect: $p_{N,T}$ should still be negligible, but the factor estimation error can be “large” if $M\subset M_{0}$ as indicated by the remainder in (A.9).
+To balance these properties in practice, it may be tempting to experiment further with the penalty.
+Indeed, the change of penalty is the key message of Theorem 1 in Bai, (2004) in the PC setting, where two conditions are satisfied: 1) $p_{N,T}\to\infty$, and 2) $p_{N,T}/\ln(T)\to 0$ which is given in footnote 3 of the study to adapt the penalty to the logarithmic sum of squared residuals, which is exactly our setting.
+Importantly, this condition does not interfere with consistency in our Theorem 1, because clearly $p_{N,T}=o(\ln(T))$ holds.
 Therefore, we compare both types of penalties in the simulations to illustrate the outcome if practitioners simply extrapolate recommendations from the PC literature.
@@
 Remark 1.
-Note that under stationarity, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}c>0$ when $M\subset M_{0}$, according to the results in DVS.
-Our Proposition 1 naturally accommodates this result under $\tau=0$, because, by following the same approximation steps, (2.5) becomes
-$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\sum_{j:\lambda>0}\ln\left(1+\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}})\right)+R_{N,T}\to_{p}c>0$
+Note that under stationarity, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}c>0$ when $M\subset M_{0}$, according to the results in DVS (and similarly in MW).
+Our Proposition 1 naturally accommodates this result under $\tau=0$, because, by following the same approximation steps, (A.9) becomes
+$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\sum_{j:\lambda>0}\ln\left(1+\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)+R_{N,T}\to_{p}c>0$
 as $(N,T)\to\infty$ because each summand is greater than 1 w.p.
 1.
-Also, $|R_{N,T}|=O_{p}(N^{-1})+O_{p}((NT)^{-1/2})$, which is the same as in DVS under $\tau=0$, as expected from the discussion of (2.5).
+Here, $\boldsymbol{\Omega}_{M.M_{0}}^{0}=\operatorname*{plim}_{(N,T)\to\infty}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}$ (positive semi-definite, non-zero).
+Also, $|R_{N,T}|=O_{p}(N^{-1})+O_{p}((NT)^{-1/2})$, which is the same as in DVS (and MW) under $\tau=0$, as expected from the discussion of (A.9).
 Remark 2.
@@
 The factor loadings are generated as $\boldsymbol{\Gamma}_{i}=[\mathbf{I}_{m},\mathbf{0}_{m\times(k-m)}]\psi_{i}$, that is, a matrix $m\times k$, with $\psi_{i}\sim N(1,1)$ and the elements of the vector $1\times m$ $\boldsymbol{\gamma}_{i}$ are drawn from $N(1,1)$.
-In the case of time- and space-independent errors, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$.
-In the case of errors correlated weakly across time and cross-section we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
+In the case of errors independent over time and cross-sectionally, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$.
+In the case of errors correlated weakly across time and/or cross-sections we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
 $\mathbf{e}_{i,t}=(\epsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}=\mathbf{P}_{\rho}\mathbf{e}_{i,t-1}+\mathbf{z}_{\mathbf{w},i,t},\hskip 5.69054pt\mathbf{P}_{\rho}=\begin{bmatrix}\rho&\mathbf{0}_{1\times k}\\ \mathbf{0}_{k\times 1}&\rho_{v}\mathbf{I}_{k}\end{bmatrix},\hskip 5.69054pt\mathbf{z}_{\mathbf{w},i,t}=\begin{bmatrix}\sqrt{\frac{m\left(1-\rho^{2}\right)}{1+2J\kappa^{2}}}\boldsymbol{\epsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa\mathbf{w}_{i}^{\prime}\right)\\ \sqrt{\frac{\left(1-\rho_{v}^{2}\right)}{1+2J_{v}\kappa_{v}^{2}}}\boldsymbol{\Upsilon}_{t}^{\prime}\left(\boldsymbol{\iota}_{N,i}+\kappa_{v}\boldsymbol{\tilde{\mathbf{w}}}_{i}^{\prime}\right)\end{bmatrix}$ (3.2)
 where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero.
-$\rho$ is set to 0.5, $\kappa=0.2$ and $J=5$.
-The idiosyncratics in $\mathbf{X}_{i}$, $\mathbf{V}_{i}$, are also drawn as potentially weakly correlated over time and cross-sections.
-Here, $\mathbf{\Upsilon}_{t}$ is a $N\times k$ matrix, which stacks $\boldsymbol{\upsilon}_{i,t}\sim N(\mathbf{0}_{k},\mathbf{I}_{k})$ over $i$, and it is independent of $\boldsymbol{\epsilon}_{t}$.
+$\rho$ and $\rho_{v}$ control the correlation over time in $\boldsymbol{\varepsilon}_{t}$ and $\mathbf{V}_{i}$, respectively.
+For this simulation exercise, we set them to zero.
+In the Supplement we explore settings with $\rho=\rho_{v}=0.5$, an extension to our theory which prescribes uncorrelated innovations.
+Across all specifications, we allow for weak correlation across units with $\kappa=\kappa_{v}=0.2$, $J=J_{v}=5$ and $\mathbf{w}_{i}=\tilde{\mathbf{w}}_{i}$.
+$\mathbf{\Upsilon}_{t}$ is a $N\times k$ matrix, which stacks $\boldsymbol{\upsilon}_{i,t}\sim N(\mathbf{0}_{k},\mathbf{I}_{k})$ over $i$, and it is independent of $\boldsymbol{\epsilon}_{t}$.
 The vector $\boldsymbol{\iota}_{N,i}\in\mathbb{R}^{N\times 1}$ contains 1 in the $i$-th coordinate, and zeros elsewhere.
-$\rho_{v}=0$ for the time IID case and $\rho_{v}=0.5$ for weakly time-correlated idiosyncratics.
-In both cases, we allow for weak correlation across cross-sections, $\kappa_{v}=0.2$, $J_{v}=5$ and $\tilde{\mathbf{w}}_{i}$ is defined as $\mathbf{w}_{i}$ above.
-Therefore, we also test correlated idiosyncratics over time, whereas our theory prescribes uncorrelated innovations for $\tau\in(0,1)$.
 The factors are drawn as $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t}$, where
-$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=diag(g_{1},...,g_{m}),g_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.3)
+$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=\mathrm{diag}(g_{1},...,g_{m}),g_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\left(\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}\right)^{1/2}\right).$ (3.3)
 In the stationary factors case, the factors are drawn as a stable VAR with $\tau=0$.
-Two cases for non-stationary factors are considered: the non-stationary case of $\tau=0.4$ and the case close to local-to-unity process $\tau=0.9$.
-All simulations have 1000 repetitions and we compare DVO and MW.
+Two cases of non-stationary factors are considered: the non-stationary case of $\tau=0.4$, and the case close to local-to-unity process $\tau=0.9$.
+All simulations have 1000 repetitions, and we compare DVS and MW.
 In the Supplement, we provide additional simulation evidence of the eigenvalue growth ratio (ER) based on Juodis, (2022) for comparison.
@@
 Nevertheless, its performance crumbles with $\tau\to 1$ and does not improve as $(N,T)\to\infty$.
-Table 3 below presents the frequency of the correctly selected number of CAs, $g=4$, with the default penalties, $p_{N,T,1}$ and $p_{N,T,2}$ from 2.2, and the penalty from Bai, (2004), $\widetilde{p}_{N,T}$, for the case of stationary factors ($\tau=0$), the mildly non-stationary case ($\tau=0.4$) and the close to local-to-unity process.
-The idiosyncratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units.
-The Supplement presents additional results with weak correlation over time.
-We start with case of stationary factors established in the literature and the default penalties.
-In the small $T$ and $N$ case, $T=N=50$, and for stationary factors, we observe a frequency between $16\%$ and $36\%$, a pattern which resembles the results in MW.
+Table 3 below presents the frequency of the correctly selected number of CAs ($g=4$) with default penalties $p_{N,T,1}$ and $p_{N,T,2}$ from (2.2), and the penalty from Bai, (2004), $\widetilde{p}_{N,T}$, for the case of stationary factors ($\tau=0$), the mildly non-stationary case ($\tau=0.4$) and the close to local-to-unity process.
+The idiosyncratic terms $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units.
+In what follows, the low selection frequency reported is dominated by underslection.
+We start with the case established in the literature, which is stationary factors and default penalties.
+For small $N$ and $T$ ($T=N=50$) and stationary factors, we observe a frequency between $16\%$ and $36\%$, a pattern which resembles the results in MW.
 [^5] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$.
 If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors.
-The selection frequency decreases to zero in the case of non-stationary factors.
 As expected, consistency, illustrated here by a selection frequency of 100%, is quickly achieved when $N$ and $T$ increase.
-In the non-stationary cases, consistency requires much larger combinations of $(N,T)$.
-For the extreme case of a process close to local-to-unit, combinations of $(N,T)>300$ are required.
+We now turn to the case of non-stationary factors.
+For small $N$ and $T$, the selection frequency decreases to zero, which is driven by underselection of the number of CAs, as we show in Section 4 in the Supplement.
+To gain consistency, much larger combinations of $(N,T)$ are required.
+For the extreme case of a process close to local-to-unity, combinations of $(N,T)>300$ are integral.
+The behavior of both criteria with respect to increases in $N$, $T$, and $\tau$ is almost identical.
 Bai, (2004) proposes to adjust the penalty by $\ln(T)$ in the presence of non-stationary factors.
 The lower panel of Table 3 presents the frequency of the correctly selected number of CAs with a penalty of $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(\frac{NT}{N+T})$ for $\mathrm{IC}_{1}$ and $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(C^{2}_{N,T})$ for $\mathrm{IC}_{2}$.
-We observe for all four ICs a selection frequency of 0 for almost all combinations of $N$ and $T$ and $\tau$.
+We observe a selection frequency of 0 for all four ICs and for almost all combinations of $N$ and $T$ and $\tau$.
 Only in the large sample cases of $N=500$ and $T>300$ and only under stationary factors the selection frequency reaches 100%.
-This strongly suggests that prescriptions from the PC literature do not directly apply to CAs, as ICs are clearly inconsistent, and over-selection should not be heavily penalized.
-Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$ (”small sample”) and $N=T=500$ (”large sample”), where we increase $\tau$ in $0.01$ steps from 0 to $0.95$.
-As in Table 3 the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ are uncorrelated over time, but weakly correlated across units.
-For the small sample, the share of misselected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint.
-Then, it increases to almost $100\%$ with swings showcasing that the criteria become more unstable as $\tau$ increases.
-As expected, the degree of non-stationarity is much less of a problem for the large sample, and we observe misselection and instability of the criteria only for a high degree of non-stationarity.
-To shed more light on the importance of the sample size, we present in Figure 2 the share of misselected number of CAs with varying number of time periods from 25 to 500, $\tau=0.9$ and $N=[100,500]$.
-The share of misselection remains high for $N=100$ for both ICs, but declines for $N=500$.
-Again, this presents further evidence of the consistency of the ICs, but large samples are required to use them effectively.
+This strongly suggests that prescriptions from the PC literature do not directly apply to CAs, as ICs perform even worse, and selecting more CAs should not be heavily penalized.
+In the Supplement, we also investigate the share of misselected number of cross-sectional averages when increasing $\tau$ using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ for a small ($N=T=100$) and large sample ($N=T=500$).
+For the small sample, the share remains relatively flat until $\tau$ reaches levels of around $0.5$.
+As expected, this is much less of a problem for the large sample and we observe under-selection of CA only for a high degree of non-stationarity.
 Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
@@
 Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ uncorrelated over time, but weakly correlated across units, see (3.2).
-(a) $IC^{DVS}_{1}$
-(b) $IC^{MW}_{1}$
-Figure 1: Share of incorrect selected cross-sectional averages with increasing $\tau$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023).
-Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time but weakly correlated across units.
-(a) $IC^{DVS}_{1}$
-(b) $IC^{MW}_{1}$
-Figure 2: Share of incorrect selected cross-sectional averages with increasing $T$, $\tau=0.9$.
-DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023).
-Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time but weakly correlated across units.
 4 Conclusion
-This study is the first to examine selection of an optimal set of CAs in CCE and related settings by ICs inspired by Bai and Ng, (2002) when latent factors are non-stationary.
-In particular, we use mild integration in order to explore varying degrees of factor persistence, and demonstrate that ICs remain consistent without any modifications.
-However, the more persistent common factors are, the worse small sample performance becomes, and ICs regain selection consistency only in very large samples (i.e.
+This study is the first to examine the selection of an optimal set of CAs in CCE and related settings by ICs inspired by Bai and Ng, (2002) when latent factors are non-stationary.
+In particular, we use mild integration to explore varying degrees of factor persistence and demonstrate that ICs remain consistent without any modifications.
+However, the more persistent common factors are, the worse their small sample performance becomes, and ICs regain selection consistency only in very large samples (i.e.
 $(N,T)>300$, according to our experiments).
-Importantly, a diverging penalty suggested by Bai, (2004) under non-stationary factors for PC renders ICs inconsistent in the case of CAs.
-Therefore, our recommendation for CCE practitioners is not to take prescriptions from the PC literature automatically, and interpret IC results with caution in presence of highly persistent data, unless $N,T$ are substantial.
+Importantly, a divergent penalty suggested by Bai, (2004) to address non-stationary factors in PC makes our ICs perform even worse in the case of CAs.
+Therefore, our recommendation for CCE/CAs practitioners is not to automatically take PC literature prescriptions and interpret IC results with caution in the presence of highly persistent data, unless $N,T$ are substantial.
 5 Supplement
@@
 Throughout our analysis, we employ the following set of assumptions.
-Assumption 1.
-  • (a)
-    $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009). In particular,
-    $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\frac{\mathbf{G}}{T^{\tau}},\quad\mathbf{G}>0\hskip 5.69054pt(diagonal),$
+Assumption 1. $\{\mathbf{f}_{t}\}$ is a mildly integrated process as defined in Magdalinos and Phillips, (2009), such that
+$\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t},\quad\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=\mathrm{diag}(g_{1},\ldots,g_{m}),\hskip 5.69054ptg_{j}\in(0,2),$
 where $\mathbf{u}_{f,t}$ is a zero-mean linear process.
-  • (b)
-    If $\tau=0$, then $\{\mathbf{f}_{t}\}$ is a stationary process with summable autocovariances.
 Assumption 2. Let $\mathbf{e}_{i,t}=(\varepsilon_{i,t},\mathbf{v}_{i,t}^{\prime})^{\prime}\in\mathbb{R}^{k+1}$. Then
@@
 which implies that $\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$.
-5.3 Proof of Proposition 1
-In the proof we focus on the IC of De Vos and Stauskas, (2024) ($j=DVS$).
+5.3 Proposition 1 and the Proof
+Proposition 1. Under Assumptions 1-4 as $(N,T)\to\infty$, we have that:
+$\textit{(a)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM_{0}\subset M,\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=O_{p}(C_{N,T}^{-2});$ (A.8)
+$\textit{(b)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM\subset M_{0},\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\tau b\ln(T)+\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)+R_{N,T},$ (A.9)
+where $\boldsymbol{\Omega}_{M.M_{0}}^{0}=\operatorname*{plim}_{(N,T)\to\infty}\boldsymbol{\Omega}_{M.M_{0}}$ (positive semi-definite, non-zero), $|R_{N,T}|=o_{p}(1)$ is the remainder, $\lambda_{j}(\mathbf{A})$ is the $j$-th eigenvalue of $\mathbf{A}$, while $b$ is a number of strictly positive eigenvalues.
+Proof.
+In the proof, we focus on the IC of De Vos and Stauskas, (2024) ($j=DVS$).
 Therefore, we will have $K=k$ throughout.
@@
 Further,
-$\left\|\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]\right\|$ $\leq\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|T^{\tau}\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|$
+$\left|\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]\right|$ $\leq\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|T^{\tau}\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|$
 $=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$
@@
 5.3.3 Proof Changes for IC of Margaritella and Westerlund, (2023)
-5.3.3.1  Homogeneous $\boldsymbol{\beta}$
+Homogeneous $\boldsymbol{\beta}$
 Note that IC of MW is based on the residuals from the first stage CCE regression under the full set of available $k+1$ CAs (note that $|M|=k+1$ now).
@@
 $\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
-$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
+$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
+$-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
 $+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
@@
 and the overspecification risk is eliminated.
-5.3.3.2  Heterogeneous $\boldsymbol{\beta}$
+Heterogeneous $\boldsymbol{\beta}$
 If the parameter vector is heterogeneous, e.g. $\boldsymbol{\beta}_{i}=\boldsymbol{\beta}+\boldsymbol{\upsilon}_{i}$, where $\boldsymbol{\upsilon}_{i}$ is IID and mean-zero, then
@@
 Finally,
-$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|=\begin{cases}O_{p}(T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.98)
+$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|=\begin{cases}O_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.100)
 and, in particular, under $M\subset M_{0}$:
@@
 $=d_{NT}+o_{p}(T^{-\tau})$ (A.101)
-with $d_{NT}\to_{p}d>0$ in analogy to the analysis in (5.3.3.1).
+with $d_{NT}\to_{p}d>0$ in analogy to the analysis in (5.3.3).
 Now, we need to assume $T^{\tau}N^{-1/2}\to 0$ in addition to $\tau\in(0,0.5)$ to isolate the $o_{p}(T^{-\tau})$ terms.
@@
 $\mathbb{P}\left(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))<0)\right)\to 0,$ (A.106)
-as expected.
+as expected. Note that the restrictions on $\tau$ discussed here are just sufficient conditions, since we do not see such need in the simulationd exercises.
 5.4 Additional Monte Carlo Results
-5.4.1 Independent Idiosyncratics
+5.4.1 Idiosyncratics Independent Over Time
 Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
@@
 Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
-5.4.4 Correlated Idiosyncratics
+5.4.4 Idiosyncratics Correlated Over Time
 Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
@@
 Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
+5.4.6 Mis-selection with Varying $\tau$
+Figure 1 presents the share of mis-selected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$ (”small sample”) and $N=T=500$ (”large sample”), where we increase $\tau$ in $0.01$ steps from 0 to $0.95$.
+As in Table 1 of the manuscript the idiosyncratics $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units.
+For the small sample, the share of mis-selected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint.
+Then, it increases to almost $100\%$ with swings showcasing that the criteria become more unstable as $\tau$ increases.
+As expected, the degree of non-stationarity is much less of a problem for the large sample, and we observe misselection and instability of the criteria only for a high degree of non-stationarity.
+To shed more light on the importance of the sample size, we present in Figure 2 the share of misselected number of CAs with varying number of time periods from 25 to 500, $\tau=0.9$ and $N=[100,500]$.
+The share of misselection remains high for $N=100$ for both ICs, but declines for $N=500$.
+Again, this presents further evidence of the consistency of the ICs, but large samples are required to use them effectively.
+(a) $IC^{DVS}_{1}$
+(b) $IC^{MW}_{1}$
+Figure 1: Share of incorrect selected cross-sectional averages with increasing $\tau$.
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023).
+Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time but weakly correlated across units.
+(a) $IC^{DVS}_{1}$
+(b) $IC^{MW}_{1}$
+Figure 2: Share of incorrect selected cross-sectional averages with increasing $T$, $\tau=0.9$.
+DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023).
+Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ uncorrelated over time but weakly correlated across units.